### v1

# Non-Gaussian Simultaneous Autoregressive Models with Missing Data

Abstract
Standard simultaneous autoregressive (SAR) models are usually assumed to have normally distributed errors, an assumption that is often violated in real-world datasets, which are frequently found to exhibit non-normal, skewed, and heavy-tailed characteristics. New SAR models are proposed to capture these non-Gaussian features. In this project, the spatial error model (SEM), a widely used SAR-type model, is considered. Three novel SEMs are introduced that extend the standard Gaussian SEM by incorporating Student’s $t$-distributed errors after a one-to-one transformation is applied to the response variable. Variational Bayes (VB) estimation methods are developed for these models, and the framework is further extended to handle missing response data. Standard variational Bayes (VB) methods perform well with complete datasets; however, handling missing data requires a Hybrid VB (HVB) approach, which integrates a Markov chain Monte Carlo (MCMC) sampler to generate missing values. The proposed VB methods are evaluated using both simulated and real-world datasets, demonstrating their robustness and effectiveness in dealing with non-normal data and missing data in spatial models. Although the method is demonstrated using SAR models, the proposed model specifications and estimation approaches are widely applicable to various types of models for handling non-Gaussian data with missing values.
Keywords: Spatial error models; Student’s $t$ errors; Yeo and Johnson (YJ) transformation; Missing not at random; Variational Bayes approximation
1 Introduction
Conventional spatial statistical models, such as simultaneous autoregressive (SAR) models, conditional autoregressive (CAR) models, and kriging models, among others (see Besag, (1974); Cressie, (1993); Anselin, (1988)), typically assume that the data being modeled follow a Gaussian distribution. However, in practice, this assumption is often violated. In this study, we propose novel SAR-type models and their estimation methods, which are particularly useful when the data are non-Gaussian.
SAR models are a broad class of spatial statistical models commonly used for analysing spatially correlated lattice data (Cressie,, 1993). Examples of SAR models includes spatial error models (SEMs), spatial autoregressive models (SAMs), and spatial Durbin models (SDMs), and many more; see Anselin, (1988), and Cressie, (1993). These models account for spatial dependence, making them useful tools for understanding spatial phenomena in fields such as ecology (Ver Hoef et al.,, 2018), social sciences (Chi and Zhu,, 2008, 2019), finance (Calabrese et al.,, 2017; Agosto et al.,, 2019).
The estimation of SAR models using methods such as the maximum likelihood estimation (MLE) (Ord,, 1975; LeSage,, 1997) and the Bayesian methods (De Oliveira and Song,, 2008; Wu,, 2018), under the assumption of Gaussian errors, has been extensively discussed in the literature. To address the issue of outliers in estimating SEM, Yildirim and Mert Kantar, (2020) proposed a robust version of the ML estimator. Estimation methods for SAR models that do not rely on strong distributional assumptions have been developed. For example, Lee, (2004) introduced the quasi-maximum likelihood estimator (QMLE) for the SAM. Furthermore, Generalised Method of Moments (GMM) estimators, which require fewer distributional assumptions than traditional ML estimators, are robust to unknown heteroskedasticity and deviations from normality. These estimators have been proposed by Kelejian and Prucha, (1998), Lee and Liu, (2010), and Breitung and Wigger, (2018). Non-parametric estimation approaches have also been discussed in the literature for modeling SEM with non-Gaussian data (Al-Momani et al.,, 2017).
Missing values frequently occur in spatial datasets. When estimating SAR models, ignoring missing response values can lead to inconsistency and bias  (Wang and Lee,, 2013; Benedetti et al.,, 2020). Various estimation methods for SAR models with randomly missing data have been well developed (see LeSage and Pace, (2004); Wang and Lee, (2013); Suesse and Zammit-Mangion, (2017); Suesse, (2018); Wijayawardhana et al., (2024)). However, research on estimating SAR models with non-randomly missing data remains relatively limited. Flores-Lagunes and Schnier, (2012) introduced a GMM estimator, while Rabovič and Čížek, (2023) proposed a partial ML method for SEM. Furthermore, Doğan and Taşpinar, (2018) and Seya et al., (2021) used the Metropolis-Hastings (MH) algorithm, which becomes computationally expensive when dealing with a large number of observations.
Variational Bayes (VB) methods have recently emerged as a faster alternative to Markov chain Monte Carlo (MCMC) for estimating complex statistical models (Chappell et al.,, 2008; Han et al.,, 2016; Ong et al.,, 2018; Gunawan et al.,, 2021, 2024). Several commonly used VB methods include mean-field variational Bayes (MFVB) (Ormerod and Wand,, 2010), integrated non-factorised variational Bayes inference (INFVB) (Han et al.,, 2013), and Gaussian variational approximation (Ong et al.,, 2018; Tan and Nott,, 2018). Although VB methods offer a more computationally efficient alternative to MCMC, their use in estimating SAR models has been relatively limited. Wu, (2018) applied two VB approaches, a hybrid MFVB method and an INFVB method, to estimate spatial autoregressive confused (SAC) and matrix exponential spatial specification (MESS) models. Similarly, Bansal et al., (2021) used MFVB and INFVB to estimate spatial count data models, incorporating a MESS model to capture spatial dependence in error terms. More recently, Wijayawardhana et al., (2025) introduced novel VB methods for estimating SAR models with missing data, demonstrating their effectiveness for moderately large datasets with a high proportion of missing values.
Our article makes several key contributions. First, we introduce novel SAR models designed to capture non-Gaussian characteristics of the response variable, such as skewness and heavy tails. These models extend the standard Gaussian SAR by incorporating Student’s t-distributed errors and applying a one-to-one transformation, the Yeo-Johnson (YJ) transformation (Yeo and Johnson,, 2000), to the response variable (see Section 2.2 for details). By combining the properties of the Student’s $t$ distribution and the YJ transformation, these models effectively account for heavy-tailedness and skewness, making them well-suited for complex real-world data. Although we illustrate these new model specifications using SEMs, the framework is applicable to any SAR model.
Second, we propose efficient VB methods for estimating these new SEMs, with and without missing values in the response variable. For the missing data case, we specifically consider the missing not at random (MNAR) mechanism, as defined by Rubin, (1976), where the probability of a missing value depends on both observed and unobserved responses. In standard VB inference, the joint posterior distribution of missing values and model parameters is approximated using a Gaussian variational approximation (Wijayawardhana et al.,, 2025). However, it has been shown that the Gaussian variational approximation with a factor covariance structure (Ong et al.,, 2018), is not sufficiently flexible to jointly approximate the posterior distributions of the SAR model parameters and missing values. Therefore, in this study, we use the standard Gaussian variational approximation of Ong et al., (2018) for SEMs without missing values, while SEMs with missing values are estimated using an extended version of the hybrid VB (HVB) algorithm of Wijayawardhana et al., (2025).
The rest of this paper is organised as follows. Section 2 presents the new SEMs. In Section 3, we present the VB methods to estimate the SEMs with missing data. In Section 4, simulation studies are conducted to evaluate the performance of the new SEMs and VB methods. Section 5 applies the SEMs to a real-world dataset. Section 6 discusses our major results and findings. The paper also has an online supplement with additional technical details and examples.
2 Simultaneous Autoregressive Models
This section introduces our proposed SAR models, focusing on the spatial error model (SEM), a widely used SAR-type model. We present three novel SEMs that extend the standard Gaussian SEM.
Let $\textbf{y}^{\ast}=(y^{\ast}_{1},y^{\ast}_{2},...,y^{\ast}_{n})^{\top}$ be the vector of response variable observed at $n$ spatial locations $s_{1},\ldots,s_{n}$, X be the $n\times(r+1)$ design matrix containing the covariates, and W be the $n\times n$ spatial weight matrix. The SEM is given by
$\textbf{y}^{\ast}=\textbf{X}\boldsymbol{\beta}+(\textbf{I}_{n}-\rho\textbf{W})^{-1}\textbf{e},$ (1)
where e is an $n\times 1$ vector of error terms, $\boldsymbol{\beta}=(\beta_{0},\beta_{1},\ldots,\beta_{r})^{\top}$ is an $(r+1)\times 1$ vector of fixed effects parameters, $\rho$ is the spatial autocorrelation parameter, which measures the strength and direction of spatial dependence (LeSage and Pace,, 2009), and $\textbf{I}_{n}$ is the $n\times n$ identity matrix.
Section 2.1 discusses SEMs with Gaussian and Student’s $t$ errors. Section 2.2 discusses two SEMs with Yeo-Johnson (Yeo and Johnson,, 2000) transformation. Section 2.3 discusses SEMs with missing data.
2.1 Spatial error models with Gaussian and Student’s $t$ errors
The standard SEM assumes Gaussian errors $\textbf{e}\sim N(0,\sigma_{\textbf{e}}^{2}\textbf{I}_{n})$, where $\sigma_{\textbf{e}}^{2}$ is the variance parameter. This assumption implies that the distribution of $\textbf{y}^{\ast}$ is also Gaussian, with the mean vector and the covariance matrix given in Table 1. We call this model the SEM-Gau.
Student’s $t$-distribution has been widely used to model heavy-tailed data in popular statistical models, such as regression (Lange et al.,, 1989; Fernández and Steel,, 1999) and mixed-effects models (Pinheiro et al.,, 2001; Wang and Lin,, 2014). Huang et al., (2021) applied Student’s $t$ errors to spatial autoregressive scalar-on-function regression models. We now discuss SEM-t, which replaces the Gaussian errors in the SEM-Gau model with a Student’s t-distribution. Let $e_{i}\sim t_{\nu}(0,\sigma^{2}_{e})$, where $e_{i}$ is the $i^{th}$ element of the error vector e in Equation (1) and $t_{\nu}(0,\sigma^{2}_{e})$ is the Student’s t distribution with $\nu$ degrees of freedom ($\nu\geq 3$), mean zero, and scale parameter $\sigma^{2}_{e}$. We write Student’s t errors, $e_{i}$, $i=1,...,n$, as a scale mixture of normals (Chan et al.,, 2019),
$\begin{split}e_{i}&\mid\tau_{i},\sigma^{2}_{e}~{}\sim N(0,\sigma^{2}_{e}\tau_{i}),\\ \tau_{i}&\mid\nu\sim IG\left(\frac{\nu}{2},\frac{\nu}{2}\right),~{}~{}~{}~{}i=1,\ldots n\end{split}$ (2)
where $IG(a,b)$ denotes the inverse gamma distribution with shape parameter $a$ and scale parameter $b$. The conditional distribution of errors presented in Equation (2), can be written in vector notation as
$\textbf{e}\mid\boldsymbol{\tau},\sigma^{2}_{e}~{}\sim N(0,\sigma^{2}_{e}\boldsymbol{\Sigma}_{\boldsymbol{\tau}}),$ (3)
where the matrix $\boldsymbol{\Sigma}_{\boldsymbol{\tau}}=\textrm{diag}(\tau_{1},\ldots,\tau_{n})$. We now incorporate the error structure from Equation (3) into the SEM described in Equation (1). As a result, the distribution of $\textbf{y}^{\ast}$ given the latent vector $\boldsymbol{\tau}$ is also Gaussian, with the mean vector and the covariance matrix given in Table 1. We refer to this model as SEM-t. The log-likelihood function of $\textbf{y}^{\ast}$ for SEM-Gau and SEM-t is expressed as:
$\text{log}~{}p_{\textbf{y}^{\ast}}(\textbf{y}^{\ast}\mid\boldsymbol{\xi})=-\frac{n}{2}\textrm{log}(2\pi)-\frac{n}{2}\textrm{log}(\sigma^{2}_{\boldsymbol{e}})+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{1}{2\sigma^{2}_{\boldsymbol{e}}}\textbf{r}^{\top}\textbf{M}\textbf{r},$ (4)
where for the SEM-Gau model, $\boldsymbol{\xi}=\boldsymbol{\phi}$, while for the SEM-t model, $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$. The definitions for r, M, and $\boldsymbol{\phi}$ are given in Table 1.
2.2 Spatial error models with transformations
This section extends the SEM-Gau and SEM-t defined in Section 2.1. We form a novel class of SEMs using an element-wise Yeo and Johnson (YJ) transformation (Yeo and Johnson,, 2000) for the response variable, $\textbf{y}^{\ast}$, in both the SEM-Gau and SEM-t models. We assume that the $i^{th}$ observed response variable $(y_{i})$ is modeled as,
${y}_{i}=t_{{\gamma}}^{-1}(y^{\ast}_{i}),$
where $t_{{\gamma}}^{-1}(.)$ is the inverse function of the YJ transformation. For $20<\gamma<$ 2, it is defined as:
$y_{i}=t_{{\gamma}}^{-1}(y^{\ast}_{i})=\begin{cases}(y^{\ast}_{i}\gamma+1)^{1/\gamma}-1&\text{if }y^{\ast}_{i}\geq 0\\ 1-(-(2-\gamma)y^{\ast}_{i}+1)^{1/(2-\gamma)}&\text{if }y^{\ast}_{i}<0.\\ \end{cases}$ (5)
As the YJ transformation makes the data more symmetric and less skewed, applying its inverse transformation moves the data away from symmetry. This implies that $\textbf{y}=t_{{\gamma}}^{-1}(\textbf{y}^{\ast})=(t^{-1}_{{\gamma}}(y^{\ast}_{1}),t^{-1}_{{\gamma}}(y^{\ast}_{2}),\dots,t^{-1}_{{\gamma}}(y^{\ast}_{n}))^{\top}$ becomes an asymmetric, non-Gaussian response variable. The SEM with the YJ transformation (YJ-SEM) is defined as:
$\textbf{y}=t_{{\gamma}}^{-1}(\textbf{y}^{\ast})=t_{{\gamma}}^{-1}(\textbf{X}\boldsymbol{\beta}+(\textbf{I}_{n}-\rho\textbf{W})^{-1}\textbf{e}).\$ (6)
The density of y for the YJ-SEM is:
$p_{\textbf{y}}{(\textbf{y}\mid\boldsymbol{\xi})}=p_{\textbf{y}^{\ast}}({t_{\boldsymbol{\gamma}}(\textbf{y})})\prod_{i=1}^{n}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right),~{}~{}~{}~{}i=1,\ldots n,$ (7)
where $p_{\textbf{y}^{\ast}}(\cdot)$ is the density of $\textbf{y}^{\ast}$; see Section 2.1, and $\frac{dt_{\gamma}(y_{i})}{dy_{i}}$ is the derivative of the YJ transformation with respect to $y_{i}$; see Section S1 of the online supplement. The YJ-SEM with normally distributed errors (denoted as YJ-SEM-Gau) assumes that the error term in Equation (6) follows a normal distribution: $\mathbf{e}\sim N(0,\sigma_{\mathbf{e}}^{2}\mathbf{I}_{n})$. The YJ-SEM with Student’s t-distributed errors (denoted as YJ-SEM-t) assumes that the error term follows the structure specified in Equation (2). The log-likelihood function for y in YJ-SEM-Gau and YJ-SEM-t is:
$\begin{split}\text{log}~{}p_{\textbf{y}}(\textbf{y}\mid\boldsymbol{\xi})&=-\frac{n}{2}\textrm{log}(2\pi)-\frac{n}{2}\textrm{log}(\sigma^{2}_{\boldsymbol{e}})+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{1}{2\sigma^{2}_{\boldsymbol{e}}}\textbf{r}^{\top}\textbf{M}\textbf{r}\\ &+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right),\end{split}$ (8)
where for the YJ-SEM-Gau model, $\boldsymbol{\xi}=\boldsymbol{\phi}$, while for the YJ-SEM-t model, $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$. The definitions for r, M, and $\boldsymbol{\phi}$ are given in Table 1. Note that $y_{i}=y^{\ast}_{i}$ for $i=1,...,n$, corresponds to an identity transformation in both SEM-Gau and SEM-t.
The three SEMs, SEM-Gau, SEM-t, and YJ-SEM-Gau, are special cases of the YJ-SEM-t model. The SEM-Gau model is a YJ-SEM-t model with the degrees of freedom parameter ($\nu$) set to $\infty$ and the YJ parameter ($\gamma$) set to 1, while the YJ-SEM-Gau model is a YJ-SEM-t model with $\nu=\infty$ and $\gamma\neq 1$. Finally, the SEM-t model is also a YJ-SEM-t model with $\gamma=1$.
Table 1: The definitions of r, M, and the parameters $\boldsymbol{\phi}$ for the four spatial error models. Each SEM has a mean vector of $\mathbf{X}\boldsymbol{\beta}$ and a covariance matrix of the form $\boldsymbol{\Sigma}=\sigma^{2}_{\textbf{e}}\textbf{M}^{-1}$, with $\textbf{A}=\textbf{I}_{n}-\rho\textbf{W}$.
Model $\boldsymbol{\phi}$ r M
SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $^{\top}\textbf{A}$ A
SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$ A
YJ-SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e},},\rho,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $^{\top}\textbf{A}$ A
YJ-SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$ A
2.3 Spatial error models with missing responses
Consider that the response vector y of an SEM presented in Sections 2.1 and 2.2 contains missing values. Let $\textbf{y}_{o}$ be the subset of y with $n_{o}$ observed units, and $\textbf{y}_{u}$ be the subset of y with $n_{u}$ unobserved units. The complete response vector is $\textbf{y}=({\textbf{y}}_{o}^{\top},{\textbf{y}}_{u}^{\top})^{\top}$. A missing data indicator vector m of length ${n}$ containing 1’s and 0’s is defined. If an element in y is missing, then the corresponding element in m is 1 and 0, otherwise. In the presence of missing values, the vector r and the matrices X and M are decomposed into distinct components as follows:
$\textbf{r}=\begin{pmatrix}\textbf{r}_{o}\\ {\textbf{r}_{u}}\end{pmatrix},~{}\textbf{X}=\begin{pmatrix}\textbf{X}_{o}\\ {\textbf{X}_{u}}\end{pmatrix},~{}\textbf{M}=\begin{pmatrix}\textbf{M}_{oo}&\textbf{M}_{ou}\\ \textbf{M}_{uo}&\textbf{M}_{uu}\end{pmatrix},$ (9)
where $\textbf{r}_{o}$ and $\textbf{r}_{u}$ are subvectors of r, corresponding to the observed and unobserved responses, respectively. Similarly, $\textbf{X}_{o}$ and $\textbf{X}_{u}$ represent the design matrices associated with the observed and unobserved responses. Additionally, the matrices $\textbf{M}_{oo}$, $\textbf{M}_{ou}$, $\textbf{M}_{uo}$, and $\textbf{M}_{uu}$ are submatrices of M.
We assume missing responses follow a missing not at random (MNAR) mechanism (Rubin,, 1976). The missing data mechanism is defined by the conditional distribution of m given y, commonly referred to as the missing data model, denoted by $p(\textbf{m}|\textbf{y},\boldsymbol{\psi},\textbf{X}^{*})$, where $\textbf{X}^{*}$ is an $n\times(q+1)$ design matrix containing the covariates of the missing data model. These covariates may be a subset of those used in SEMs. Assuming the conditional independence of m given y, $\textbf{X}^{*}$, and $\boldsymbol{\psi}$, the density is the product of $p(m_{i}\mid y_{i},\textbf{x}^{*}_{i},\boldsymbol{\psi})$ for $i=1,\ldots,n$, where $m_{i}$ and $y_{i}$ are the $i^{th}$ elements of m and y, and ${\textbf{x}_{i}^{*}}^{\top}$ is the $i^{th}$ row of $\textbf{X}^{*}$. The parameter vector $\boldsymbol{\psi}=(\boldsymbol{\psi}_{\textbf{x}}^{\top},\psi_{\textbf{y}})^{\top}$ consists of the fixed effects vector associated with covariates $\textbf{X}^{*}$, denoted as $\boldsymbol{\psi}_{\textbf{x}}=(\psi_{0},\psi_{1},\psi_{2},\ldots,\psi_{q})^{\top}$, and the fixed effect corresponding to y, denoted as $\psi_{\textbf{y}}$. A logistic regression model is used as the missing data model and is given by:
$p(\textbf{m}\mid\textbf{y},\textbf{X}^{*},\boldsymbol{\psi})=\prod_{i=1}^{n}\frac{e^{({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\psi_{{y}}){m_{i}}}}{1+e^{{\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+\textbf{y}_{i}\psi_{\textbf{y}}}}.$ (10)
3 Variational Bayes inference for SEMs with missing data
Section 3.1 introduces the variational Bayes (VB) inference method. Section 3.2 discusses the efficient hybrid variational Bayes (HVB) methods for estimating the SEMs with missing data. Section 3.3 discusses the Deviance Information Criterion (DIC) to select the best SEMs. The VB method for the proposed SEMs in Section 2, when there are no missing data, is provided in Section S2 of the online supplement.
3.1 Variational Bayes inference
Variational Bayes (VB) inference approximates a Bayesian posterior distribution by formulating it as an optimisation problem, offering a more computationally efficient alternative to the computationally intensive Markov chain Monte Carlo (MCMC) methods for complex statistical models. This section presents the VB method for SEMs with missing data.
We assume that some values in the response vector of an SEM are missing under the missing not at random (MNAR) mechanism. Under this assumption, Bayesian estimation of the SEM requires consideration of the joint distribution of the missing data indicator vector m and the response vector y (Wijayawardhana et al.,, 2025), which is denoted by $p(\textbf{y},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})$. Using the selection model factorisation (Little and Rubin,, 2019), $p(\textbf{y},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})$ is expressed as
$p(\textbf{y},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})=p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\textbf{y}\mid\boldsymbol{\xi}).$ (11)
where $p(\textbf{y}\mid\boldsymbol{\xi})$ is the likelihood of y. The log-likelihood for y for different SEMs is given in Section 2. Additionally, $p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})$ denotes the density of the missing data model, given in Equation (10).
Consider Bayesian inference to estimate the joint posterior of $\boldsymbol{\xi}$, $\boldsymbol{\psi}$, and the missing values $\textbf{y}_{u}$. For the SEM-Gau and YJ-SEM-Gau models, $\boldsymbol{\xi}=\boldsymbol{\phi}$, while for the SEM-t and YJ-SEM-t models, $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$; see Table 1 for the definition of $\boldsymbol{\phi}$ for each SEM. Let the joint posterior distribution of $\boldsymbol{\xi}$, $\boldsymbol{\psi}$, and $\textbf{y}_{u}$ be denoted as $p(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m})$. We let $p(\boldsymbol{\xi})$ and $p(\boldsymbol{\psi})$ as the prior distributions of $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$, respectively. Using the selection model factorisation in Equation (11), the joint posterior distribution $p(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m})$ is given by
$p(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m})\propto p(\textbf{m}\mid{\textbf{y}},\boldsymbol{\psi})p({\textbf{y}}\mid\boldsymbol{\xi})p(\boldsymbol{\xi})p(\boldsymbol{\psi}).$ (12)
We define $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\xi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\xi})p(\boldsymbol{\psi})$. The term $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ for each SEM with missing data, along with the total number of parameters to be estimated ($s_{m}$) for each model are provided in Table 2.
Table 2: The term $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\xi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\xi})p(\boldsymbol{\psi})$, and the total number of model parameters $s_{m}$ (including the length of the latent vector $\boldsymbol{\tau}$) for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models with missing values. The distribution $p(\boldsymbol{\tau}\mid\boldsymbol{\phi})=\prod_{i=1}^{n}p(\tau_{i}\mid\nu)$ represents the $n$ independent inverse gamma latent variables used in SEMs with Student-$t$ errors (see Equation (2)).
Model $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ $s_{m}$
SEM-Gau $p({\textbf{y}}\mid\boldsymbol{\phi})p(\textbf{m}\mid{\textbf{y}},\boldsymbol{\psi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$ $r+q+5$
SEM-t $p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\textbf{m}\mid{\textbf{y}},\boldsymbol{\psi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$ $r+q+6+n$
YJ-SEM-Gau $p(\textbf{y}\mid\boldsymbol{\phi})p(\textbf{m}\mid{\textbf{y}},\boldsymbol{\psi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$ $r+q+6$
YJ-SEM-t $p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\textbf{m}\mid{\textbf{y}},\boldsymbol{\psi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$ $r+q+7+n$
[…]
$\mathcal{L}(\boldsymbol{\lambda})=E_{q}\left[\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})-\text{log}~{}q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})\right],$ (13)
where $E_{q}\left[\cdot\right]$ denotes the expectation with respect to $q_{\boldsymbol{\lambda}}$. Table 2 provides expressions for $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ for different SEMs.
To maximise ELBO given in Equation (S6) with respect to variational parameters, $\boldsymbol{\lambda}$, stochastic gradient ascent (SGA) methods are used (Nott et al.,2012Titsias and Lázaro-Gredilla,20142015). The SGA method updates the initial value for $\boldsymbol{\lambda}$ (say $\boldsymbol{\lambda}^{(0)}$) according to the iterative scheme,
[…]
3.2 Hybrid variational Bayes
To efficiently handle missing values $\textbf{y}_{u}$ in SEMs described in Section 2, we extend the hybrid VB method of Wijayawardhana et al., (2025) that uses the Gaussian variational approximation with a factor covariance structure to approximate the posterior distribution of the parameters, while employing an MCMC sampler to generate samples of $\textbf{y}_{u}$ from its conditional distribution, given $\textbf{y}_{o}$, m, $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$. We first factorise the variational distribution that approximates $\boldsymbol{\xi},\boldsymbol{\psi}$, and $\textbf{y}_{u}$, $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ as follows:
$q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p({\textbf{y}}_{u}\mid{\textbf{y}}_{o},\textbf{m},\boldsymbol{\xi,\boldsymbol{\psi}})q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi}),$ (15)
where $p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ represents the conditional distribution of $\textbf{y}_{u}$, given the observed data ($\textbf{y}_{o}$, m), and the model parameters ($\boldsymbol{\xi}$ and $\boldsymbol{\psi}$). The term $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ is the Gaussian variational approximation of Ong et al., (2018), used to approximate the posterior distribution of $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$. We rewrite the ELBO in Equation (S6) by substituting the factorised variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ from Equation (15). Using Bayes’ rule, it is straightforward to show
$\mathcal{L}(\boldsymbol{\lambda})=E_{q}\left[\text{log}~{}p({\textbf{y}}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})+\text{log}~{}p(\boldsymbol{\xi},\boldsymbol{\psi})-\text{log}~{}q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\right]=\mathcal{L}^{0}(\boldsymbol{\lambda}),$ (16)
where $\mathcal{L}^{0}(\boldsymbol{\lambda})$ is the ELBO resulting from approximating only the posterior distribution of parameters $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$, $p(\boldsymbol{\xi},\boldsymbol{\psi}\mid\textbf{y}_{o},\textbf{m})$, directly using the variational distribution $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$; see Wijayawardhana et al., (2025) for the complete proof.
We assume that the variational distribution, $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$, follows a multivariate normal distribution with a factor covariance structure (Ong et al.,, 2018), $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\sim N\left((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top};\boldsymbol{\mu}_{s_{m}},\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2}\right)$, where $\boldsymbol{\mu}_{s_{m}}$ is an $s_{m}\times 1$ vector of variational means, $\textbf{B}_{s_{m}}$ is an $s_{m}\times p$ matrix with the upper triangular elements set to zero, and $\textbf{D}_{s_{m}}$ is an $s_{m}\times s_{m}$ diagonal matrix with positive diagonal elements $\textbf{d}_{s_{m}}=(d_{1},\dots,d_{s_{m}})^{\top}$. For all SEMs with missing values, the value of $s_{m}$ is provided in Table 2. The vector of variational parameters is given by $\boldsymbol{\lambda}_{s_{m}}=\left(\boldsymbol{\mu}_{s_{m}}^{\top},\text{vech}(\textbf{B}_{s_{m}})^{\top},\textbf{d}_{s_{m}}^{\top}\right)^{\top}$, where the ’vech’ operator vectorises a matrix by stacking its columns from left to right while removing all the elements above the diagonal (the superdiagonal elements) of the matrix.
In the SGA algorithm (Equation (14)), we compute an unbiased estimate of the ELBO gradient, $\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\lambda)$, denoted as $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\lambda)}$, at each iteration using the reparameterisation trick (Kingma and Welling,, 2013). To do this, we first generate samples from the variational distribution $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ by drawing ${\boldsymbol{\delta}^{0}}=(\boldsymbol{\eta}^{0},\boldsymbol{\epsilon}^{0})\sim N(\mathbf{0},\mathbf{I}_{s_{m}+p})$, where $\boldsymbol{\eta}^{0}$ is a $p$-dimensional vector and $\boldsymbol{\epsilon}^{0}$ is an $s_{m}$-dimensional vector. Next, we compute $(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}=t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})=\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}$. The density of $\boldsymbol{\delta}^{0}$ is denoted by $f_{\boldsymbol{\delta}^{0}}(\boldsymbol{\delta}^{0})$. Let $\boldsymbol{\delta}=({\boldsymbol{\delta}^{0}}^{\top},\textbf{y}_{u}^{\top})^{\top}$, where the product density is expressed as $f_{\boldsymbol{\delta}}(\boldsymbol{\delta})=f_{\boldsymbol{\delta}^{0}}(\boldsymbol{\delta}^{0})p({\textbf{y}}_{u}\mid{\textbf{y}}_{o},\textbf{m},t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}}))$. Finally, the transformation from $\boldsymbol{\delta}$ to the parameter and augmented missing value space is given by $((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top},{\textbf{y}}_{u}^{\top})^{\top}=t(\boldsymbol{\delta},\boldsymbol{\lambda}_{s_{m}})=(t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top},\textbf{y}_{u}^{\top})^{\top}=((\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0})^{\top},\textbf{y}_{u}^{\top})^{\top}$.
The reparameterisation gradient of the ELBO in Equation (16) is derived by differentiating under the integral sign as shown below:
$\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})=\mathbb{E}_{f_{\boldsymbol{\delta}}}\left[\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top}}{d\boldsymbol{\lambda}_{s_{m}}}(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})-\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi}))\right],$ (17)
where $\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})}{d\boldsymbol{\lambda_{s_{m}}}}$ is the derivative of the transformation $t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})=\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}$ with respect to the variational parameters $\boldsymbol{\lambda}_{s_{m}}=(\boldsymbol{\mu}_{s_{m}}^{\top},\text{vech}(\textbf{B}_{s_{m}})^{\top},\textbf{d}_{s_{m}}^{\top})^{\top}$; see Section S4.1 of the online supplement for the proof. The expressions for $\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top}}{d\boldsymbol{\lambda}_{s_{m}}}$ and $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ can be found in Section S4.1 of the online supplement. In addition, the expressions for $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ are provided in Section S4.3 for different SEMs.
Algorithm 1 outlines the HVB algorithm. For all four SEMs with missing data, the conditional distribution of missing data; $p(\mathbf{y}_{u}\mid\mathbf{y}_{o},\mathbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ is not available in closed form. To sample from $p(\mathbf{y}_{u}\mid\mathbf{y}_{o},\mathbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ in step 5 of Algorithm 1, we use the Markov chain Monte Carlo (MCMC) steps outlined in Algorithm 2.
Algorithm 1 Hybrid variational Bayes (HVB) algorithm.
1:  Initialise $\boldsymbol{\lambda}_{s_{m}}^{(0)}=(\boldsymbol{\mu}_{s_{m}}^{\top(0)},\textrm{vech}{(\textbf{B}_{s_{m}})}^{\top(0)},\textbf{d}_{s_{m}}^{\top(0)})^{\top}$ and set $t=0$
2:  repeat
3:     Generate $({\boldsymbol{\eta}^{0}}^{(t)},{\boldsymbol{\epsilon}^{0}}^{(t)})\sim N(\textbf{0},\textbf{I}_{s_{m}+p})$
4:     Generate $(\boldsymbol{\xi}^{(t)^{\top}},\boldsymbol{\psi}^{(t)^{\top}})^{\top}\sim q_{\boldsymbol{\lambda}_{s_{m}}^{(t)}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ using its reparameterised representation.
5:     Generate $\textbf{y}_{u}^{(t)}~{}\sim p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi}^{(t)},\boldsymbol{\psi}^{(t)})$
6:     Construct unbiased estimates $\widehat{{\nabla_{\boldsymbol{\mu}_{s_{m}}}\mathcal{L}(\boldsymbol{\lambda})}},\widehat{{\nabla_{\textrm{vech}{(\textbf{B}_{s_{m}})}}\mathcal{L}(\boldsymbol{\lambda})}},\text{and}\widehat{{\nabla_{\textbf{d}_{s_{m}}}\mathcal{L}(\boldsymbol{\lambda})}}$ using Equations  (S44), (S45) and  (S46) in Section S4.1 of the online supplement.
[…]
11:     Set $\boldsymbol{\lambda}_{s_{m}}^{(t+1)}=(\boldsymbol{\mu}_{s_{m}}^{\top(t+1)},\textrm{vech}(\textbf{B}_{s_{m}})^{\top(t+1)},\textbf{d}_{s_{m}}^{\top(t+1)})^{\top}$, and $t=t+1$
12:  until some stopping rule is satisfied
Algorithm 2 MCMC steps within the $t^{th}$ iteration of the HVB algorithm.
1:  Initialise missing values $\textbf{y}_{u,0}~{}\sim p(\textbf{y}_{u}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{o})$
2:  for i=1, …, $N_{1}$ do
3:     Sample $\widetilde{\textbf{y}}_{u}$ from the proposal distribution $p(\widetilde{\textbf{y}}_{u}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{o})$.
4:     Sample $u$ from uniform distribution, $u~{}\sim\mathcal{U}(0,1)$
5:     Calculate $a=\text{min}\left(1,\frac{p(\textbf{m}\mid\widetilde{\textbf{y}},\boldsymbol{\psi}^{(t)})}{p(\textbf{m}\mid\textbf{y}_{{i-1}},\boldsymbol{\psi}^{(t)})}\right)$, where $\widetilde{\textbf{y}}=(\textbf{y}_{o}^{\top},{\widetilde{\textbf{y}}_{u}}^{\top})$ and $\textbf{y}_{i-1}=(\textbf{y}^{\top}_{o},{\textbf{y}^{\top}_{u,{i-1}}})$
6:     if $a>u$ then
7:        $\textbf{y}_{u,i}=\widetilde{\textbf{y}}_{u}$
8:     else
9:        $\textbf{y}_{u,i}=\textbf{y}_{u,i-1}$
10:     end if
11:  end for
12:  Output $\textbf{y}^{(t)}_{u}=\textbf{y}_{u,N_{1}}$
The MCMC steps in Algorithm 2 generate samples from the proposal distribution $p(\widetilde{\textbf{y}}_{u}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{o})$, which follows a multivariate Gaussian with the mean vector given by $\textbf{X}_{u}\boldsymbol{\beta}-\textbf{M}_{uu}^{-1}\textbf{M}_{uo}\textbf{r}_{o}$ and the covariance matrix given by $\sigma^{2}_{\textbf{e}}\textbf{M}_{uu}^{-1}$, for both SEM-Gau and SEM-t; see Table 1 and Equation (9) for details on the partitioning of r, X and M across different SEMs.
For YJ-SEM-Gau and YJ-SEM-t, the proposals are generated in two steps. First, we sample $\widetilde{\textbf{y}}_{u}^{*}$ from the conditional distribution $p({\widetilde{\textbf{y}}_{u}^{*}}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{o}^{*})$, which follows a multivariate Gaussian with the mean vector given by $\textbf{X}_{u}\boldsymbol{\beta}-\textbf{M}_{uu}^{-1}\textbf{M}_{uo}(\textbf{y}_{o}^{*}-\textbf{X}_{o}\boldsymbol{\beta})$ and the covariance matrix given by $\sigma^{2}_{\textbf{e}}\textbf{M}_{uu}^{-1}$, where $\textbf{y}_{o}^{*}=t_{\gamma}(\textbf{y}_{o})$. Then, we apply the inverse Yeo-Johnson (YJ) transformation to obtain the final proposal: $\widetilde{\textbf{y}}_{u}=t_{\gamma}^{-1}(\widetilde{\textbf{y}}_{u}^{*})$.
As $n$ and $n_{u}$ increase, the HVB algorithm implemented using the MCMC scheme outlined in Algorithm 2 fails to estimate the parameters accurately due to a low acceptance rate. To address this, we partition $\mathbf{y}_{u}$ into $k$ blocks and update one block at a time. The MCMC steps for sampling the missing values one block at a time are detailed in Algorithm S2 in Section S4.2 of the online supplement. We refer to the modified HVB algorithm with block-wise updates as HVB-AllB (which uses MCMC steps in Algorithm S2), and the original version without block-wise updates as HVB-NoB (which uses MCMC steps in Algorithm 2) in the following sections.
3.3 Bayesian model comparison
To assess model fit in both the simulation study (Section 4) and the real-world analysis (Section 5), we employ extended versions of the standard Deviance Information Criterion (DIC, Spiegelhalter et al., (2002)) proposed by Celeux et al., (2006). This section first outlines DIC methods for models without missing data, then for those with missing data.
For full data SEMs, both $\text{DIC}_{1}$ and $\text{DIC}_{2}$ are computed to evaluate model fit, with lower values indicating better fit. $\text{DIC}_{1}$ is defined as:
$\text{DIC}_{1}=-4E_{\boldsymbol{\phi}}[\text{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})]+2~{}\text{log}~{}p(\textbf{y}\mid\overline{\boldsymbol{\phi}}),$ (18)
where $\overline{\boldsymbol{\phi}}$ is the posterior mean of ${\boldsymbol{\phi}}$, and $p(\textbf{y}\mid\boldsymbol{\phi})$ denotes the density of y given the model parameters $\boldsymbol{\phi}$. Both the expectation and the posterior mean $\overline{\boldsymbol{\phi}}$ are computed using samples drawn from the approximate posterior distribution estimated by the proposed VB method. For all SEMs, the logarithm of the density $p(\textbf{y}\mid\boldsymbol{\phi})$ is provided in Section S5 of the online supplement.
$\text{DIC}_{2}$ is defined as:
$\text{DIC}_{2}=-4E_{\boldsymbol{\phi}}[\text{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})]+2~{}\text{log}~{}p(\textbf{y}\mid\widehat{\boldsymbol{\phi}}),$ (19)
where $\widehat{\boldsymbol{\phi}}$ is the value of ${\boldsymbol{\phi}}$ that maximises the function $p(\textbf{y}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$, estimated using samples drawn from the approximate posterior obtained by the proposed VB method. The expectation is also computed using these variational posterior samples.
$\text{DIC}_{5}$ is used to assess model fit for SEMs with missing data and is defined as:
$\text{DIC}_{5}=-4E_{\boldsymbol{\phi},\boldsymbol{\psi},\textbf{y}_{u}}[\text{log}~{}p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})]+2~{}\text{log}~{}p(\textbf{y}_{o},\widehat{\textbf{y}}_{u},\textbf{m}\mid\widehat{\boldsymbol{\phi}},\widehat{\boldsymbol{\psi}}),$ (20)
where $p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})$ is the joint density of y and m, conditional on model parameters $\boldsymbol{\phi}$ and $\boldsymbol{\psi}$. The values $\widehat{\boldsymbol{\phi}}$, $\widehat{\boldsymbol{\psi}}$, and $\widehat{\textbf{y}}_{u}$ are the posterior sample values of $\boldsymbol{\phi}$, $\boldsymbol{\psi}$ and $\textbf{y}_{u}$ that maximise the density $p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})p(\boldsymbol{\phi},\boldsymbol{\psi})$, where the samples are drawn from the approximate posterior estimated by the proposed HVB method. The expectation is also computed using these variational posterior samples. Note that $p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})=p(\textbf{y}\mid\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})$; see the selection model factorisation in Equation (11).
4 Simulation study
The simulation study is divided into two subsections. Section 4.1 compares the posterior distributions obtained using variational Bayes (VB) methods with those of Hamiltonian Monte Carlo (HMC, Neal, (1996)). The posterior densities estimated using the HMC method are considered as ground truth to assess the accuracy of the VB methods. In Section 4.1, we utilise a small simulated dataset ($n=625$), as implementing HMC is computationally intensive for SEMs, particularly when handling large datasets with missing data (Wijayawardhana et al.,, 2025). Section 4.2 uses two moderately large simulated datasets, each considered with and without missing values, and each exhibiting distinct characteristics, to evaluate the robustness of the proposed models and VB methods.
We now describe the prior distributions used in this study for both the VB and HMC methods, applied in the simulation and real-world examples presented in Sections 4.1, 4.2, and 5. To map the parameters $\sigma^{2}_{\textbf{e}}$, $\rho$, $\nu$, and $\gamma$ onto the real line, we apply the following transformations: $\omega^{\prime}=\text{log}(\sigma^{2}_{\textbf{e}})$, $\rho^{\prime}=\text{log}(1+\rho)-\text{log}(1-\rho)$, $\nu^{\prime}=\text{log}(\nu-3)$, and $\gamma^{\prime}=\text{log}(\gamma)-\text{log}(2-\gamma)$. Table S3 gives the prior distributions for the transformed parameters.
Parameter $\boldsymbol{\beta}$ $\omega^{\prime}$ $\rho^{\prime}$ $\nu^{\prime}$ $\gamma^{\prime}$ $\boldsymbol{\psi}$
Prior distribution $N(\textbf{0},\sigma^{2}_{\boldsymbol{\beta}}\textbf{I})$ $N(0,\sigma^{2}_{\omega^{\prime}})$ $N(0,\sigma^{2}_{\rho^{\prime}})$ $N(0,\sigma^{2}_{\nu^{\prime}})$ $N(0,\sigma^{2}_{\gamma^{\prime}})$ $N(\textbf{0},\sigma^{2}_{\boldsymbol{\psi}}\textbf{I})$
Hyperparameters $\sigma^{2}_{\boldsymbol{\beta}}=10^{2}$ $\sigma^{2}_{\omega^{\prime}}=10^{2}$ $\sigma^{2}_{\rho^{\prime}}=10^{2}$ $\sigma^{2}_{\nu^{\prime}}=10^{2}$ $\sigma^{2}_{\gamma^{\prime}}=10^{2}$ $\sigma^{2}_{\boldsymbol{\psi}}=10^{2}$
Table 3: Prior distributions of model parameters
4.1 Assessing the accuracy of variational Bayes approximations
This section compares the accuracy of the proposed VB approximation methods with the computationally intensive, exact HMC methods. The comparison focuses solely on the YJ-SEM-Gau model, both with and without missing data. To facilitate this comparison, a dataset with 625 observations is simulated from the YJ-SEM-Gau model with five covariates generated from standard normal distributions. The weight matrix, W is constructed based on a $25\times 25$ regular grid, where neighbouring units are defined using the rook neighbourhood criteria, as described in Lloyd, (2010).
The model parameters are specified as follows: the fixed effects ($\boldsymbol{\beta}$) are randomly drawn from discrete uniform values between $-3$ and $3$ (excluding 00). The error variance parameter is set to $\sigma^{2}_{\mathbf{e}}=1$ and the spatial autocorrelation parameter is set to $\rho=0.8$. The Yeo Johnson (YJ) parameter is specified as $\gamma=1.25$, inducing moderate left skewness into the simulated response variable. See Figure S1 in Section S7.1 of the online supplement for the kernel density plot of the response variables.
The VB algorithm, outlined in Algorithm S1 in Section S2 of the online supplement, and the HMC algorithm are used to fit the YJ-SEM-Gau model to the simulated dataset. The initial values for the VB algorithm are set as follows: The variational mean vector ($\boldsymbol{\mu}$) is initialised using the estimates of $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ obtained from fitting the SEM-Gau model via maximum likelihood estimation (MLE) method. The initial value for $\gamma$ is set to 1. For the variational covariance matrix parameters, the elements of B and the diagonal elements of D are all initialised to 0.01. We use $p=4$ factors. The results do not improve when we increase the number of factors. We used similar starting values for the HMC algorithms as well.
Both the VB and HMC algorithms are run for 10,000 iterations, at which point convergence is achieved. For the HMC algorithm, the first 5,000 iterations are treated as burn-in and discarded. Convergence for the VB algorithm is assessed visually by inspecting the plots of variational means over iterations. For the HMC algorithms, convergence is assessed by inspecting trace plots of the model parameters (see Section S9.1 of the online supplement).
The HMC algorithm directly generates posterior samples of the model parameters. In contrast, the VB algorithm first estimates the variational parameters, $\boldsymbol{\lambda}$, by running the variational optimisation. Then, 10,000 posterior draws of the model parameters are generated from the variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$.
Figure 1 compares the posterior densities of selected model parameters estimated using the VB method and the HMC algorithm implemented via RSTAN (Stan Development Team,, 2023) to fit the simulated dataset. For the fixed effect parameter (${\beta}_{1}$), the spatial autocorrelation parameter ($\rho$), the variance parameter ($\sigma^{2}_{\textbf{e}}$), and the YJ parameter $\gamma$, the posterior densities estimated by both methods are nearly identical and include the true values.
Figure 1: Comparison of posterior densities of selected parameters in the YJ-SEM-Gau model, obtained using the HMC and VB methods on the simulated dataset without missing values. The vertical line represents the true parameter value.
Next, we introduce missing values into the response variable y under the MNAR mechanism, using the logistic regression model described in Section 2.3. To do this, we generate an additional covariate $\textbf{x}^{*}$ from a log-normal distribution with a log-scale mean of 0 and a log-scale standard deviation of 1. This covariate is included in the missing data model, which is specified as a logistic regression model with $\textbf{x}^{*}$ and the response variable y as covariates (see Equation (10)). The regression coefficients are set as follows: $\psi_{0}=-1.0$ (intercept), $\psi_{\textbf{x}^{*}}=0.5$ (the coefficient of $\textbf{x}^{*}$), and $\psi_{\textbf{y}}=-0.1$ (the coefficient of y). This configuration yields approximately $50\%$ missing values in y, corresponding to $n_{u}=320$.
The HVB-NoB algorithm (without block-wise update) described in Algorithm 1 and the HMC algorithm (implemented via RSTAN) are used to estimate the YJ-SEM-Gau model. The tuning parameter, $N_{1}$ of the HVB-NoB algorithm, is set to 10. The initial values for the algorithm are set as follows: the variational mean vector ($\boldsymbol{\mu}$) is initialised using the estimates of $\boldsymbol{\beta}$, $\sigma^{2}_{\mathbf{e}}$, and $\rho$ obtained by fitting the SEM-Gau model via MLE method. The initial value of the transformation parameter $\gamma$ is set to 1. All fixed-effect parameters of the missing value model ($\boldsymbol{\psi}$) are initialised to 0.1. For the variational covariance matrix, all elements of $\mathbf{B}$ and the diagonal elements of $\mathbf{D}$ are initialised to 0.01. We use $p=4$ factors, as increasing the number of factors did not lead to improved results. Similar initial values are used for the HMC algorithm. Both the HVB-NoB and HMC algorithms are run for 10,000 iterations. Convergence diagnostics plots are provided in Section S9.1 of the online supplement.
Now, we briefly explain how the posterior sample of model parameters, $\boldsymbol{\xi}$, $\boldsymbol{\psi}$, and missing values $\boldsymbol{y}_{u}$ are generated in the HVB-NoB method. Once the algorithm converges, the set of variational parameters, $\boldsymbol{\lambda}_{s_{m}}=\left(\boldsymbol{\mu}_{s_{m}}^{\top},\text{vech}(\textbf{B}_{s_{m}})^{\top},\textbf{d}_{s_{m}}^{\top}\right)^{\top}$, is obtained. Given these variational parameters, $10,000$ draws from the variational distribution, $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\sim N\left((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top};\boldsymbol{\mu}_{s_{m}},\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2}\right)$, are generated, which form the posterior sample of $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$. Based on this posterior sample of model parameters, the posterior sample of $\boldsymbol{y}_{u}^{(i)}$ is generated from the conditional distribution $p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi}^{(i)},\boldsymbol{\psi}^{(i)})$, where $\boldsymbol{\xi}^{(i)}$ and $\boldsymbol{\psi}^{(i)}$ are the $i^{th}$ posterior sample for $i=1,\dots,10,000$.
Figure 2 compares the posterior densities of selected model parameters estimated using HVB-NoB and HMC for fitting the YJ-SEM-Gau model on the simulated dataset with missing data. For most model parameters, the posterior densities obtained from both methods are nearly identical. However, the posterior density of $\gamma$ estimated using the HVB-NoB method is slightly overestimated compared to that of the HMC method.
Figure 2: Comparison of posterior densities of selected parameters in the YJ-SEM-Gau model, obtained using the HMC and HVB-NoB methods on the simulated dataset with missing values. The vertical line represents the true parameter value.
Notably, the computation time per iteration of the VB and HVB-NoB algorithms is significantly lower than that of the HMC algorithm. The computation time per iteration for the VB algorithm is approximately 0.0361, while for the HMC algorithm, it is 0.4680 when estimating the YJ-SEM-Gau model without missing data. Similarly, the computation time per iteration for the HVB-NoB algorithm is approximately 0.0542 seconds, while for the HMC algorithm, it is 41.1658 seconds for estimating the YJ-SEM-Gau model with missing data. The high computational cost of the HMC algorithm in estimating SEMs (such as SEM-Gau) with missing data is also highlighted in the findings of Wijayawardhana et al., (2025).
4.2 Evaluating the performance of proposed SEMs
This section focuses on evaluating the robustness of the proposed SEMs and the VB methods using moderately large simulated datasets. We use a spatial weight matrix from the spData R package (Bivand et al.,, 2023), based on $25,357$ houses sold between 1993 and 1998 in Lucas County, Ohio, USA. Further details on this dataset are provided in Section 5.
We simulate two datasets from the YJ-SEM-t model using the spatial weight matrix $\textbf{W}_{1998}$, which corresponds to 4,378 houses sold in 1998 in Lucas County, Ohio, USA. In the first dataset, we set the parameters $\gamma=0.5$ (right skew) and $\nu=4$ (heavy tails). In the second dataset, we set the parameters $\gamma=1$ (no skew) and $\nu=30$ (light tails). The other parameters are set as follows: 6 fixed effects ($\boldsymbol{\beta}$) are randomly drawn from the discrete uniform values between -3 and 3 (excluding 0), $\sigma^{2}_{\textbf{e}}=0.5$, $\rho=0.8$, and all covariates are generated from a standard normal distribution, $N(0,1)$. Figure 3 shows kernel density plots of the response variables for both datasets.
Figure 3: Kernel density plots of the simulated data: The left panel shows the kernel density of the response variable from YJ-SEM-t with $\nu=4$ and $\gamma=0.5$, displaying right skewness and heavy tails (referred to as dataset 1). The right panel shows the kernel density of the response variable from YJ-SEM-t with $\nu=30$ and $\gamma=1$, which appears nearly symmetric with light tails (referred to as dataset 2).
Section S6 of the online supplement fits the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models to simulated datasets 1 and 2 without missing data, and evaluates how effectively these models, along with their VB estimation methods, capture the characteristics of these datasets. Section 4.2.1 introduces missing values into both datasets and evaluates the performance of the models and estimation method (the HVB method) in the presence of missing values.
4.2.1 SEMs with missing data
This section evaluates the accuracy of the HVB method for estimating the SEMs with missing values in the response variable using simulated data. For each simulated dataset, missing values in the response variable y are introduced according to the MNAR mechanism using a logistic regression model defined in Section 2.3. To construct this model, an additional covariate $\mathbf{x}^{*}$ is generated from a log-normal distribution with with a log-scale mean of 0 and a log-scale standard deviation of 1, and the model is regressed on $\mathbf{x}^{*}$ and the response variable y. We set $\psi_{0}=1$ (intercept), $\psi_{\textbf{x}^{*}}=-1$ (the coefficient for $\textbf{x}^{\ast}$), and $\psi_{\textbf{y}}=-0.1$ (the coefficient for y). This configuration results in approximately $40\%$ missing values, leading to $n_{u}=1,752$.
We fit SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models for the simulated datasets with missing values. As $n$ and $n_{u}$ are large, the HVB-AllB algorithm (Algorithm S2) described in Section S4.2 of the online supplement is used. The tuning parameters for the HVB-AllB algorithm is set as follows: the number of MCMC iterations ($N_{1}$) is set to 10, and the block size ($k^{*}$) is defined as $10\%$ of $n_{u}$, resulting in 10 or 11 blocks.
The initial values for the variational parameters corresponding to $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, $\rho$, and $\gamma$ are set as described in Section 4.1, while $\nu$ is initialised with a value of 4. The initial values for the missing data ($\mathbf{y}_{u}$) are simulated from the conditional distribution $p(\mathbf{y}_{u}\mid\boldsymbol{\xi}^{(0)},\mathbf{y}_{o})$, where $\boldsymbol{\xi}^{(0)}$ represents the vector of initial parameter values $\boldsymbol{\xi}$.
We ran the HVB-AllB algorithm for 10,000 iterations for the SEM-Gau and YJ-SEM-Gau models across both simulated datasets. For the SEM-t and YJ-SEM-t models, more iterations were required to achieve convergence due to the additional complexity introduced by estimating the latent variables $\boldsymbol{\tau}$ and the degrees of freedom parameter $\nu$. Specifically, for simulated dataset 1, SEM-t was run for 20,000 iterations and YJ-SEM-t for 10,000 iterations. For simulated dataset 2, SEM-t required 30,000 iterations, while YJ-SEM-t was run for 20,000 iterations.
Table 4 presents the posterior means and 95% credible intervals for some of the model parameters of the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models obtained using the HVB-AllB method applied to the simulated dataset 1 with missing values, along with the computational cost for one VB iteration. The table also includes the $\text{DIC}_{5}$ values for each model, calculated using the formula in Equation (20). Corresponding posterior density plots of the parameters are given in Section S7.2.2 of the online supplement.
We start by comparing the estimated values of the fixed effects (specifically $\beta_{0}$ and ${\beta}_{1}$ for the SEM and $\psi_{0}$ and ${\psi}_{1}$ for the missing data model), the variance parameter $\sigma_{\textbf{e}}^{2}$, and the spatial autocorrelation parameter $\rho$, as these parameters are common to all four models. The posterior means for $\beta_{0}$, $\beta_{1}$, $\psi_{0}$, $\psi_{1}$, and $\rho$ are nearly identical for both YJ-SEM-Gau and YJ-SEM-t, closely matching the true values. In contrast, the estimates from SEM-Gau and SEM-t show significant deviations from the true values. The estimated posterior means of $\sigma_{\textbf{e}}^{2}$ from SEM-Gau, SEM-t, and YJ-SEM-Gau differ considerably from the true value, while the posterior mean from YJ-SEM-t is much closer to the true value. The models incorporating the YJ transformation (YJ-SEM-Gau and YJ-SEM-t) successfully recover the parameter $\gamma$. The estimates of $\nu$ obtained from the SEM-t and YJ-SEM-t models are slightly inaccurate. Furthermore, the posterior means of the estimated missing values from YJ-SEM-t and YJ-SEM-Gau align more closely with the true missing values compared to those from SEM-Gau and SEM-t; see Figure S7 in Section S7.2.2 of the online supplement.
For the simulated dataset 1 with missing values in the response variable, the YJ-SEM-t model provides the best fit, as indicated by the lowest $\text{DIC}_{5}$ value, among all four models. This result is expected, as the dataset was generated from the YJ-SEM-t model, which incorporates both heavy tails and skewness. The second-best performer is YJ-SEM-Gau, which has the next lowest DIC value. Furthermore, similar to the case without missing values discussed in Section S6 of the online supplement, the relatively small difference in the $\text{DIC}_{5}$ values between YJ-SEM-t and YJ-SEM-Gau suggests that YJ-SEM-Gau achieves a model fit nearly comparable to that of YJ-SEM-t for this dataset.
The left panel of Figure 4 shows the kernel densities of the true missing values ($\mathbf{y}_{u}$) and the kernel density of posterior means of the missing values at locations $s_{1},s_{2},\ldots,s_{n}$, estimated from different SEMs for the simulated dataset 1. The densities from both YJ-SEM-Gau and YJ-SEM-t are nearly identical and closely align with the density of the true missing values.
The right panel of Figure 4 presents the posterior density of the maximum missing value ($\textbf{y}_{u_{\textrm{max}}}$) predicted by each SEM. The true value of the maximum missing value lies closer to the posterior densities produced by YJ-SEM-Gau and YJ-SEM-t than to those from SEM-Gau and SEM-t. These outcomes are as expected, since YJ-SEM-Gau and YJ-SEM-t yield similar parameter estimates and DIC values (see Table 4). Figure S8 in Section S7.2.2 of the online supplement provides similar density plots for the simulated dataset 2.
Figure 4: Left panel: The kernel density of the true missing values ($\textbf{y}_{u}$) and the kernel density of posterior means of the missing values ($\textbf{y}_{u}$) obtained using the HVB-AllB method for different SEMs for the simulated dataset 1 with missing values. Right panel: The posterior density of the maximum missing value ($\textbf{y}_{u_{\textrm{max}}}$) obtained by different SEMs for the simulated dataset 1 with missing values. The true value of the maximum missing value is indicated by the vertical line.
SEM-Gau SEM-t YJ-SEM-Gau YJ-SEM-t
${\beta}_{0}=-1$ 1.9268 (1.6342, 2.2270) 1.2929 (1.1456, 1.4382) -0.9862 (-1.0256, -0.9462) -0.9680 (-1.0238, -0.9123)
${\beta}_{1}=1$ 1.8848 (1.6344, 2.1381) 1.3230 (1.1675, 1.4797) 0.9938 (0.9554, 1.0332) 0.9835 (0.9338, 1.0324)
$\sigma^{2}_{\textbf{e}}=0.5$ 77.4912 (74.2316, 80.7856) 24.4549 (23.2015, 25.7503) 0.9994 (0.9331, 1.0701) 0.7028 (0.6615, 0.7451)
$\rho=0.8$ 0.2839 (0.2527, 0.3144) 0.1948 (0.1722, 0.2177) 0.8007 (0.7858, 0.8149) 0.8044 (0.7922, 0.8159)
$\nu=4$ NA 3.3400 (3.2319, 3.4772) NA 9.5082 (9.0327, 10.0109)
$\gamma=0.5$ NA NA 0.4965 (0.4822, 0.5113) 0.4990 (0.4882, 0.5099)
$\psi_{0}=1$ 0.8251 (0.7048, 0.9441) 0.9171 (0.8126, 1.0203) 0.9337 (0.8193, 1.0442) 0.9515 (0.8525, 1.0522)
${\psi}_{\textbf{y}}=-0.1$0.0793 (-0.0907, -0.0679) -0.0997 (-0.1178, -0.0812) -0.0794 (-0.0924, -0.0663) -0.0790 (-0.0891, -0.0686)
$\text{DIC}_{5}$ 36009.5 36302.7 18278.55 17855.59
CT 0.5894 1.2250 0.6000 1.339
Table 4: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 1 with missing data, along with the computation time (CT) in seconds for one HVB iteration. The table also presents the $\text{DIC}_{5}$ values for each model. Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
Table 5 reports the posterior means and $95\%$ credible intervals for key parameters of the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models, obtained using the HVB-AllB method on the simulated dataset 2 with missing values, along with the computational cost for one VB iteration. The table also presents the corresponding $\text{DIC}_{5}$ values for each model. Posterior density plots for these parameters are provided in Section S7.2.2 of the online supplement. The estimated posterior means for $\beta_{0}$, $\beta_{1}$, $\psi_{0}$, and $\sigma^{2}_{\textbf{e}}$ across all four models closely match their true values. Similarly, the estimated posterior mean of $\gamma$ for the YJ-SEM-Gau and YJ-SEM-t models is also close to the true value. The parameter $\nu$ is accurately estimated in both the SEM-t and YJ-SEM-t models. In addition, all four models produce posterior means for the estimated missing values that closely match the true values; see Figure S8 in Section S7.2.2 of the online supplement.
SEM-Gau SEM-t YJ-SEM-Gau YJ-SEM-t
${\beta}_{0}=-3$3.0131 (-3.0495, -2.9764) -3.0150 (-3.0497, -2.9792) -2.9511 (-2.9820, -2.9199) -2.9467 (-2.9767, -2.9161)
${\beta}_{1}=2$ 1.9942 (1.9582, 2.0309) 1.9884 (1.9605, 2.0165) 1.9803 (1.9621, 1.9979) 1.9834 (1.9568, 2.0096)
$\sigma^{2}_{\textbf{e}}=0.5$ 0.5528 (0.5222, 0.5849) 0.5143 (0.4892, 0.5406) 0.5450 (0.5155, 0.5759) 0.5096 (0.4826, 0.5377 )
$\rho=0.8$ 0.8002 (0.7857, 0.8141) 0.7990 (0.7874, 0.8101) 0.8003 (0.7892, 0.8113) 0.7971 (0.7841, 0.8095)
$\nu=30$ NA 30.2880 (29.1096, 31.4718) NA 29.7195 (27.9820, 31.6032)
$\gamma=1$ NA NA 1.0032 (0.9933, 1.0129) 1.0017 (0.9961, 1.0074)
$\psi_{0}=0.5$ 0.4949 (0.3847, 0.6044) 0.5036 (0.4219, 0.5841) 0.5477 (0.4319, 0.6619) 0.5467 (0.4742, 0.6204)
${\psi}_{\textbf{y}}=-0.1$0.0956 (-0.1210, -0.0706) -0.1090 (-0.1320, -0.0864) -0.0965 (-0.1106, -0.0827) -0.0972 (-0.1137, -0.0809)
$\text{DIC}_{5}$ 14713.89 14221.78 14695.32 14164.86
CT 0.5755 1.2166 0.6048 1.3486
Table 5: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 2 with missing data, along with the computation time (CT) in seconds for one HVB iteration. The table also presents the $\text{DIC}_{5}$ values for each model. Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
Based on the $\text{DIC}_{5}$ values, the YJ-SEM-t model offers the best fit for the simulated dataset 2 with missing values. In contrast, for the simulated dataset 2 without missing values, the SEM-Gau model provides the best fit; see Section S6 of the online supplement. For the simulated dataset 2 with missing data, the SEM-t model has the second-lowest $\text{DIC}_{5}$, indicating it is the second-best performing model.
As shown in Tables 4 and 5, the HVB-AllB algorithms for SEM-t and YJ-SEM-t, which incorporate Student’s $t$ errors, require significantly more computation time than those for SEM-Gau and YJ-SEM-Gau. This increased computational cost arises from the greater complexity of VB (HVB-AllB) optimisation in models with $t$ errors. Specifically, these models introduce an additional length $n$ latent variable vector and include the degrees of freedom parameter $\nu$, both of which increase the dimensionality and complexity of the VB optimisation.
Our simulations demonstrate that the standard Gaussian VB method for full data and the hybrid VB (HVB-NoB and HVB-AllB) methods for missing data offer accurate yet computationally efficient approximations for estimating both Gaussian and non-Gaussian SEMs, outperforming HMC in terms of computational cost. The YJ-SEM-Gau and YJ-SEM-t models are particularly effective for modeling data with pronounced skewness and heavy tails. In most scenarios, the DIC values show that YJ-SEM-Gau performs nearly as well as YJ-SEM-t. Although estimating YJ-SEM-t incurs significantly higher computational cost compared to SEM-Gau (the standard SEM with Gaussian errors), YJ-SEM-Gau has a similar computational cost to SEM-Gau while offering substantial improvements in modeling non-Gaussian data.
5 Real data application
We apply the proposed SEMs and VB methods to analyse a dataset of 25,357 single-family homes sold between 1993 and 1998, along with a spatial weight matrix of size $357\times 25,$ 357 × 25 , 357 from Lucas County, Ohio, USA. This dataset is available in the R package spData (Bivand et al.,, 2023). The data set includes various house characteristics: house age, the lot size in square feet, the number of rooms, the total living area in square feet, the number of bedrooms, and binary indicators for each year from 1993 to 1998 to represent the year of house sale. For our analysis, we focus on the house price data from 1998, which contains $4,378$ observations. We call this dataset Lucas-1998-HP. House prices are typically positive and often exhibit skewness. To transform house prices into the entire real line and reduce skewness, researchers commonly apply logarithmic transformations before modeling (LeSage and Pace,, 2004; Wijayawardhana et al.,, 2024). Accordingly, we apply the natural logarithmic transformation to the house prices in the Lucas-1998-HP dataset. Despite this transformation, the resulting variable still exhibits left skewness (see Figure S9 in Section S8 of the online supplement).
To evaluate the impact of missing data on Gaussian and non-Gaussian SEMs, we use the Lucas-1998-HP dataset, a complete real-world dataset. We first fit the SEMs to the full dataset using the VB algorithm. Then, we intentionally introduce missing values into the response variable and refit the SEMs using the proposed HVB-AllB method. This setup allows for a direct comparison of parameter estimates obtained with and without missing data.
For the Lucas-1998-HP dataset, we fit all four models described in Section 2. The set of independent variables includes various powers of house age (age, $\text{age}^{2}$, and $\text{age}^{3}$), the natural logarithm of the lot size in square feet ($\ln(\text{lotsize})$), the number of rooms (rooms), the natural logarithm of the total living area in square feet ($\ln(\text{LTA})$), and the number of bedrooms. The natural logarithm of house price in hundreds of thousands is selected as the dependent variable. Results for the complete dataset (without missing values) are presented in Section 5.1, while Section 5.2 summarises the findings with missing data.
5.1 Full Data SEMs
This section presents the results of analysing the Lucas-1998-HP dataset by fitting all four SEMs introduced in Section 2, using the VB algorithms outlined in Section S2 of the online supplement. The starting values for the VB algorithms are obtained as follows: The variational mean vector ($\boldsymbol{\mu}$) is set to the estimates for $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ obtained from fitting the SEM-Gau model via maximum likelihood estimation (MLE). The initial value for $\nu$ is set to 4, and the initial value for $\gamma$ is set to 1. For the variational covariance matrix parameters, the elements of B and the diagonal elements of D are all initialised to 0.01. We use $p=4$ factors.
As in the simulation study with full data described in Section S6 of the online supplement, models with Student-$t$ errors required more iterations to achieve convergence. The VB algorithms were run for 20,000 iterations for the SEM-Gau and YJ-SEM-Gau models and 75,000 iterations for the SEM-t and YJ-SEM-t models, at which point convergence was attained. Convergence was assessed visually by examining the plots of variational means over iterations, which are provided in Section S9.2 of the online supplement.
The estimated posterior means and $95\%$ credible intervals for selected model parameters across different models are summarised in Table 6, along with the corresponding DIC values. According to the $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values, the YJ-SEM-t model is the most appropriate SEM for the Lucas-1998-HP dataset, followed by the YJ-SEM-Gau model as the second most suitable.
SEM-Gau SEM-t YJ-SEM-Gau YJ-SEM-t
$intercept$0.4336 (-0.4469, -0.4207) -0.3865 (-0.3979, -0.3752) -0.3020 (-0.3278, -0.2765) -0.2849 (-0.3087, -0.2612)
$\beta_{\text{age}}$ 0.1650 (0.0557, 0.2742) -0.4197 (-0.5236, -0.3162) -0.1643 (-0.2926, -0.0321) -0.3201 (-0.3984, -0.2405)
$\beta_{\text{rooms}}$ 0.0068 (-0.0176, 0.0313) 0.0188 (-0.0018, 0.0395) 0.0311 (-0.0281, 0.0893) 0.0238 (-0.0056, 0.0524)
$\beta_{\text{log(lotsize)}}$ 0.1614 (0.1480, 0.1752) 0.1193 (0.1033, 0.1350) 0.1441 (0.1068, 0.1815) 0.1306 (0.1192, 0.1420)
$\sigma^{2}_{\textbf{e}}$ 0.1577 (0.1516, 0.1641) 0.0670 (0.0633, 0.0708) 0.1031 (0.0967, 0.1099) 0.0695 (0.0647, 0.0744)
$\rho$ 0.6291 (0.5605, 0.6902) 0.5475 (0.4900, 0.6015) 0.6019 (0.5236, 0.6712) 0.5990 (0.5261, 0.6651)
$\nu$ NA 3.0249 (3.0000, 3.1324) NA 7.9482 (7.5327, 8.3940)
$\gamma$ NA NA 1.5264 (1.4783, 1.5728) 1.5529 (1.5145, 1.5904)
$\text{DIC}_{1}$ 4470.316 4878.664 4329.431 3720.687
$\text{DIC}_{2}$ 4469.655 5031.204 4332.07 3766.714
Table 6: Posterior means and 95% credible intervals for selected parameters of different SEMs for the Lucas-1998-HP dataset (no missing values). The table also includes DIC values for each model. Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
We now compare estimates of key parameters across different SEMs. A crucial aspect of SEM estimation is identifying the underlying spatial correlation, represented by the parameter $\rho$. For the SEM-Gau, YJ-SEM-Gau, and YJ-SEM-t models, the estimated values of $\rho$ are relatively close, with posterior mean values of 0.6291, 0.6019, and 0.5990, respectively. This suggests a moderate spatial correlation in the Lucas-1998-HP dataset. However, for the SEM-t model, the estimated $\rho$ is noticeably lower, with a posterior mean value of 0.5475, compared to the other three models.
Similar to linear regression models, the fixed effects $\boldsymbol{\beta}$ in SEMs represent the influence of covariates on the response variable, which in this case is the logarithm of the house price. An interesting disparity arises in the estimated posterior mean of the fixed effect corresponding to the covariate ’age’ ($\beta_{\text{age}}$). For the SEM-Gau model, $\beta_{\text{age}}$ is positive, while for the other three models, $\beta_{\text{age}}$ is negative. This suggests that for the SEM-Gau model, the age of the house is associated with an increase in house prices, while for the other models, the age of the house is linked to a decrease in house prices. The estimated posterior means of the fixed effect for the number of rooms ($\beta_{\text{rooms}}$) are 0.0068, 0.0188, 0.0311, and 0.0238 for the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models, respectively. These estimates suggest that the effect of the number of rooms on house prices varies across models. Notably, the YJ-SEM-Gau model shows the strongest positive influence, while the SEM-Gau model exhibits the weakest.
Additionally, the SEM-t, YJ-SEM-Gau, and YJ-SEM-t models provide further insights into the characteristics of the dataset. The posterior mean of $\nu=3.0249$ for SEM-t and $\nu=7.9482$ for YJ-SEM-t suggest that the log-transformed house price distribution exhibits heavy tails. Furthermore, the posterior means of $\gamma$, $1.5264$ for YJ-SEM-Gau and $1.5529$ for YJ-SEM-t, suggest left skewness in the response variable, effectively capturing the asymmetry in the distribution of log-transformed house prices. These distributional features are visually evident in the kernel density plot of the response variable (see Figure S9 in Section S8 of the online supplement).
5.2 SEMs with missing data
This section discusses the analysis of the Lucas-1998-HP dataset with missing data in the response variable. We use the age of the house and the response variable of SEMs (logarithm of the house price) as the covariates in the missing data model. The missing data model has three parameters, $\boldsymbol{\psi}=(\psi_{0},\psi_{\textbf{x}^{*}},\psi_{\textbf{y}})$. We set the values $\psi_{0}=0.1$, $\psi_{\textbf{x}^{*}}=0.1$, and $\psi_{\textbf{y}}=-0.1$. With these settings, approximately 50% of the responses are missing, resulting in $n_{u}=2,348$.
Since both $n$ and $n_{u}$ are large, we utilise the HVB-AllB algorithm. We set the number of MCMC iterations $N_{1}$ to 10 and the block size to 10% of $n_{u}$, resulting in 11 blocks. The starting values for the HVB-AllB algorithms are obtained as follows: the variational mean vector ($\boldsymbol{\mu}$) is set to the estimates of $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ obtained from fitting the SEM-Gau model with only the observed data, $\textbf{y}_{o}$, using MLE. The initial value for $\nu$ is set to 4, and the initial value for $\gamma$ is set to 1. For the parameters of the missing data generating model ($\psi_{0}$, $\psi_{\textbf{x}^{*}}$, and $\psi_{\textbf{y}}$), all coefficients are set to an initial value of 0.01. For the variational covariance matrix parameters, the elements of B and the diagonal elements of D are all initialised to 0.01. The initial values for the missing data ($\mathbf{y}_{u}$) are simulated from the conditional distribution $p(\mathbf{y}_{u}\mid\boldsymbol{\xi}^{(0)},\mathbf{y}_{o})$, where $\boldsymbol{\xi}^{(0)}$ represents the vector of initial parameter values for $\boldsymbol{\xi}$. We use $p=4$ factors.
Similar to the simulation study with missing data, models with Student’s $t$ errors required a relatively large number of iterations to converge. Therefore, we ran the HVB-AllB algorithm for 10,000 iterations for the SEM-Gau and YJ-SEM-Gau models, 40,000 iterations for the SEM-t model, and 20,000 iterations for the YJ-SEM-t model. Convergence plots can be found in Section S9.2 of the online supplement.
The estimated posterior means and 95% credible intervals for selected model parameters across different models are summarised in Table 7, along with the corresponding $\text{DIC}_{5}$ values. According to the $\text{DIC}_{5}$ values, the YJ-SEM-t model is the most appropriate SEM for the Lucas-1998-HP dataset with missing data, followed by the YJ-SEM-Gau model as the second best model.
SEM-Gau SEM-t YJ-SEM-Gau YJ-SEM-t
$intercept$0.4189 (-0.4413, -0.3960) -0.3721 (-0.3886, -0.3562) -0.2856 (-0.3008, -0.2703) -0.2752 (-0.2956, -0.2544)
$\beta_{\text{age}}$ 0.1935 (0.0335, 0.3510) -0.4640 (-0.5990, -0.3324) -0.1473 (-0.2598, -0.0345) -0.3768 (-0.4870, -0.2687)
$\beta_{\text{rooms}}$ 0.0010 (-0.0409, 0.0432) 0.0179 (-0.0132, 0.0491) 0.0341 (0.0021, 0.0655) 0.0438 (0.0053, 0.0816)
$\beta_{\text{log(lotsize)}}$ 0.1360 (0.1097, 0.1620) 0.1124 (0.0952, 0.1293) 0.1173 (0.1020, 0.1323) 0.1219 (0.0786, 0.1651)
$\sigma^{2}_{\textbf{e}}$ 0.1463 (0.1378, 0.1553) 0.0696 (0.0649, 0.0746) 0.0894 (0.0847, 0.0943) 0.0657 (0.0617, 0.0699)
$\rho$ 0.5872 (0.4810, 0.6818) 0.5686 (0.4553, 0.6695) 0.5494 (0.4337, 0.6502) 0.5377 (0.4325 0.6322)
$\nu$ NA 4.1165 (3.9476, 4.3046) NA 10.6394 (10.2060, 11.1003)
$\gamma$ NA NA 1.5814 (1.5297, 1.6299) 1.5998 (1.5552, 1.6415)
$\psi_{0}=0.1$ 0.1128 (0.0193, 0.2064) 0.2075 (0.1361, 0.2806) 0.1110 (0.0349, 0.1865) 0.1317 (0.0595, 0.2033)
$\text{DIC}_{5}$ 8074.382 7734.356 7309.394 6657.172
Table 7: Posterior means and 95% credible intervals for selected parameters of different SEMs for the Lucas-1998-HP dataset with missing data. The table also includes $\text{DIC}_{5}$ values for each model. Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
The estimated parameters from different SEMs are now briefly discussed. For the spatial autocorrelation parameter, $\rho$, the posterior mean estimates across all SEMs are slightly lower than those obtained without missing values, shown in Table 6. However, these estimates still indicate a moderate spatial correlation in the Lucas-1998-HP dataset, consistent with the conclusions drawn from SEMs fitted without missing values in Section 5.1.
The posterior mean estimates for the fixed effects ($\boldsymbol{\beta}$) in the SEMs lead to the same conclusions as those drawn from the Lucas-1998-HP dataset without missing values. For example, the posterior mean estimate of $\beta_{\text{age}}$ in the SEM-Gau model suggests a positive effect of age on house prices, whereas the estimates from the other three models indicate a negative effect. Moreover, the estimated intercepts in the missing data model, ${\psi}_{0}$, are close to the true value of $0.1$ for all models except SEM-t.
The posterior mean estimates of $\nu$ for the SEM-t and YJ-SEM-t models indicate moderate to strong heavy-tailed behavior. Meanwhile, the estimated $\gamma$ values for the YJ-SEM-Gau and YJ-SEM-t models suggest left skewness in the data. These conclusions are consistent with the findings from the analysis without missing data discussed in Section 5.1.
6 Conclusion
Our article introduces three novel simultaneous autoregressive (SAR) models for non-Gaussian spatial data. Specifically, we extend the conventional spatial error model (SEM-Gau), a widely used SAR model, to accommodate skewed and heavy tailed features of real datasets. The novel SEMs are SEM-t, YJ-SEM-Gau, and YJ-SEM-t. The SEM-t assumes that the error terms follow a Student’s t distribution, allowing heavier tails than a Gaussian distribution. The YJ-SEM-Gau retains Gaussian errors but applies the Yeo-Johnson (YJ) transformation to handle asymmetric response variables. The YJ-SEM-t combines these approaches by assuming Student’s t errors while incorporating the YJ transformation to account for skewness in the response variable. These extensions improve the robustness of SAR models in analysing complex spatial datasets. Additionally, we develop efficient variational Bayes (VB) methods to estimate these models with and without missing data. For cases with missing data, our approach assumes a missing not at random (MNAR) mechanism. Standard VB methods are inadequate for accurately estimating these SEMs in the presence of missing data; therefore, we employ hybrid VB (HVB) algorithms to address this challenge. The posterior distributions obtained from the proposed VB and HVB approximation methods are compared with those of the exact Hamiltonian Monte Carlo (HMC) method.
The empirical results show that: (1) The VB and HVB methods yield posterior density estimates for most model parameters that are similar to those obtained using the HMC method when estimating SEMs with and without missing values; (2) For models with Student’s $t$ errors (SEM-t and YJ-SEM-t), the posterior approximation of the degrees of freedom parameter ($\nu$) from the VB and HVB methods exhibits slight deviations compared to the true value; (3) For both VB and HVB algorithms, models with Student’s t errors require more computation time per iteration than models with Gaussian errors (SEM-Gau and YJ-SEM-Gau); (4) In general, models with Gaussian errors tend to converge in fewer iterations compared to those with Student’s t errors; (5) YJ-SEM-t is able to model complex response variable with skewness and heavy tails and is the best model for the Lucas-1998-HP dataset, both with and without missing data.
S1 Yeo and Johnson transformation and its derivatives
This section provides the Yeo and Johnson (YJ) transformation (Yeo and Johnson,, 2000) and its derivatives, which are used to construct the YJ-SEM-Gau and YJ-SEM-t models discussed in Section 2.2 of the main paper.
Let the random variable $y_{i}$ be asymmetric. The Yeo–Johnson (YJ) transformation, denoted $t_{\gamma}(y_{i})$, is used to make $y_{i}$ more symmetric, and is defined by:
$y_{i}^{*}=t_{{\gamma}}(y_{i})=\begin{cases}\frac{(y_{i}+1)^{\gamma}-1}{\gamma}&\text{if }y_{i}\geq 0,\\ -\frac{(-y_{i}+1)^{2-\gamma}-1}{2-\gamma}&\text{if }y_{i}<0,\end{cases}$ (S1)
where $20<\gamma<$ 2.
The density function of SEMs with YJ transformations includes the derivative of the YJ transformation with respect to $y_{i}$, as shown in Equation (7) of the main paper. This derivative is equal to
$\frac{dt_{{\gamma}}(y_{i})}{dy_{i}}=\begin{cases}{(y_{i}+1)^{\gamma-1}}&\text{if }y_{i}\geq 0,\\ (-y_{i}+1)^{1-\gamma}&\text{if }y_{i}<0.\end{cases}$ (S2)
S2 Variational Bayes approximation for SEMs with full data
Table S1: The definitions of r, M, and the parameters $\boldsymbol{\phi}$ for the four spatial error models. Each SEM has a mean vector of $\mathbf{X}\boldsymbol{\beta}$ and a covariance matrix of the form $\boldsymbol{\Sigma}=\sigma^{2}_{\textbf{e}}\textbf{M}^{-1}$, with $\textbf{A}=\textbf{I}_{n}-\rho\textbf{W}$.
Model $\boldsymbol{\phi}$ r M
SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $^{\top}\textbf{A}$ A
SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$ A
YJ-SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e},},\rho,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $^{\top}\textbf{A}$ A
YJ-SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$ A
Consider Bayesian inference for the parameter vector of SEMs, $\boldsymbol{\xi}$. For SEM-Gau and YJ-SEM-Gau, we have $\boldsymbol{\xi}=\boldsymbol{\phi}$, while for SEM-t and YJ-SEM-t, $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$. See Table S1 for the definition of $\boldsymbol{\phi}$ for each SEM. Let the prior distribution of $\boldsymbol{\xi}$ be denoted by $p(\boldsymbol{\xi})$. Then, the posterior distribution of $\boldsymbol{\xi}$ given the data y is denoted by $p(\boldsymbol{\xi}\mid\textbf{y})$ and is given by:
$p(\boldsymbol{\xi}\mid\textbf{y})\propto p(\textbf{y}\mid\boldsymbol{\xi})p(\boldsymbol{\xi}),$ (S3)
where the term $p(\textbf{y}\mid\boldsymbol{\xi})$ represents the likelihood of y, and the log-likelihood of y for different SEMs are given in Section 2 of the main paper. We define $p(\textbf{y}\mid\boldsymbol{\xi})p(\boldsymbol{\xi})=h(\boldsymbol{\xi})$. The terms $h(\boldsymbol{\xi})$ for each SEM, along with the total length of the parameter vector to be estimated in each model (i.e., the length of the vector $\boldsymbol{\xi}$), are provided in Table S2.
Table S2: The term $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\xi})p(\boldsymbol{\xi})$, and the total number of model parameters $s$ (including the length of the latent vector $\boldsymbol{\tau}$ if exists) for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models. The distribution $p(\boldsymbol{\tau}\mid\boldsymbol{\phi})=\prod_{i=1}^{n}p(\tau_{i}\mid\nu)$ represents the $n$ independent inverse gamma latent variables used in SEMs with Student-$t$ errors (see Equation (2) on the main paper).
Model $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\xi})p(\boldsymbol{\xi})$ $s$
SEM-Gau $p(\textbf{y}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$ $r+3$
SEM-t $p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$ $r+4+n$
YJ-SEM-Gau $p(\textbf{y}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$ $r+4$
YJ-SEM-t $p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$ $r+5+n$
Now we explain how variational Bayes (VB) inference is used to approximate the posterior distribution in Equation (S3). We use the variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$, indexed by the variational parameter $\boldsymbol{\lambda}$ to approximate the posterior $p(\boldsymbol{\xi}\mid\textbf{y})$. The VB approach approximates this posterior distribution by minimising the Kullback-Leibler (KL) divergence between $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$and $p(\boldsymbol{\xi}\mid\textbf{y})$. The KL divergence between these two distributions is
[…]
where $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\xi})p(\boldsymbol{\xi})$. Table S2 presents the expressions for $h(\boldsymbol{\xi})$ for each of the SEMs. The ELBO in Equation (S5) can be written as an expectation with respect to $q_{\boldsymbol{\lambda}}$,
$\mathcal{L}(\boldsymbol{\lambda})=E_{q}\left[\text{log}~{}h(\boldsymbol{\xi})-\text{log}~{}q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\right],$ (S6)
where $E_{q}\left[\cdot\right]$ denotes the expectation with respect to $q_{\boldsymbol{\lambda}}$.
We employ the Gaussian variational approximation, where we select $q_{\boldsymbol{\lambda}}$ to be a multivariate normal distribution. As a result, the variational parameters $\boldsymbol{\lambda}$ consist of both the mean vector and the distinct elements of the covariance matrix. We impose a factorised covariance structure (Ong et al.,, 2018) on the covariance matrix of $q_{\boldsymbol{\lambda}}$, which reduces the number of distinct elements in the covariance matrix. Under a factor covariance structure for the covariance matrix, the variational distribution is parameterised as $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\sim N(\boldsymbol{\xi};\boldsymbol{\mu},\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})$, where $\boldsymbol{\mu}$ is the $s\times 1$ mean vector, B is an $s\times p$ full rank matrix with $p<<s$, and D is an $s\times s$ diagonal matrix having diagonal elements $\textbf{d}=(d_{1},\ldots,d_{s})$. For all SEMs, the total number of parameters, $s$ is provided in Table S2. We further impose the restriction that the upper triangular elements of B are all zero.
Since the ELBO in Equation (S6) does not have a closed-form solution for all four models, we use stochastic gradient ascent (SGA) methods (Nott et al.,,2012;Titsias and Lázaro-Gredilla,,2014,2015) to maximise the ELBO with respect to the variational parameters, $\boldsymbol{\lambda}$. Let $\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$ be the gradient of the objective function to be optimised, $\mathcal{L}(\boldsymbol{\lambda})$, and $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})}$ an unbiased estimate of $\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$. The SGA method updates $\boldsymbol{\lambda}$ iteratively, starting from an initial value $\boldsymbol{\lambda}^{(0)}$, according to the following scheme:
[…]
Minimising the variance of the gradient estimator, $\widehat{{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})}}$, is crucial for both stability and fast convergence of the SGA algorithm, which we address using the reparameterisation trick (Kingma and Welling,, 2013; Rezende et al.,, 2014).
We now briefly explain how the reparameterisation trick, combined with a factor covariance structure for the Gaussian variational distribution, is used to obtain efficient gradient estimates in this article. More detailed information can be found in Ong et al., (2018).
To apply the reparameterisation trick, we begin by sampling from $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$. This involves first drawing $\boldsymbol{\zeta}=(\boldsymbol{\eta}^{\top},\boldsymbol{\epsilon}^{\top})^{\top}$, where $\boldsymbol{\eta}$ is a $p$-dimensional vector and $\boldsymbol{\epsilon}$ is an $s$-dimensional vector, from a fixed distribution $f_{\boldsymbol{\zeta}}(\boldsymbol{\zeta})$ that is independent of the variational parameters. Next, we compute $\boldsymbol{\xi}=u(\boldsymbol{\zeta},\boldsymbol{\lambda})=\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon}$. We let, $f_{\boldsymbol{\zeta}}(\boldsymbol{\zeta})$ to follow a standard normal distribution. i.e. $\boldsymbol{\zeta}=(\boldsymbol{\eta}^{\top},\boldsymbol{\epsilon}^{\top})^{\top}~{}\sim N(\textbf{0},\textbf{I}_{s+p})$, where 0 is the zero mean vector of size $s+p$ and $\textbf{I}_{s+p}$ is the identity matrix of the same size. Then, the expectation in Equation (S6) is expressed with respect to the distribution $f_{\boldsymbol{\zeta}}$ as
$\begin{split}\mathcal{L}(\boldsymbol{\lambda})&=E_{q}\left[\text{log}~{}h(\boldsymbol{\xi})-\text{log}~{}q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\right]\\ &=E_{f_{\boldsymbol{\zeta}}}\left[\text{log}~{}h(u(\boldsymbol{\zeta,\boldsymbol{\lambda}}))-\text{log}~{}q_{\boldsymbol{\lambda}}(u(\boldsymbol{\zeta,\boldsymbol{\lambda}}))\right],\end{split}$ (S8)
and differentiating $\mathcal{L}(\boldsymbol{\lambda})$ under the integral sign, we obtain
$\begin{split}\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\zeta}}}\left[\nabla_{\boldsymbol{\lambda}}~{}\text{log}~{}h(u(\boldsymbol{\zeta,\boldsymbol{\lambda}}))-\nabla_{\boldsymbol{\lambda}}~{}\text{log}~{}q_{\boldsymbol{\lambda}}(u(\boldsymbol{\zeta,\boldsymbol{\lambda}}))\right],\\ &=E_{f_{\boldsymbol{\zeta}}}\left[\frac{du(\boldsymbol{\zeta,\boldsymbol{\lambda}})^{\top}}{d\boldsymbol{\lambda}}\{\nabla_{\boldsymbol{\xi}}\text{log}~{}h(\boldsymbol{\xi})-\nabla_{\boldsymbol{\xi}}\text{log}~{}q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\}\right],\\ \end{split}$ (S9)
where $\frac{du(\boldsymbol{\zeta,\boldsymbol{\lambda}})}{d\boldsymbol{\lambda}}$ is the derivative of the transformation $u(\boldsymbol{\zeta,\boldsymbol{\lambda}})=\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon}$ with respect to the variational parameters $\boldsymbol{\lambda}=(\boldsymbol{\mu}^{\top},\text{vech}(\textbf{B})^{\top},\textbf{d}^{\top})^{\top}$, where the ”vech” operator vectorises a matrix by stacking its columns from left to right while removing all the elements above the diagonal (the superdiagonal elements) of the matrix. The gradients $\frac{du(\boldsymbol{\zeta},\boldsymbol{\lambda})}{d\boldsymbol{\lambda}}$ and $\nabla_{\boldsymbol{\xi}}\log q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$ are provided in Section S2.1. The expressions for $\nabla_{\boldsymbol{\xi}}\log h(\boldsymbol{\xi})$ differ for each SEM and can be found in Section S2.2.
Algorithm S1 outlines the VB algorithm for estimating SEMs with full data. The unbiased estimate of the gradient, $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})}$, presented in step 4 of Algorithm S1 can be constructed using the reparameterisation trick described earlier, with a single sample drawn from $f_{\boldsymbol{\zeta}}(\boldsymbol{\zeta})$.
Algorithm S1 Variational Bayes algorithm
1:  Initialise $\boldsymbol{\lambda}^{(0)}=(\boldsymbol{\mu}^{\top(0)},\textrm{vech}{(\textbf{B})}^{\top(0)},\textbf{d}^{\top(0)})^{\top}$ and set $t=0$
2:  repeat
3:     Generate $(\boldsymbol{\eta}^{(t)},\boldsymbol{\epsilon}^{(t)})\sim N(\textbf{0},\textbf{I}_{s+p})$
4:     Construct unbiased estimates $\widehat{{\nabla_{\boldsymbol{\mu}}\mathcal{L}(\boldsymbol{\lambda})}},\widehat{{\nabla_{\textrm{vech}(\textbf{B})}\mathcal{L}(\boldsymbol{\lambda})}},\text{and}~{}\widehat{{\nabla_{\textbf{d}}\mathcal{L}(\boldsymbol{\lambda})}}$ using Equations  (S11), (S12) and  (S13) in Section S2.1 at $\boldsymbol{\lambda}^{(t)}$.
[…]
9:     Set $\boldsymbol{\lambda}^{(t+1)}=(\boldsymbol{\mu}^{\top(t+1)},\textrm{vech}(\textbf{B})^{\top(t+1)},\textbf{d}^{\top(t+1)})$, and $t=t+1$
10:  until some stopping rule is satisfied
S2.1 Expressions for $\frac{du(\boldsymbol{\zeta},\boldsymbol{\lambda})}{d\boldsymbol{\lambda}}$
In the reparameterisation gradient given in Equation (S9), $\frac{du(\boldsymbol{\zeta,\boldsymbol{\lambda}})}{d\boldsymbol{\lambda}}$ is the derivative of the transformation $u(\boldsymbol{\zeta,\boldsymbol{\lambda}})=\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon}$ with respect to the variational parameters $\boldsymbol{\lambda}=(\boldsymbol{\mu}^{\top},\text{vech}(\textbf{B})^{\top},\textbf{d}^{\top})^{\top}$. We write that $u(\boldsymbol{\zeta,\boldsymbol{\lambda}})=\boldsymbol{\mu}+(\boldsymbol{\eta}^{\top}\otimes\textbf{I}_{s})\text{vech}(\textbf{B})+\textbf{d}\circ\boldsymbol{\epsilon}$, where $\otimes$ represents the Kronecker product, and $\textbf{I}_{s}$ is the identity matrix of size $s$. It can be shown that $\nabla_{\boldsymbol{\xi}}\text{log}~{}q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})=-(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\boldsymbol{\xi}-\boldsymbol{\mu})$,
$\frac{du(\boldsymbol{\zeta},\boldsymbol{\lambda})}{d\boldsymbol{\mu}}=\textbf{I}_{s}\\ ~{}~{}~{}~{}~{}\text{and}~{}~{}~{}~{}~{}\frac{du(\boldsymbol{\zeta},\boldsymbol{\lambda})}{d\text{vech}(\textbf{B})}=\boldsymbol{\eta}^{\top}\otimes\textbf{I}_{s}.$ (S10)
The derivatives of the lower bound with respect to variational parameters are:
$\begin{split}\nabla_{\boldsymbol{\mu}}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\zeta}}}[\nabla_{\boldsymbol{\xi}}~{}\text{log}~{}h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})],\end{split}$ (S11)
$\begin{split}\nabla_{\text{vech}(\textbf{B})}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\zeta}}}[\nabla_{\boldsymbol{\xi}}\text{log}~{}h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\eta}^{\top}\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\eta}^{\top}],\end{split}$ (S12)
and
$\begin{split}\nabla_{\textbf{d}}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\zeta}}}[\text{diag}(\nabla_{\boldsymbol{\xi}}\text{log}~{}h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\epsilon}^{\top}\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\epsilon}^{\top})],\end{split}$ (S13)
where $\text{diag}(\cdot)$ is the vector of diagonal elements extracted from a square matrix. The analytical expressions for $\nabla_{\boldsymbol{\xi}}\text{log}~{}h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})=\nabla_{\boldsymbol{\xi}}\text{log}~{}h(\boldsymbol{\xi})$ in Equations (S11)-(S13) for different SEMs are provided in Section S2.2.
S2.2 Expressions for $\nabla_{\boldsymbol{\xi}}\text{log}~{}h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})=\nabla_{\boldsymbol{\xi}}\text{log}~{}h(\boldsymbol{\xi})$
The term $h(\boldsymbol{\xi})$ represents the likelihood multiplied by the prior distribution (see Section S2). For different SEMs, the expressions for $h(\boldsymbol{\xi})$ are given in Table S2. Table S3 shows the prior distributions used for the model parameters in this study. In this section, we compute the derivatives of $\log h(\boldsymbol{\xi})$ with respect to the model parameter vector $\boldsymbol{\xi}$, denoted as $\nabla_{\boldsymbol{\xi}}\log h(\boldsymbol{\xi})$. These derivatives are necessary for computing the reparameterisation gradient, as outlined in Equation (S9), for each SEM.
Parameter $\boldsymbol{\beta}$ $\omega^{\prime}$ $\rho^{\prime}$ $\nu^{\prime}$ $\gamma^{\prime}$ $\boldsymbol{\psi}$
Prior distribution $N(\textbf{0},\sigma^{2}_{\boldsymbol{\beta}}\textbf{I})$ $N(0,\sigma^{2}_{\omega^{\prime}})$ $N(0,\sigma^{2}_{\rho^{\prime}})$ $N(0,\sigma^{2}_{\nu^{\prime}})$ $N(0,\sigma^{2}_{\gamma^{\prime}})$ $N(\textbf{0},\sigma^{2}_{\boldsymbol{\psi}}\textbf{I})$
Hyperparameters $\sigma^{2}_{\boldsymbol{\beta}}=10^{2}$ $\sigma^{2}_{\omega^{\prime}}=10^{2}$ $\sigma^{2}_{\rho^{\prime}}=10^{2}$ $\sigma^{2}_{\nu^{\prime}}=10^{2}$ $\sigma^{2}_{\gamma^{\prime}}=10^{2}$ $\sigma^{2}_{\boldsymbol{\psi}}=10^{2}$
Table S3: Prior distributions of model parameters
S2.2.1 Derivatives for SEM-Gau
For SEM-Gau, $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}$. The logarithm of $h(\boldsymbol{\xi})$ is $\text{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})+\text{log}~{}p(\boldsymbol{\phi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, and $\rho$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3. This leads to
$\text{log}~{}h(\boldsymbol{\xi})\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}},$ (S14)
where the values of $\sigma^{2}_{\boldsymbol{\beta}}$, $\sigma^{2}_{\omega^{\prime}}$, and $\sigma^{2}_{\rho^{\prime}}$ are each set to 100, as detailed in Table S3, while r and M are provided in Table S1.
The derivative of $\text{log}~{}h(\boldsymbol{\xi})$ in Equation (S14) with respect to $\boldsymbol{\beta}$ is
$\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\boldsymbol{\beta}}=e^{-\omega^{\prime}}\textbf{r}^{\top}\textbf{M}\textbf{X}-\frac{\boldsymbol{\beta}^{\top}}{\sigma^{2}_{\boldsymbol{\beta}}},$ (S15)
the derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $\omega^{\prime}$ is
$\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\omega^{\prime}}=-\frac{n}{2}+\frac{e^{-\omega^{\prime}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}-\frac{\omega^{\prime}}{\sigma^{2}_{\omega^{\prime}}},$ (S16)
and the derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $\rho^{\prime}$ is
$\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\rho^{\prime}}=\left(\frac{1}{2}\text{tr}\left\{(\textbf{A}^{\top}\textbf{A})^{-1}\left(\frac{\partial\textbf{A}^{\top}\textbf{A}}{\partial\rho}\right)\right\}-\frac{e^{-\omega^{\prime}}}{2}\textbf{r}^{\top}\left(\frac{\partial\textbf{A}^{\top}\textbf{A}}{\partial\rho}\right)\textbf{r}\right)\left(\frac{\partial\rho}{\partial\rho^{\prime}}\right)-\frac{\rho^{\prime}}{\sigma^{2}_{\rho^{\prime}}},$ (S17)
where $\text{tr}\{\cdot\}$ denotes the trace operator, which computes the sum of the diagonal elements of a matrix,
$\begin{split}\left(\frac{\partial\textbf{A}^{\top}\textbf{A}}{\partial\rho}\right)&=-(\textbf{W}^{\top}+\textbf{W})+2\rho\textbf{W}^{\top}\textbf{W},\\ \left(\frac{\partial\rho}{\partial\rho^{\prime}}\right)&=\frac{2e^{\rho^{\prime}}}{(1+e^{\rho^{\prime}})^{2}}.\\ \end{split}$
S2.2.2 Gradients for SEM-t
For SEM-t, $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu)^{\top}$. The logarithm of $h(\boldsymbol{\xi})$ is $p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\text{log}~{}p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\text{log}~{}p(\boldsymbol{\phi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$ and $\nu$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3. This leads to
$\begin{split}\text{log}~{}h(\boldsymbol{\xi})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}\\ &+\sum_{i=1}^{n}\textrm{log}~{}p(\tau_{i}\mid\nu)+\textrm{log}\left|\frac{\partial\tau_{i}}{\partial\tau_{i}^{\prime}}\right|-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\nu^{\prime}}^{2}}{2\sigma^{2}_{\nu^{\prime}}},\end{split}$ (S18)
where the values of $\sigma^{2}_{\boldsymbol{\beta}}$, $\sigma^{2}_{\omega^{\prime}}$, $\sigma^{2}_{\rho^{\prime}}$, and $\sigma^{2}_{\nu^{\prime}}$ are each set to 100, as detailed in Table S3, while r and M are provided in Table S1. The term $\left|\frac{\partial\tau_{i}}{\partial\tau_{i}^{\prime}}\right|$ represents the derivative of the inverse transformation $\tau_{i}=e^{\tau_{i}^{\prime}}$, which maps the latent variable $\tau_{i}$, for $i=1,\ldots,n$, onto the real line using the transformation $\tau_{i}^{\prime}=\log(\tau_{i})$.
The derivative of $\text{log}~{}h(\boldsymbol{\xi})$ in Equation (S18) with respect to $\boldsymbol{\beta}$ is
$\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\boldsymbol{\beta}}=e^{-\omega^{\prime}}\textbf{r}^{\top}\textbf{M}\textbf{X}-\frac{\boldsymbol{\beta}^{\top}}{\sigma^{2}_{\boldsymbol{\beta}}},$ (S19)
the derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $\omega^{\prime}$ is
$\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\omega^{\prime}}=-\frac{n}{2}+\frac{e^{-\omega^{\prime}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}-\frac{\omega^{\prime}}{\sigma^{2}_{\omega^{\prime}}},$ (S20)
the derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $\rho^{\prime}$ is
$\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\rho^{\prime}}=\left(\frac{1}{2}\text{tr}\left\{(\textbf{A}^{\top}\textbf{A})^{-1}\left(\frac{\partial\textbf{A}^{\top}\textbf{A}}{\partial\rho}\right)\right\}+e^{-\omega^{\prime}}(\textbf{A}\textbf{r})^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{W}\textbf{r}\right)\left(\frac{\partial\rho}{\partial\rho^{\prime}}\right)-\frac{\rho^{\prime}}{\sigma^{2}_{\rho^{\prime}}},$ (S21)
and the derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $\nu^{\prime}$ is
$\begin{split}\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\nu^{\prime}}&=n\left(\frac{\partial~{}\textrm{log}~{}\Gamma\left(\frac{e^{\nu^{\prime}+3}}{2}\right)}{\partial\nu^{\prime}}\right)-\frac{e^{\nu^{\prime}}}{2}\sum_{i=1}^{n}\left(\tau_{i}^{\prime}+e^{-\tau_{i}^{\prime}}\right)-\frac{ne^{\nu^{\prime}}}{2}\textrm{log}\left(\frac{2}{e^{\nu^{\prime}}+3}\right)+\frac{ne^{\nu^{\prime}}}{2}\\ &-\frac{\nu^{\prime}}{\sigma^{2}_{\nu^{\prime}}},\\ \end{split}$ (S22)
where $\frac{\partial~{}\textrm{log}~{}\Gamma\left(\frac{e^{\nu^{\prime}+3}}{2}\right)}{\partial\nu^{\prime}}$ is the derivative of the log-gamma function, which is computed numerically using R.
The derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $i^{th}$ element of the vector $\boldsymbol{\tau}^{\prime}$, $\tau_{i}^{\prime}$ is
$\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\tau_{i}^{\prime}}=-\frac{1}{2}+\frac{1}{2}e^{-\omega^{\prime}}({\textbf{r}_{\textbf{A}}}_{i})^{2}e^{\tau_{i}^{\prime}}+\left(\frac{e^{\nu^{\prime}+3}}{2}\right)\left(-1+e^{-\tau_{i}^{\prime}}\right),$ (S23)
where ${\textbf{r}_{\textbf{A}}}_{i}$ is the $i^{th}$ element of the vector Ar.
S2.2.3 Gradients for YJ-SEM-Gau
For YJ-SEM-Gau, $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\gamma)^{\top}$. The logarithm of $h(\boldsymbol{\xi})$ is $\text{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})+\text{log}~{}p(\boldsymbol{\phi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$ and $\gamma$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3. This leads to
$\begin{split}\text{log}~{}h(\boldsymbol{\xi})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right)\\ &-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\gamma^{\prime}}^{2}}{2\sigma^{2}_{\gamma^{\prime}}},\end{split}$ (S24)
where the values of $\sigma^{2}_{\boldsymbol{\beta}}$, $\sigma^{2}_{\omega^{\prime}}$, $\sigma^{2}_{\rho^{\prime}}$ , and $\sigma^{2}_{\gamma^{\prime}}$ are each set to 100, as detailed in Table S3, while r and M are provided in Table S1.
The derivative of $\text{log}~{}h(\boldsymbol{\xi})$ in Equation (S24) with respect to $\boldsymbol{\beta}$ is
$\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\boldsymbol{\beta}}=e^{-\omega^{\prime}}\textbf{r}^{\top}\textbf{M}\textbf{X}-\frac{\boldsymbol{\beta}^{\top}}{\sigma^{2}_{\boldsymbol{\beta}}},$ (S25)
the derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $\omega^{\prime}$ is
$\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\omega^{\prime}}=-\frac{n}{2}+\frac{e^{-\omega^{\prime}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}-\frac{\omega^{\prime}}{\sigma^{2}_{\omega^{\prime}}},$ (S26)
the derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $\rho^{\prime}$ is
$\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\rho^{\prime}}=\left(\frac{1}{2}\text{tr}\left\{(\textbf{A}^{\top}\textbf{A})^{-1}\left(\frac{\partial\textbf{A}^{\top}\textbf{A}}{\partial\rho}\right)\right\}-\frac{e^{-\omega^{\prime}}}{2}\textbf{r}^{\top}\left(\frac{\partial\textbf{A}^{\top}\textbf{A}}{\partial\rho}\right)\textbf{r}\right)\left(\frac{\partial\rho}{\partial\rho^{\prime}}\right)-\frac{\rho^{\prime}}{\sigma^{2}_{\rho^{\prime}}},$ (S27)
and the derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $\gamma^{\prime}$ is
$\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\gamma^{\prime}}=\left(-e^{-\omega^{\prime}}\textbf{r}^{\top}\textbf{M}\left(\frac{\partial t_{\gamma}(\textbf{y})}{\partial\gamma}\right)+\sum_{i=0}^{n}\frac{\partial\textrm{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right)}{\partial\gamma}\right)\left(\frac{\partial\gamma}{\partial\gamma^{\prime}}\right)-\frac{\gamma^{\prime}}{\sigma^{2}_{\gamma^{\prime}}},$ (S28)
where $\frac{\partial t_{\gamma}(\textbf{y})}{\partial\gamma}=\left(\frac{\partial t_{\gamma}(y_{1})}{\partial\gamma},\frac{\partial t_{\gamma}({y}_{2})}{\partial\gamma},\ldots,\frac{\partial t_{\gamma}({y}_{n})}{\partial\gamma}\right)^{\top}$, with the $i^{th}$ elements of the vector is equal to
$\frac{\partial t_{\gamma}({y}_{i})}{\partial\gamma}=\begin{cases}\frac{(y_{i}+1)^{\gamma}(\gamma\textrm{log}(y_{i}+1)-1)+1}{\gamma^{2}}&\text{if }y_{i}\geq 0,\\ \frac{(2-\gamma)(-y_{i}+1)^{(2-\gamma)}\textrm{log}(-y_{i}+1)-(-y_{i}+1)^{2-\gamma}+1}{(2-\gamma)^{2}}&\text{if }y_{i}<0,\end{cases}$
$\frac{\partial\textrm{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right)}{\partial\gamma}=\begin{cases}\textrm{log}~{}(y_{i}+1)&\text{if }y_{i}\geq 0,\\ -\textrm{log}~{}(y_{i}+1)&\text{if }y_{i}<0,\end{cases}$
and
$\frac{\partial\gamma}{\partial\gamma^{\prime}}=\frac{2e^{\gamma^{\prime}}}{(1+e^{\gamma^{\prime}})^{2}}.$
S2.2.4 Gradients for YJ-SEM-t
For YJ-SEM-t, $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$. The logarithm of $h(\boldsymbol{\xi})$ is $\text{log}~{}p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\text{log}~{}p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\text{log}~{}p(\boldsymbol{\phi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, $\nu$, and $\gamma$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3. This leads to
$\begin{split}\text{log}~{}h(\boldsymbol{\xi})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right)\\ &+\sum_{i=1}^{n}\textrm{log}~{}p(\tau_{i}\mid\nu)+\textrm{log}\left|\frac{\partial\tau_{i}}{\partial\tau_{i}^{\prime}}\right|-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\nu^{\prime}}^{2}}{2\sigma^{2}_{\nu^{\prime}}},\end{split}$ (S29)
where the values of $\sigma^{2}_{\boldsymbol{\beta}}$, $\sigma^{2}_{\omega^{\prime}}$, $\sigma^{2}_{\rho^{\prime}}$, $\sigma^{2}_{\nu^{\prime}}$, and $\sigma^{2}_{\gamma^{\prime}}$ are each set to 100, as detailed in Table S3, while r and M are provided in Table S1.
The derivative of $\text{log}~{}h(\boldsymbol{\xi})$ in Equation (S29) with respect to $\boldsymbol{\beta}$ is
$\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\boldsymbol{\beta}}=e^{-\omega^{\prime}}\textbf{r}^{\top}\textbf{M}\textbf{X}-\frac{\boldsymbol{\beta}^{\top}}{\sigma^{2}_{\boldsymbol{\beta}}},$ (S30)
the derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $\omega^{\prime}$ is
$\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\omega^{\prime}}=-\frac{n}{2}+\frac{e^{-\omega^{\prime}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}-\frac{\omega^{\prime}}{\sigma^{2}_{\omega^{\prime}}},$ (S31)
the derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $\rho^{\prime}$ is
$\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\rho^{\prime}}=\left(\frac{1}{2}\text{tr}\left\{(\textbf{A}^{\top}\textbf{A})^{-1}\left(\frac{\partial\textbf{A}^{\top}\textbf{A}}{\partial\rho}\right)\right\}+e^{-\omega^{\prime}}(\textbf{A}\textbf{r})^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{W}\textbf{r}\right)\left(\frac{\partial\rho}{\partial\rho^{\prime}}\right)-\frac{\rho^{\prime}}{\sigma^{2}_{\rho^{\prime}}},$ (S32)
the derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $\nu^{\prime}$ is
$\begin{split}\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\nu^{\prime}}&=n\left(\frac{\partial~{}\textrm{log}~{}\Gamma\left(\frac{e^{\nu^{\prime}+3}}{2}\right)}{\partial\nu^{\prime}}\right)-\frac{e^{\nu^{\prime}}}{2}\sum_{i=1}^{n}\left(\tau_{i}^{\prime}+e^{-\tau_{i}^{\prime}}\right)-\frac{ne^{\nu^{\prime}}}{2}\textrm{log}\left(\frac{2}{e^{\nu^{\prime}}+3}\right)+\frac{ne^{\nu^{\prime}}}{2}\\ &-\frac{\nu^{\prime}}{\sigma^{2}_{\nu^{\prime}}},\\ \end{split}$ (S33)
and the derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $\gamma^{\prime}$ is
$\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\gamma^{\prime}}=\left(-e^{-\omega^{\prime}}\textbf{r}^{\top}\textbf{M}\left(\frac{\partial t_{\gamma}(\textbf{y})}{\partial\gamma}\right)+\sum_{i=0}^{n}\frac{\partial\textrm{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right)}{\partial\gamma}\right)\left(\frac{\partial\gamma}{\partial\gamma^{\prime}}\right)-\frac{\gamma^{\prime}}{\sigma^{2}_{\gamma^{\prime}}}.$ (S34)
The derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $i^{th}$ element of the vector $\boldsymbol{\tau}^{\prime}$, $\tau_{i}^{\prime}$ is
$\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\tau_{i}^{\prime}}=-\frac{1}{2}+\frac{1}{2}e^{-\omega^{\prime}}({\textbf{r}_{\textbf{A}}}_{i})^{2}e^{\tau_{i}^{\prime}}+\left(\frac{e^{\nu^{\prime}+3}}{2}\right)\left(-1+e^{-\tau_{i}^{\prime}}\right),$ (S35)
where ${\textbf{r}_{\textbf{A}}}_{i}$ is the $i^{th}$ element of the vector Ar.
S3 Calculate adaptive learning rates using ADADELTA
The adaptive learning rates (step sizes) for the VB algorithms are calculated using the ADADELTA algorithm (Zeiler,, 2012). The ADADELTA algorithm is now briefly described. Different step sizes are used for each element in variational parameters $\boldsymbol{\lambda}$.
The update for the $i^{th}$ element of $\boldsymbol{\lambda}$ is
$\boldsymbol{\lambda}^{(t+1)}_{i}=\boldsymbol{\lambda}^{(t)}_{i}+\Delta\boldsymbol{\lambda}^{(t)}_{i},$ (S36)
where, the step size $\Delta\boldsymbol{\lambda}^{(t)}_{i}$ is $a_{i}^{(t)}g_{\boldsymbol{\lambda}_{i}}^{(t)}$. The term $g_{\boldsymbol{\lambda}_{i}}^{(t)}$ denotes the $i^{th}$ component of $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$ and $a_{i}^{(t)}$ is defined as:
$\mathcal{a}_{i}^{(t)}=\sqrt{\frac{E\left(\boldsymbol{\Delta}_{\boldsymbol{\lambda}_{i}}^{2}\right)^{(t-1)}+\alpha}{E\left(g^{2}_{\boldsymbol{\lambda}_{i}}\right)^{(t)}+\alpha}},$ (S37)
where $\alpha$ is a small positive constant, $E\left(\boldsymbol{\Delta}_{\boldsymbol{\lambda_{i}}}^{2}\right)^{(t)}$ and $E\left(g^{2}_{\boldsymbol{\lambda}_{i}}\right)^{(t)}$ are decayed moving average estimates of ${\boldsymbol{\Delta}{\boldsymbol{\lambda_{i}}}^{(t)}}^{2}$ and ${g_{\boldsymbol{\lambda}_{i}}^{(t)}}^{2}$, defined by
$E\left(\boldsymbol{\Delta}_{\boldsymbol{\lambda}_{i}}^{2}\right)^{(t)}=\upsilon E\left(\boldsymbol{\Delta}_{\boldsymbol{\lambda}_{i}}^{2}\right)^{(t-1)}+(1-\upsilon){\boldsymbol{\Delta}{\boldsymbol{\lambda_{i}}}^{(t)}}^{2},$ (S38)
and
$E\left(g^{2}_{\boldsymbol{\lambda}_{i}}\right)^{(t)}=\upsilon E\left(g^{2}_{\boldsymbol{\lambda}_{i}}\right)^{(t-1)}+(1-\upsilon){g_{\boldsymbol{\lambda}_{i}}^{(t)}}^{2},$ (S39)
where the variable $\upsilon$ is a decay constant. We use the default tuning parameter choices $\alpha=10^{-6}$ and $\upsilon=0.95$, and initialize $E\left(\boldsymbol{\Delta}_{\boldsymbol{\lambda}_{i}}^{2}\right)^{(0)}=E\left(g^{2}_{\boldsymbol{\lambda}_{i}}\right)^{(0)}=0$.
S4 Additional Derivation of HVB algorithms for estimating SEMs with missing data
S4.1 Derivation of the reparameterisation gradient for HVB algorithm
Since $\mathcal{L}(\boldsymbol{\lambda})=E_{q}\left[\text{log}~{}p(\textbf{y}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})+\text{log}~{}p(\boldsymbol{\xi},\boldsymbol{\psi})-\text{log}~{}q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\right]=\mathcal{L}^{0}(\boldsymbol{\lambda})$ as shown in Equation (16) of the main paper. Then, the reparameterisation gradient of $\mathcal{L}$ is the same as that of $\mathcal{L}^{0}$, given by
$\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$ $=E_{f_{\boldsymbol{\delta}^{0}}}\Bigg{[}\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top}}{d\boldsymbol{\lambda_{s_{m}}}}\Big{(}\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}p(\boldsymbol{\xi},\boldsymbol{\psi})$ (S40)
$\qquad+\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}p(\textbf{y}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})-\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\theta})\Big{)}\Bigg{]},$
where, the random vector $\boldsymbol{\delta}^{0}$ has density $f_{\boldsymbol{\delta}^{0}}$, which follows a standard normal, and does not depend on $\boldsymbol{\lambda_{s_{m}}}$, and $t^{0}$ is the one-to-one vector-valued transformation from $\boldsymbol{\delta}^{0}=({\boldsymbol{\eta}^{0}}^{\top},{\boldsymbol{\epsilon}^{0}}^{\top})^{\top}$ to the parameter vector, such that $(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}=t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})=\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}$.
Then, the Fisher’s identity is given by
$\begin{split}\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}p(\textbf{y}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})&=\int\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\left[\text{log}~{}(p(\textbf{y}_{o},\textbf{m}\mid\textbf{y}_{u},\boldsymbol{\xi},\boldsymbol{\psi})p(\textbf{y}_{u}\mid\boldsymbol{\xi}))\right]\\ &p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi},\boldsymbol{\psi})d\textbf{y}_{u},\end{split}$ (S41)
see Poyiadjis et al., (2011). Substituting this expression into Equation (S40), and writing $E_{f_{\boldsymbol{\delta}}}(\cdot)$ for expectation with respect to $f_{\boldsymbol{\delta}}(\boldsymbol{\delta})=f_{\boldsymbol{\delta}^{0}}(\boldsymbol{\delta}^{0})p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ and because $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}_{o},\textbf{m}\mid\textbf{y}_{u},\boldsymbol{\xi},\boldsymbol{\psi})p(\textbf{y}_{u}\mid\boldsymbol{\xi})p(\boldsymbol{\xi},\boldsymbol{\psi})$, we get
$\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$ $={E}_{f_{\boldsymbol{\delta}}}\Bigg{[}\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top}}{d\boldsymbol{\lambda}_{s_{m}}}\Big{(}\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log p(\boldsymbol{\xi},\boldsymbol{\psi})$ (S42)
$\qquad+\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log p(\textbf{y}_{u}\mid\boldsymbol{\xi})+\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log p(\textbf{y}_{o},\textbf{m}\mid\textbf{y}_{u},\boldsymbol{\xi},\boldsymbol{\psi})$
$\qquad-\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\Big{)}\Bigg{]}$
$={E}_{f_{\boldsymbol{\delta}}}\left[\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top}}{d\boldsymbol{\lambda}_{s_{m}}}\left(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})-\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\right)\right].$
The term $\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})}{d\boldsymbol{\lambda_{s_{m}}}}$ in Equation (S42) is the derivative of the transformation $t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})=\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}$ with respect to the variational parameters $\boldsymbol{\lambda}_{s_{m}}=(\boldsymbol{\mu}_{s_{m}}^{\top},\text{vech}(\textbf{B}_{s_{m}})^{\top},\textbf{d}_{s_{m}}^{\top})^{\top}$. We can express that $t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})=\boldsymbol{\mu}_{s_{m}}+(\boldsymbol{\eta}^{0}\otimes\textbf{I}_{s_{m}})\text{vech}(\textbf{B}_{s_{m}})+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}$, where $\textbf{I}_{s_{m}}$ is the identity matrix of size $s_{m}$, and it can be further shown that $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}q^{0}_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi})=-(\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2})^{-1}((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}-\boldsymbol{\mu}_{s_{m}})$,
$\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})}{d\boldsymbol{\mu}_{s_{m}}}=\textbf{I}_{s_{m}}\\ ~{}~{}~{}~{}~{}\text{and}~{}~{}~{}~{}~{}\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})}{d\text{vech}(\textbf{B}_{s_{m}})}={\boldsymbol{\eta}^{0}}^{\top}\otimes\textbf{I}_{s_{m}}.$ (S43)
Then the derivatives of $\mathcal{L}(\boldsymbol{\lambda})$ with respect to $\boldsymbol{\mu}_{s_{m}},\textrm{vech}(\textbf{B}_{s_{m}})$, and $\textbf{d}_{s_{m}}$ are:
$\begin{split}\nabla_{\boldsymbol{\mu}_{s_{m}}}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\delta}}}[\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi^{\top})^{\top}}}~{}\text{log}~{}h(\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u})\\ &+(\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2})^{-1}(\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0})],\end{split}$ (S44)
$\begin{split}\nabla_{\textrm{vech}(\textbf{B}_{s_{m}})}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\delta}}}[\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi^{\top})^{\top}}}~{}\text{log}~{}h(\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u}){\boldsymbol{\eta}^{0}}^{\top}\\ &+(\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2})^{-1}(\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}){\boldsymbol{\eta}^{0}}^{\top}],\end{split}$ (S45)
$\begin{split}\nabla_{\textbf{d}_{s_{m}}}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\delta}}}[\text{diag}(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi^{\top})^{\top}}}\text{log}~{}h(\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u}){\boldsymbol{\epsilon}^{0}}^{\top}\\ &+(\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2})^{-1}(\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}){\boldsymbol{\epsilon}^{0}}^{\top})].\end{split}$ (S46)
The expressions for $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u})=\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ in Equations (S44)-(S46) can be found in Section S4.3. The expectations in these gradients are estimated using a single sample of $\boldsymbol{\delta}=({\boldsymbol{\delta}^{0}}^{\top},\textbf{y}_{u}^{\top})^{\top}$, which is drawn from $f_{\boldsymbol{\delta}^{0}}$, and $p(\textbf{y}_{u}\mid\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{o},\textbf{m})=p(\textbf{y}_{u}\mid t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}}),\textbf{y}_{o},\textbf{m})$; see  Wijayawardhana et al., (2025) for further details.
S4.2 HVB-AllB algorithm
This section presents the HVB-AllB algorithm for estimating SEMs with missing values, which is particularly suited for cases where both $n$ and $n_{u}$ are large.
The MCMC steps in Algorithm 2 of the main paper generates samples from the conditional distribution $p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$. However, as $n$ and $n_{u}$ increase, the HVB algorithm implemented using these MCMC steps does not estimate the parameters accurately because of the low acceptance percentage. To improve the acceptance percentage, we partition $\textbf{y}_{u}$ into $k$ blocks and update one block at a time.
We start with partitioning the unobserved responses vector into $k$ blocks, such that $\textbf{y}_{u}=(\textbf{y}_{u_{1}}^{\top},\dots,\textbf{y}_{u_{k}}^{\top})^{\top}$. Using proposals from $p(\textbf{y}_{u_{j}}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{o},\textbf{y}_{u}^{(-j)})$, the Algorithm S2 outlines the MCMC steps for sampling the missing values one block at a time, for $j=1,..,k$, where $\textbf{y}_{u_{j}}$ is the updated block and $\textbf{y}_{u}^{(-j)}=(\textbf{y}_{u_{1}}^{\top},\dots,\textbf{y}_{u_{j-1}}^{\top},\textbf{y}_{u_{j+1}}^{\top},\dots,\textbf{y}_{u_{k}}^{\top})^{\top}$ is the remaining blocks. The complete response vector y can now be written as $\textbf{y}=({\textbf{y}^{\top}_{s_{j}}},\textbf{y}_{u_{j}}^{\top})^{\top}$, where ${\textbf{y}_{s_{j}}}=(\textbf{y}_{o}^{\top},{\textbf{y}_{u}^{(-j)}}^{\top})^{\top}$. Based on this partitioning of y, the following partitioned vector and matrices are defined:
$\textbf{r}=\begin{pmatrix}\textbf{r}_{s_{j}}\\ \textbf{r}_{u_{j}}\end{pmatrix},~{}\textbf{X}=\begin{pmatrix}\textbf{X}_{s_{j}}\\ \textbf{X}_{u_{j}}\end{pmatrix},~{}\textbf{M}=\begin{pmatrix}\textbf{M}_{{s_{j}}{s_{j}}}&\textbf{M}_{{s_{j}}{u_{j}}}\\ \textbf{M}_{{u_{j}}{s_{j}}}&\textbf{M}_{{u_{j}}{u_{j}}}\end{pmatrix},$ (S47)
where $\textbf{r}_{s_{j}}$ is the corresponding sub-vector of r for the observed responses ($\textbf{y}_{o}$), and the unobserved responses that are not in the $j^{th}$ block (${\textbf{y}_{u}^{(-j)}}$). i.e. $\textbf{r}_{s_{j}}=({{\textbf{r}}}^{\top}_{o},{{\textbf{r}}^{\top}_{u^{(-j)}}})^{\top}$ and ${\textbf{r}}_{u_{j}}$ is the corresponding sub vector of r for $j^{th}$ block of unobserved responses, $\textbf{y}_{u_{j}}$. Similarly, $\textbf{X}_{s_{j}}$ and $\textbf{X}_{u_{j}}$ are represent the sub-matrices of X, and $\textbf{M}_{{s_{j}}{s_{j}}}$, $\textbf{M}_{{s_{j}}{u_{j}}}$, $\textbf{M}_{{u_{j}}{s_{j}}}$, and $\textbf{M}_{{u_{j}}{u_{j}}}$ are sub-matrices of M.
Algorithm S2 MCMC steps within the $t^{th}$ iteration of the HVB algorithm under MNAR. The missing values are updated one block at a time.
1:  Initialise missing values $\textbf{y}_{u,0}=(\textbf{y}^{\top}_{u_{1},0},\dots,\textbf{y}^{\top}_{u_{k},0})^{\top}~{}\sim p(\textbf{y}_{u}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{o})$
2:  for i=1, …, $N_{1}$ do
3:     for j=1, …, $k$ do
4:        Sample $\widetilde{\textbf{y}}_{u_{j}}$ from the proposal distribution $p(\widetilde{\textbf{y}}_{u_{j}}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{o},\textbf{y}_{u,{i-1}}^{(-j)})$, where $\textbf{y}_{u,i-1}^{(-j)}=(\textbf{y}^{\top}_{u_{1},i},\textbf{y}^{\top}_{u_{2},i},\ldots,\textbf{y}^{\top}_{u_{j-1},i},\textbf{y}^{\top}_{u_{j+1},i-1},\ldots,,\textbf{y}^{\top}_{u_{k},i-1})^{\top}$.
5:        Sample $u$ from uniform distribution, $u~{}\sim\mathcal{U}(0,1)$
6:        Calculate $a=\text{min}\left(1,\frac{p(\textbf{m}\mid\widetilde{\textbf{y}},\boldsymbol{\psi}^{(t)})}{p(\textbf{m}\mid\textbf{y}_{i-1},\boldsymbol{\psi}^{(t)})}\right)$, where $\widetilde{\textbf{y}}=(\textbf{y}_{o}^{\top},{\textbf{y}_{u}^{(-j)}}^{\top},{\widetilde{\textbf{y}}_{u_{j}}}^{\top})^{\top}$ and $\textbf{y}_{i-1}=(\textbf{y}_{o}^{\top},{\textbf{y}_{u}^{(-j)}}^{\top},{\textbf{y}_{u_{j},{i-1}}}^{\top})^{\top}$
7:        if $a>u$ then
8:           $\textbf{y}_{u_{j},i}=\textbf{y}_{u_{j}}^{*}$
9:        else
10:           $\textbf{y}_{u_{j},i}=\textbf{y}_{{u_{j},i-1}}$
11:        end if
12:     end for
13:     $\textbf{y}_{u,i}=(\textbf{y}_{u_{1},i}^{\top},\ldots,\textbf{y}_{u_{k},i}^{\top})^{\top}$
14:  end for
15:  Output $\textbf{y}^{(t)}_{u}=\textbf{y}_{u,N_{1}}$
In step 4 of the Algorithm S2, proposals from the conditional distributions, $p(\widetilde{\textbf{y}}_{u_{j}}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{s_{j}})$ needs to be generated. For the SEM-Gau, and SEM-t, this distribution follows a multivariate Gaussian distribution with the mean $\textbf{X}_{u_{j}}\boldsymbol{\beta}-\textbf{M}_{{u_{j}}{u_{j}}}^{-1}\textbf{M}_{{u_{j}}{s_{j}}}\textbf{r}_{s_{j}}$ and the covariance matrix $\sigma^{2}_{\textbf{e}}\textbf{M}_{{u_{j}}{u_{j}}}^{-1}$.
For YJ-SEM-Gau and YJ-SEM-t, the proposals are generated in two steps. First, we sample $\widetilde{\textbf{y}}_{u_{j}}^{\ast}$ from the full conditional distribution $p({\widetilde{\textbf{y}}^{\ast}}_{u_{j}}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{s_{j}})$, which follows a multivariate Gaussian with the mean vector given by $\textbf{X}_{u_{j}}\boldsymbol{\beta}-\textbf{M}_{{u_{j}}{u_{j}}}^{-1}\textbf{M}_{{u_{j}}{s_{j}}}(\textbf{y}^{\ast}_{s_{j}}-\textbf{X}_{s_{j}}\boldsymbol{\beta})$ and the covariance matrix $\sigma^{2}_{\textbf{e}}\textbf{M}_{{u_{j}}{u_{j}}}^{-1}$, where $\textbf{y}^{\ast}_{s_{j}}=t_{\gamma}(\textbf{y}_{s_{j}})$. Then, we apply the inverse Yeo-Johnson (YJ) transformation to obtain the final proposal: $\widetilde{\textbf{y}}_{u_{j}}=t_{\gamma}^{-1}(\widetilde{\textbf{y}}_{u_{j}}^{*})$.
S4.3 Expressions for $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$
This section presents the derivatives $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$, which are required to compute the gradients of the reparameterised ELBO, defined in Equation (17) of the main paper for different SEMs with missing data.
S4.3.1 Derivatives for SEM-Gau
For SEM-Gau with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}$. The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})+\textrm{log}~{}p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~{}p(\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\psi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, and $\rho$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3. This leads to
$\begin{split}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}\\ &+\sum_{i=1}^{n}m_{i}(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{\boldsymbol{\psi}^{\top}\boldsymbol{\psi}}{2\sigma^{2}_{\boldsymbol{\psi}}},\end{split}$ (S48)
where the values of $\sigma^{2}_{\boldsymbol{\beta}}$, $\sigma^{2}_{\omega^{\prime}}$, $\sigma^{2}_{\rho^{\prime}}$, and $\sigma^{2}_{\boldsymbol{\psi}}$ are each set to 100, as detailed in Table S3, while r and M are provided in Table S1.
The derivatives of $\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ in Equation (S48) with respect to $\boldsymbol{\beta}$, ${\omega^{\prime}}$, and ${\rho^{\prime}}$ are similar to that of without missing data given in Equations (S15),  (S16), and (S17), respectively. The derivative of $\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ with respect to $\boldsymbol{\psi}$ is
$\frac{\partial\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})}{\partial\boldsymbol{\psi}}=\sum_{i=1}^{n}\left(m_{i}-\frac{e^{{\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y}}}{1+e^{{\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y}}}\right)\textbf{z}_{i}-\frac{\boldsymbol{\psi}^{\top}}{\sigma^{2}_{\boldsymbol{\psi}}},$ (S49)
where $\textbf{z}_{i}=({\textbf{x}^{*}_{i}},y_{i})$ is the vector containing the $i^{th}$ row vector of matrix $\textbf{X}^{*}$, and the $i^{th}$ element of the vector y, see Section 2.3 of the main paper.
S4.3.2 Derivatives for SEM-t
For SEM-t with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\nu,\rho)^{\top}$. The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~{}p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\textrm{log}~{}p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~{}p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\psi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, and $\nu$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3. This leads to
$\begin{split}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}\\ &+\sum_{i=1}^{n}m_{i}(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &+\sum_{i=1}^{n}\textrm{log}~{}p(\tau_{i}\mid\nu)+\textrm{log}\left|\frac{\partial\tau_{i}}{\partial\tau_{i}^{\prime}}\right|-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\nu^{\prime}}^{2}}{2\sigma^{2}_{\nu^{\prime}}},\end{split}$ (S50)
where the values of $\sigma^{2}_{\boldsymbol{\beta}}$, $\sigma^{2}_{\omega^{\prime}}$, $\sigma^{2}_{\rho^{\prime}}$, $\sigma^{2}_{\nu^{\prime}}$, and $\sigma^{2}_{\boldsymbol{\psi}}$ are each set to 100, as detailed in Table S3, while r and M are provided in Table S1.
The derivatives of $\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ in Equation (S50) with respect to $\boldsymbol{\beta}$, ${\omega^{\prime}}$, ${\rho^{\prime}}$, ${\nu^{\prime}}$ and $\tau_{i}^{\prime}$ are similar to that of without missing data given in Equations (S19),  (S20), (S21), (S22), and (S23), respectively. The derivative of $\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ with respect to $\boldsymbol{\psi}$ is similar to that of SEM-Gau given in Equation (S49).
S4.3.3 Derivatives for YJ-SEM-Gau
For YJ-SEM-Gau with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\gamma)^{\top}$. The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})+\textrm{log}~{}p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~{}p(\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\psi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, and $\gamma$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3. This leads to
$\begin{split}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right)\\ &+\sum_{i=1}^{n}m_{i}(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\gamma^{\prime}}^{2}}{2\sigma^{2}_{\gamma^{\prime}}}-\frac{\boldsymbol{\psi}^{\top}\boldsymbol{\psi}}{2\sigma^{2}_{\boldsymbol{\psi}}},\end{split}$ (S51)
where the values of $\sigma^{2}_{\boldsymbol{\beta}}$, $\sigma^{2}_{\omega^{\prime}}$, $\sigma^{2}_{\rho^{\prime}}$, $\sigma^{2}_{\nu^{\prime}}$, $\sigma^{2}_{\gamma^{\prime}}$, and $\sigma^{2}_{\boldsymbol{\psi}}$ are each set to 100, as detailed in Table S3, while r and M are provided in Table S1.
The derivatives of $\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ in Equation (S51) with respect to $\boldsymbol{\beta}$, ${\omega^{\prime}}$, ${\rho^{\prime}}$, and ${\gamma^{\prime}}$ are similar to that of without missing data given in Equations (S25),  (S26), (S27), and (S28), respectively. The derivative of $\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ with respect to $\boldsymbol{\psi}$ is similar to that of SEM-Gau given in Equation (S49).
S4.3.4 Derivatives for YJ-SEM-t
For YJ-SEM-t with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$. The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~{}p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\textrm{log}~{}p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~{}p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\psi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, $\nu$, and $\gamma$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3. This leads to
$\begin{split}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right)\\ &+\sum_{i=1}^{n}m_{i}(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &+\sum_{i=1}^{n}\textrm{log}~{}p(\tau_{i}\mid\nu)+\textrm{log}\left|\frac{\partial\tau_{i}}{\partial\tau_{i}^{\prime}}\right|-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\nu^{\prime}}^{2}}{2\sigma^{2}_{\nu^{\prime}}}-\frac{{\gamma^{\prime}}^{2}}{2\sigma^{2}_{\gamma^{\prime}}},\end{split}$ (S52)
where the values of $\sigma^{2}_{\boldsymbol{\beta}}$, $\sigma^{2}_{\omega^{\prime}}$, $\sigma^{2}_{\rho^{\prime}}$, $\sigma^{2}_{\nu^{\prime}}$, $\sigma^{2}_{\gamma^{\prime}}$, and $\sigma^{2}_{\boldsymbol{\psi}}$ are each set to 100, as detailed in Table S3, while r and M are provided in Table S1.
The derivatives of $\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ in Equation S52 with respect to $\boldsymbol{\beta}$, ${\omega^{\prime}}$, ${\rho^{\prime}}$, ${\nu^{\prime}}$, ${\gamma^{\prime}}$ and $\tau_{i}^{\prime}$ are similar to that of without missing data given in Equations (S30),  (S31), (S32), (S33), (S34), and (S35), respectively. The derivative of $\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ with respect to $\boldsymbol{\psi}$ is similar to that of SEM-Gau given in Equation (S49).
S5 Bayesian model comparison of spatial error models
This section presents expressions for the term log $p(\mathbf{y}\mid\boldsymbol{\phi})$, which are used to compute the Deviance Information Criterion (DIC) (Celeux et al.,, 2006) for the various SEMs discussed in Section 3.3 of the main paper.
Note that for SEM-Gau, and YJ-SEM-Gau, $\boldsymbol{\xi}=\boldsymbol{\phi}$, meaning for these two models $p(\mathbf{y}\mid\boldsymbol{\phi})=p(\mathbf{y}\mid\boldsymbol{\xi})$. For the SEM-Gau model, $p(\mathbf{y}\mid\boldsymbol{\xi})$ is Gaussian; see Section 2.1 of the main paper. The log-likelihood is given by:
$\text{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})=-\frac{n}{2}\textrm{log}(2\pi)-\frac{n}{2}\textrm{log}(\sigma^{2}_{\boldsymbol{e}})+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{1}{2\sigma^{2}_{\boldsymbol{e}}}\textbf{r}^{\top}\textbf{M}\textbf{r},$ (S53)
where M, and r are given in Table S1.
For the YJ-SEM-Gau, $p(\mathbf{y}\mid\boldsymbol{\phi})=p(\mathbf{y}\mid\boldsymbol{\xi})$ is derived in Section 2.2 of the main paper, and the log-likelihood is given by:
$\begin{split}\text{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})&=-\frac{n}{2}\textrm{log}(2\pi)-\frac{n}{2}\textrm{log}(\sigma^{2}_{\boldsymbol{e}})+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{1}{2\sigma^{2}_{\boldsymbol{e}}}\textbf{r}^{\top}\textbf{M}\textbf{r}\\ &+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right),\end{split}$ (S54)
where M, and r are given in Table S1, while $\frac{dt_{\gamma}(y_{i})}{dy_{i}}$ is given in Equation (S2).
For the SEM-t model,
[…]
where $p(\mathbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})$ is multivariate Gaussian, with its mean vector and covariance matrix specified in Table S1. Each $p({\tau}_{i}\mid\boldsymbol{\phi})$ (for $i=1,\ldots,n$) follows an inverse gamma distribution with both shape and scale parameters equal to $\nu/2$; see Section 2.1 of the main paper. This hierarchical representation implies that the marginal distribution $p(\mathbf{y}\mid\boldsymbol{\phi})$ is a multivariate Student’s $t$-distribution with mean vector $\textbf{X}\boldsymbol{\beta}$, scale matrix $\sigma^{2}_{\boldsymbol{e}}(\mathbf{A}^{\top}\mathbf{A})^{-1}$, and degrees of freedom $\nu$. The log-likelihood is given by:
$\begin{split}\text{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})&=\textrm{log}~{}\Gamma\left(\frac{\nu+2}{2}\right)-\textrm{log}~{}\Gamma\left(\frac{\nu}{2}\right)+\frac{1}{2}\textrm{log}|\textbf{A}^{\top}\textbf{A}|\\ &-\frac{n}{2}\textrm{log}(\pi\nu\sigma^{2}_{\textbf{e}})-\frac{(n+\nu)}{2}\textrm{log}\left(1+\frac{1}{\nu\sigma^{2}_{\boldsymbol{e}}}\textbf{r}^{\top}\textbf{A}^{\top}\textbf{A}\textbf{r}\right),\end{split}$ (S56)
where $\textbf{r}=\textbf{y}-\textbf{X}\boldsymbol{\beta}$.
Similarly, logarithm of $p(\textbf{y}\mid\boldsymbol{\phi})$ for YJ-SEM-t can be derived as
$\begin{split}\text{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})&=\textrm{log}~{}\Gamma\left(\frac{\nu+2}{2}\right)-\textrm{log}~{}\Gamma\left(\frac{\nu}{2}\right)+\frac{1}{2}\textrm{log}|\textbf{A}^{\top}\textbf{A}|\\ &-\frac{n}{2}\textrm{log}(\pi\nu\sigma^{2}_{\textbf{e}})-\frac{(n+\nu)}{2}\textrm{log}\left(1+\frac{1}{\nu\sigma^{2}_{\boldsymbol{e}}}\textbf{r}^{\top}\textbf{A}^{\top}\textbf{A}\textbf{r}\right)\\ &+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right),\end{split}$ (S57)
where $\textbf{r}=t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$.
S6 Simulation study with full data SEMs
This section fits the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models to simulated datasets 1 and 2 (without missing data) in Section 4.2 of the main paper, and evaluates how effectively these models, along with their VB estimation methods, capture the characteristics of these datasets.
The initial values for the VB algorithms are set as follows: The variational mean vector ($\boldsymbol{\mu}$) is initialised using the estimates of $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ obtained from fitting the SEM-Gau model via maximum likelihood estimation (MLE). The initial values for $\nu$ and $\gamma$ are set to 4 and 1, respectively. For the variational covariance matrix parameters, the elements of B and the diagonal elements of D are all initialised to 0.01. We use $p=4$ factors.
After some experimentation, we found that the models with Student’s $t$ errors required a relatively large number of iterations to converge. This increased computational demand is attributed to the added complexity of estimating both the latent variables $\boldsymbol{\tau}$ and the degrees of freedom parameter $\nu$. For simulated dataset 1, the VB algorithms were run for 10,000 iterations for both SEM-Gau and YJ-SEM-Gau, and 20,000 iterations for SEM-t and YJ-SEM-t. For simulated dataset 2, convergence was achieved after 10,000 iterations for SEM-Gau and YJ-SEM-Gau, and 40,000 iterations for SEM-t and YJ-SEM-t. Convergence is assessed visually by inspecting the variational mean plots over iterations, which are presented in Section S9.1. Then, we generate 10,000 draws of model parameters from the variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$.
Table S4 shows the posterior means and the 95% credible intervals for some of the model parameters of the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models obtained using the Gaussian variational approximation method described in Section S2 applied to the simulated dataset 1. The table also includes the computation times (in seconds) for one VB iteration, as well as the $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values for each model, calculated using Equations (18) and (19) of the main paper, respectively. The plots of posterior densities of the parameters are provided in Section S7.2.1.
We begin by comparing the estimated values of the fixed effects (specifically $\beta_{0}$ and $\beta_{1}$), the variance parameter $\sigma_{\textbf{e}}^{2}$, and the spatial autocorrelation parameter $\rho$, as these parameters are present in all four models. Accurate estimation of these parameters is crucial for interpreting the underlying spatial relationships and the effect of covariates on the response variable in SEMs. The posterior means for $\beta_{0}$, $\beta_{1}$ and $\rho$ are nearly identical for YJ-SEM-Gau and YJ-SEM-t, closely matching the true values of $-1$ for $\beta_{0}$, $1$ for $\beta_{1}$, and $0.8$ for $\rho$. In contrast, the estimates from SEM-Gau and SEM-t show significant deviations from the true values. The estimated posterior means of $\sigma_{\textbf{e}}^{2}$ from SEM-Gau, SEM-t, and YJ-SEM-Gau deviate considerably from the true value, while the posterior mean obtained from YJ-SEM-t is much closer to the true value, as expected. The models incorporating the YJ transformation, YJ-SEM-Gau and YJ-SEM-t, successfully recover the parameter $\gamma$. The degrees of freedom ($\nu$) estimated by the SEM-t and YJ-SEM-t models show slight deviations from the true value.
For the simulated dataset 1, the model fit assessment indicates that YJ-SEM-t is the best-performing model, as it achieves the lowest $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values among all four models. This outcome is expected, given that the dataset was simulated from YJ-SEM-t. The second-best model is YJ-SEM-Gau, which has the next lowest DIC values. Notably, the difference in DIC between YJ-SEM-t and YJ-SEM-Gau is minimal, suggesting that YJ-SEM-Gau provides a model fit nearly as good as YJ-SEM-t for this dataset.
SEM-Gau SEM-t YJ-SEM-Gau YJ-SEM-t
${\beta}_{0}=-1$ 3.1962 (2.9423, 3.4453) 0.9017 (0.8048, 0.9978) -1.0319 (-1.0709, -0.9931) -1.0065 (-1.0391, -0.9742)
${\beta}_{1}=1$ 1.5352 (1.3089, 1.7603) 0.9097 (0.8079, 1.0136) 0.9941 (0.9669, 1.0212) 0.9880 (0.9568, 1.0192)
$\sigma^{2}_{\textbf{e}}=0.5$ 58.4686 (56.0301, 60.9582) 8.2516 (7.7781, 8.7454) 1.0362 (0.9909, 1.0823) 0.6318 (0.6090, 0.6554)
$\rho=0.8$ 0.2274 (0.1969, 0.2577) 0.1626 (0.1416, 0.1835) 0.7945 (0.7832, 0.8053) 0.7925 (0.7798, 0.8048)
$\nu=4$ NA 3.0049 (3.0000, 3.0063) NA 5.9993 (5.7126, 6.3156)
$\gamma=0.5$ NA NA 0.4964 (0.4918, 0.5010) 0.4985 (0.4954, 0.5016)
$\text{DIC}_{1}$ 30328.85 31535.42 14734.73 14732.85
$\text{DIC}_{2}$ 30328.34 31655.98 14737.32 14732.32
CT 0.1322 0.1206 0.6715 0.7252
Table S4: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 1 (full data), along with the computation time (CT) in seconds for one VB iteration. The table also presents the DIC values for each model. Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
Table S5 presents the posterior means and 95% credible intervals for selected model parameters obtained from the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models applied to the simulated dataset 2, along with the computational cost for one VB iteration. The table also includes the $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values for each model. Corresponding posterior density plots of the parameters are provided in Section S7.2.1. Unlike the simulated dataset 1, which is skewed and heavy tails, the simulated dataset 2 is more symmetrical with lighter tails. The posterior mean estimates for parameters $\beta_{0}$, $\beta_{1}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ across all four models closely match their true values. Likewise, the posterior mean estimate of $\nu$ for SEM-t and YJ-SEM-t is close to the true value of $30$, while the posterior mean estimate of $\gamma$ for YJ-SEM-Gau and YJ-SEM-t is close to the true value of $1$. The overall accuracy of parameter estimation across all models is further supported by the nearly identical $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values, with SEM-Gau having the lowest values, indicating the best fit for the simulated dataset 2, as expected.
According to the computation times presented in Tables S4 and S5, VB algorithms for both SEM-t and YJ-SEM-t, which incorporate Student’s $t$ errors, require significantly more computation time compared to SEM-Gau and YJ-SEM-Gau. This is because, in models with $t$ errors, the VB optimisation is substantially more complex than in models with Gaussian errors. Specifically, for models with Student’s $t$ errors, the set of model parameters is $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$, whereas for models with Gaussian errors, it simplifies to $\boldsymbol{\xi}=\boldsymbol{\phi}$. Here, $\boldsymbol{\tau}$ is a latent variable vector of length $n$; see Section 2.1 of the main paper.
SEM-Gau SEM-t YJ-SEM-Gau YJ-SEM-t
${\beta}_{0}=-3$2.9976 (-3.0258, -2.9692) -3.0008 (-3.0305, -2.9707) -2.9781 (-3.0108, -2.9449) -2.9808 (-3.0033, -2.9584)
${\beta}_{1}=2$ 1.9968 (1.9773, 2.0161) 1.9947 (1.9774, 2.0121) 1.9920 (1.9739, 2.0104) 1.9950 (1.9716, 2.0183)
$\sigma^{2}_{\textbf{e}}=0.5$ 0.5414 (0.5175, 0.5661) 0.4921 (0.4734, 0.5113) 0.5378 (0.5167, 0.5598) 0.4981 (0.4775, 0.5192)
$\rho=0.8$ 0.7987 (0.7880, 0.8091) 0.7970 (0.7856, 0.8080) 0.7984 (0.7873, 0.8091) 0.7977 (0.7857, 0.8092)
$\nu=30$ NA 22.7544 (21.8282, 23.7190) NA 26.1797 (24.8903, 27.5045)
$\gamma=1$ NA NA 1.0028 (0.9998, 1.0057) 1.0035 ( 0.9987, 1.0083)
$\text{DIC}_{1}$ 11315.17 11320.44 11475.1 11324.83
$\text{DIC}_{2}$ 11314.81 11321.4 11535.12 11324.85
CT 0.1280 0.6718 0.1197 0.7241
Table S5: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 2 (full data), along with the computation time (CT) in seconds for one VB iteration. The table also presents the DIC values for each model. Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
S7 Additional figures for simulation study in Section 4
This section provides additional figures related to the simulation study presented in Section 4 of the main paper, and Section S6.
S7.1 Comparison of VB and HVB methods
Section 4.1 of the main paper compares the estimated posterior densities of the model parameters for the YJ-SEM-Gau model using VB (including HVB) and HMC (Neal,, 1996) methods, both with and without missing data, based on a simulated dataset. To generate the data, the model parameters are set as follows: the fixed effects ($\boldsymbol{\beta}$) are randomly drawn from discrete uniform values between $-3$ and $3$ (excluding 00). The error variance parameter is set to $\sigma^{2}_{\mathbf{e}}=1$ and the spatial autoregressive parameter is $\rho=0.8$. The Yeo-Johnson transformation parameter is set to $\gamma=1.25$, which introduces moderate left skewness in the simulated response variable. The kernel density plot of the simulated response variable is provided in Figure S1. As expected, the density is moderately skewed to the left.
Figure S1: Kernel density plot of the simulated response variable generated using the YJ-SEM-Gau model with $\gamma=1.25$ in the simulation study of Section 4.1 of the main paper.
S7.2 Evaluating the performance of proposed SEMs
S7.2.1 Simulation study with full data
This section presents further figures from Section S6.
Figures S2 and S3 compare the posterior densities of selected model parameters obtained using the proposed VB method, applied to different SEMs (without missing values) for simulated dataset 1 and simulated dataset 2, respectively.
As shown in Figure S2, the YJ-SEM-t model is able to recover the true parameter values more accurately than the other three models, as expected, since the simulated dataset 1 was simulated using the YJ-SEM-t model. In contrast, as shown in Figure S3, all four models successfully recover the true values of the parameters.
Figure S2: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the simulated dataset 1, using the VB method. The vertical dotted lines indicate the true parameter values.
Figure S3: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the simulated dataset 2, using the VB method. The vertical dotted lines indicate the true parameter values.
S7.2.2 Simulation study with missing data
This section presents further figures from Section 4.2 of the main paper.
Figures S4 and S5 compare the posterior densities of selected model parameters estimated using the proposed HVB method, applied to different SEMs with missing values, for simulated dataset 1 and simulated dataset 2, respectively.
As shown in Figure S4, the YJ-SEM-t model is able to recover the true parameter values more accurately than the other three models, as expected, since the simulated dataset 1 was simulated using the YJ-SEM-t model. In contrast, as shown in Figure S5, all four models successfully recover the true values of the parameters.
Figure S4: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the simulated dataset 1 with missing data, using the HVB method. The vertical dotted lines indicate the true parameter values.
Figure S5: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the simulated dataset 2 with missing data, using the HVB method. The vertical dotted lines indicate the true parameter values.
The left panel of Figure S6 shows the kernel densities of the true missing values ($\mathbf{y}_{u}$) and the kernel density of posterior means of the missing values at locations $s_{1},s_{2},\ldots,s_{n}$, estimated from different SEMs for simulated dataset 2. All four models produce nearly identical densities of posterior means of missing values that closely match the density of the true missing values.
The right panel of Figure S6 presents the posterior density of the maximum missing value ($\textbf{y}_{u_{\textrm{max}}}$) predicted by each SEM. The true value of $\textbf{y}_{u_{\textrm{max}}}$ lies within the posterior distribution estimated by all models. These results are consistent with expectations, as all four models produce similar parameter estimates and DIC values (see Table 5).
Figure S6: Left panel: The kernel densities of the true missing values ($\textbf{y}_{u}$) and the posterior means of the missing values ($\textbf{y}_{u}$) obtained using the HVB-AllB method for different SEMs for the simulated dataset 2 with missing values.
Right panel: The posterior density of the maximum missing value ($\textbf{y}_{u_{\textrm{max}}}$) predicted by different SEMs for the simulated dataset 2 with missing values. The true value of the maximum missing value is indicated by the vertical line.
Figure S7 compares the posterior means and standard deviations of the missing values ($y_{u_{1}},\ldots,y_{u_{n_{u}}}$) estimated by different SEMs for the simulated dataset 1 with missing values. The left panel compares the posterior means of the predicted missing values with their true values. The right panel compares the posterior standard deviations of the missing values estimated by the YJ-SEM-t model (x-axis) against those obtained from SEM-Gau, SEM-t, and YJ-SEM-Gau models (y-axis).
The posterior means of the missing values estimated by the YJ-SEM-Gau and YJ-SEM-t models are closer to the true missing values compared to those estimated by the SEM-Gau and SEM-t models. Additionally, the posterior standard deviations from the YJ-SEM-Gau and YJ-SEM-t models are nearly identical, whereas those from the SEM-Gau and SEM-t models deviate from the standard deviations obtained by the YJ-based models.
Figure S7: Comparison of posterior means and standard deviations of missing values for the simulated dataset 1: The left plot shows the comparison of the posterior means of predicted missing values from the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t with true missing values. The right plot compares the posterior standard deviations of the missing values from SEM-Gau, SEM-t, and YJ-SEM-Gau against those obtained from YJ-SEM-t.
Figure S8 compares the posterior means and standard deviations of the missing values ($y_{u_{1}},\ldots,y_{u_{n_{u}}}$) estimated by different SEMs for the simulated dataset 2 with missing values. The left panel compares the posterior means of the predicted missing values with their true values. The right panel compares the posterior standard deviations of the missing values estimated by the YJ-SEM-t model (x-axis) against those obtained from SEM-Gau, SEM-t, and YJ-SEM-Gau models (y-axis).
The posterior means of the missing values estimated by all four models are close to the true values. Additionally, the posterior standard deviations from the SEM-Gau, SEM-t, and YJ-SEM-Gau models closely align with those from the YJ-SEM-t model, with YJ-SEM-Gau showing the closest match.
Figure S8: Comparison of posterior means and standard deviations of predicted missing values for the simulated dataset 2: The left plot shows the comparison of the posterior means of missing values from the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t with true missing values. The right plot compares the posterior standard deviations of the missing values from SEM-Gau, SEM-t, and YJ-SEM-Gau against those obtained from YJ-SEM-t.
S8 Additional figures for real data application in Section 5
This section provides additional figures related to the real data application discussed in Section 5 of the main paper.
Figure S9 shows kernel density plots of house prices from the Lucas-1998-HP dataset. The left panel displays the distribution of house prices (in hundreds of thousands), while the right panel shows the distribution of their natural logarithms. For a detailed description of the Lucas-1998-HP dataset, see Section 5 of the main paper. As expected, house prices being strictly positive, exhibit right skewness, whereas their logarithms appear to be left-skewed.
Figure S9: Kernel density plots of house prices from the Lucas-1998-HP data: The left panel displays the kernel density of house prices (in hundreds of thousands). The right panel shows the kernel density of the logarithm of house prices (in hundreds of thousands).
Figures S10 and S11 compare the posterior densities of selected model parameters, estimated using the proposed VB and HVB-AllB methods, across different SEMs (without and with missing values) applied to the Lucas-1998-HP dataset.
Figure S10: Posterior densities of some of the parameters of the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the Lucas-1998-HP dataset without missing values, using the VB method.
Figure S11: Posterior densities of some of the parameters of the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the Lucas-1998-HP dataset with missing values, using the HVB-AllB method.
S9 Convergence analysis
For the variational Bayes (VB) and hybrid VB (HVB) algorithms, convergence is assessed by examining the trajectories of the variational means of model parameters across iterations. For the Hamiltonian Monte Carlo (HMC) algorithm (Neal et al.,, 2011), convergence is evaluated by inspecting the trace plots of the posterior draws of model parameters.
Section S9.1 presents the convergence analysis plots for the simulation study described in Section 4 of the main paper. Similarly, Section S9.2 provides the convergence analysis plots for the real-data example discussed in Section 5 of the main paper.
S9.1 Convergence analysis of simulation study
S9.1.1 Convergence analysis for Section 4.1
This section presents convergence plots from the simulation study designed to compare the performance of the VB approximation with the HMC algorithm, as described in Section 4.1 of the main paper. In this study, data are simulated from the YJ-SEM-Gau model, and the model is estimated using both VB and HMC methods.
Figure S12 presents trace plots of posterior samples obtained using the HMC algorithm for selected parameters of the YJ-SEM-Gau model with full data, after excluding burn-in samples. Figure S13 displays the trajectories of the variational means for selected parameters across VB iterations. Both plots indicate that the algorithms have converged.
Figure S12: Trace plots of posterior samples obtained using HMC for selected parameters of the YJ-SEM-Gau model (without missing values), after excluding burn-in samples.
Figure S13: Trajectories of the variational means of selected parameters of the YJ-SEM-Gau model (without missing values) across VB iterations.
Next, missing values are introduced into the simulated dataset, and the YJ-SEM-Gau model is fitted using both the HVB-NoB and HMC algorithms.
Figure S14 presents trace plots of posterior samples obtained using the HMC algorithm for selected parameters of the YJ-SEM-Gau model with missing values, after excluding burn-in samples. Figure S15 displays the trajectories of the variational means for selected parameters across HVB-NoB iterations. Both plots indicate that the algorithms have converged.
Figure S14: Trace plots of posterior samples obtained using HMC for selected parameters of the YJ-SEM-Gau model (with missing values), after excluding burn-in samples.
Figure S15: Trajectories of the variational means of selected parameters of the YJ-SEM-Gau model (with missing values) across HVB-NoB iterations.
S9.1.2 Convergence analysis for Section 4.2
This section presents convergence plots from the simulation study conducted to evaluate the performance of the proposed SEMs, as described in Section 4.2 of the main paper.
Figure S16 shows trajectories of the variational means for selected model parameters of different SEMs fitted to the simulated dataset 1, without missing values. It can be observed that for each parameter in each model, the variational means converge by the end of the final iterations, indicating that the VB algorithms for all models have successfully converged.
Figure S16: Trajectories of the variational means of selected parameters across VB iterations for different SEMs fitted on the simulated dataset 1.
Figure S17 shows trajectories of the variational means for selected model parameters of different SEMs fitted to the simulated dataset 2, without missing values. It can be observed that for each parameter in each model, the variational means converge by the end of the final iterations, indicating that the VB algorithms for all models have successfully converged.
Figure S17: Trajectories of the variational means of selected parameters across VB iterations for different SEMs fitted on the simulated dataset 2.
Figure S18 shows trajectories of the variational means for selected model parameters of different SEMs fitted to the simulated dataset 1, with missing values. It can be observed that for each parameter in each model, the variational means converge by the end of the final iterations, indicating that the HVB-AllB algorithms for all models have successfully converged.
Figure S18: Trajectories of the variational means of selected parameters across HVB-AllB iterations for different SEMs fitted on the simulated dataset 1 with missing values.
Figure S19 shows trajectories of the variational means for selected model parameters of different SEMs fitted to the simulated dataset 2, with missing values. It can be observed that for each parameter in each model, the variational means converge by the end of the final iterations, indicating that the HVB-AllB algorithms for all models have successfully converged.
Figure S19: Trajectories of the variational means of selected parameters across HVB-AllB iterations for different SEMs fitted on the simulated dataset 2 with missing values.
S9.2 Convergence analysis of real example
Figure S20 shows trajectories of the variational means for selected model parameters of different SEMs fitted to the Lucas-1998-HP dataset (full data). It can be observed that for each parameter in each model, the variational means converge by the end of the final iterations, indicating that the HVB-AllB algorithms for all models have successfully converged.
Figure S20: Trajectories of the variational means of selected parameters across HVB-AllB iterations for different SEMs fitted on the Lucas-1998-HP dataset without missing values.
Figure S21 shows trajectories of the variational means for selected model parameters of different SEMs fitted to the Lucas-1998-HP dataset, with missing values. It can be observed that for each parameter in each model, the variational means converge by the end of the final iterations, indicating that the VB algorithms for all models have successfully converged.
Figure S21: Trajectories of the variational means of selected parameters across HVB-AllB iterations for different SEMs fitted on the Lucas-1998-HP dataset with missing values.

### v2

@@
-Non-Gaussian Simultaneous Autoregressive Models with Missing Data
+Bayesian Inference for Non-Gaussian Simultaneous Autoregressive Models with Missing Data
 Abstract
-Standard simultaneous autoregressive (SAR) models are usually assumed to have normally distributed errors, an assumption that is often violated in real-world datasets, which are frequently found to exhibit non-normal, skewed, and heavy-tailed characteristics.
+Standard simultaneous autoregressive (SAR) models typically assume normally distributed errors, an assumption often violated in real-world datasets that frequently exhibit non-normal, skewed, or heavy-tailed characteristics.
 New SAR models are proposed to capture these non-Gaussian features.
-In this project, the spatial error model (SEM), a widely used SAR-type model, is considered.
-Three novel SEMs are introduced that extend the standard Gaussian SEM by incorporating Student’s $t$-distributed errors after a one-to-one transformation is applied to the response variable.
-Variational Bayes (VB) estimation methods are developed for these models, and the framework is further extended to handle missing response data.
-Standard variational Bayes (VB) methods perform well with complete datasets; however, handling missing data requires a Hybrid VB (HVB) approach, which integrates a Markov chain Monte Carlo (MCMC) sampler to generate missing values.
-The proposed VB methods are evaluated using both simulated and real-world datasets, demonstrating their robustness and effectiveness in dealing with non-normal data and missing data in spatial models.
+The spatial error model (SEM), a widely used SAR-type model, is considered.
+Three novel SEMs are introduced, extending the standard Gaussian SEM.
+These extensions incorporate Student’s $t$-distributed errors to accommodate heavy-tailed behaviour, one-to-one transformations of the response variable to address skewness, or a combination of both.
+Variational Bayes (VB) estimation methods are developed for these models, and the framework is further extended to handle missing response data under the missing not at random (MNAR) mechanism.
+Standard VB methods perform well with complete datasets; however, handling missing data requires a hybrid VB (HVB) approach, which integrates a Markov chain Monte Carlo (MCMC) sampler to generate missing values.
+The proposed VB methods are evaluated using both simulated and real-world datasets, demonstrating their robustness and effectiveness in dealing with non-Gaussian data and missing data in spatial models.
 Although the method is demonstrated using SAR models, the proposed model specifications and estimation approaches are widely applicable to various types of models for handling non-Gaussian data with missing values.
-Keywords: Spatial error models; Student’s $t$ errors; Yeo and Johnson (YJ) transformation; Missing not at random; Variational Bayes approximation
+Keywords: Spatial error models; Student’s $t$ errors; Yeo and Johnson (YJ) transformation; Missing not at random; Variational Bayes
 1 Introduction
-Conventional spatial statistical models, such as simultaneous autoregressive (SAR) models, conditional autoregressive (CAR) models, and kriging models, among others (see Besag, (1974); Cressie, (1993); Anselin, (1988)), typically assume that the data being modeled follow a Gaussian distribution.
-However, in practice, this assumption is often violated.
-In this study, we propose novel SAR-type models and their estimation methods, which are particularly useful when the data are non-Gaussian.
-SAR models are a broad class of spatial statistical models commonly used for analysing spatially correlated lattice data (Cressie,, 1993).
-Examples of SAR models includes spatial error models (SEMs), spatial autoregressive models (SAMs), and spatial Durbin models (SDMs), and many more; see Anselin, (1988), and Cressie, (1993).
-These models account for spatial dependence, making them useful tools for understanding spatial phenomena in fields such as ecology (Ver Hoef et al.,, 2018), social sciences (Chi and Zhu,, 2008, 2019), finance (Calabrese et al.,, 2017; Agosto et al.,, 2019).
-The estimation of SAR models using methods such as the maximum likelihood estimation (MLE) (Ord,, 1975; LeSage,, 1997) and the Bayesian methods (De Oliveira and Song,, 2008; Wu,, 2018), under the assumption of Gaussian errors, has been extensively discussed in the literature.
-To address the issue of outliers in estimating SEM, Yildirim and Mert Kantar, (2020) proposed a robust version of the ML estimator.
-Estimation methods for SAR models that do not rely on strong distributional assumptions have been developed.
-For example, Lee, (2004) introduced the quasi-maximum likelihood estimator (QMLE) for the SAM.
-Furthermore, Generalised Method of Moments (GMM) estimators, which require fewer distributional assumptions than traditional ML estimators, are robust to unknown heteroskedasticity and deviations from normality.
-These estimators have been proposed by Kelejian and Prucha, (1998), Lee and Liu, (2010), and Breitung and Wigger, (2018).
-Non-parametric estimation approaches have also been discussed in the literature for modeling SEM with non-Gaussian data (Al-Momani et al.,, 2017).
-Missing values frequently occur in spatial datasets.
+Simultaneous autoregressive (SAR) models are a broad class of spatial statistical models commonly used for analysing spatially correlated lattice data, with applications spanning ecology, social sciences, and finance (Cressie,, 1993; Chi and Zhu,, 2008; Calabrese et al.,, 2017; Ver Hoef et al.,, 2018).
+Examples of SAR models include spatial error model (SEM), spatial autoregressive models (SAM), and spatial Durbin model (SDM), among others (Anselin,, 1988; Cressie,, 1993).
+Similar to other conventional spatial models, SAR models typically assume Gaussian-distributed data; see  Besag, (1974); Anselin, (1988); Cressie, (1993).
+However, this normality assumption is frequently violated in practice, as real-world datasets often exhibit non-normal characteristics such as skewness and heavy-tailed distributions.
+Several studies have examined the estimation of SAR models when the underlying error distributions are non-Gaussian, leading to methods that do not rely on strong distributional assumptions.
+For example, quasi-maximum likelihood estimators (QMLE) have been proposed for the SAM (Lee,, 2004), while various Generalised Method of Moments (GMM) approaches have been developed for different types of SAR models, which are robust to unknown heteroskedasticity and deviations from Gaussianity (Kelejian and Prucha,, 1998; Lee and Liu,, 2010; Breitung and Wigger,, 2018).
+However, these approaches require the response variable to be fully observed and therefore cannot accommodate missing data.
+Missing values frequently arise in spatial datasets.
 When estimating SAR models, ignoring missing response values can lead to inconsistency and bias (Wang and Lee,, 2013; Benedetti et al.,, 2020).
 Various estimation methods for SAR models with randomly missing data have been well developed (see LeSage and Pace, (2004); Wang and Lee, (2013); Suesse and Zammit-Mangion, (2017); Suesse, (2018); Wijayawardhana et al., (2024)).
-However, research on estimating SAR models with non-randomly missing data remains relatively limited.
-Flores-Lagunes and Schnier, (2012) introduced a GMM estimator, while Rabovič and Čížek, (2023) proposed a partial ML method for SEM.
-Furthermore, Doğan and Taşpinar, (2018) and Seya et al., (2021) used the Metropolis-Hastings (MH) algorithm, which becomes computationally expensive when dealing with a large number of observations.
-Variational Bayes (VB) methods have recently emerged as a faster alternative to Markov chain Monte Carlo (MCMC) for estimating complex statistical models (Chappell et al.,, 2008; Han et al.,, 2016; Ong et al.,, 2018; Gunawan et al.,, 2021, 2024).
-Several commonly used VB methods include mean-field variational Bayes (MFVB) (Ormerod and Wand,, 2010), integrated non-factorised variational Bayes inference (INFVB) (Han et al.,, 2013), and Gaussian variational approximation (Ong et al.,, 2018; Tan and Nott,, 2018).
-Although VB methods offer a more computationally efficient alternative to MCMC, their use in estimating SAR models has been relatively limited.
-Wu, (2018) applied two VB approaches, a hybrid MFVB method and an INFVB method, to estimate spatial autoregressive confused (SAC) and matrix exponential spatial specification (MESS) models.
-Similarly, Bansal et al., (2021) used MFVB and INFVB to estimate spatial count data models, incorporating a MESS model to capture spatial dependence in error terms.
-More recently, Wijayawardhana et al., (2025) introduced novel VB methods for estimating SAR models with missing data, demonstrating their effectiveness for moderately large datasets with a high proportion of missing values.
-Our article makes several key contributions.
+However, research on estimating SAR models when data are missing not at random (MNAR) remains relatively scarce.
+Flores-Lagunes and Schnier, (2012) proposed a GMM estimator for this setting, while Doğan and Taşpinar, (2018) and Seya et al., (2021) employed the Metropolis–Hastings (MH) algorithm.
+Bayesian estimation of SAR models using Markov chain Monte Carlo (MCMC) methods can be computationally demanding, particularly when the data contain missing values (Wijayawardhana et al.,, 2025).
+Variational Bayes (VB) provides a computationally efficient alternative, yet it remains underutilised in the context of SAR models.
+For instance, Wu, (2018) and Bansal et al., (2021) applied VB to fully observed SAR models, while Wijayawardhana et al., (2025) proposed hybrid VB (HVB) approaches for SAR models with missing data.
+Moreover, Bayesian inference methods, including both VB and MCMC, offer full posterior uncertainty quantification, in contrast to GMM, which yields only point estimates with asymptotic standard errors.
+Our article makes several key contributions aimed at addressing limitations in the current literature on SAR models, particularly in handling non-Gaussian responses, missing data, and the computational challenges arising from these complex scenarios.
 First, we introduce novel SAR models designed to capture non-Gaussian characteristics of the response variable, such as skewness and heavy tails.
-These models extend the standard Gaussian SAR by incorporating Student’s t-distributed errors and applying a one-to-one transformation, the Yeo-Johnson (YJ) transformation (Yeo and Johnson,, 2000), to the response variable (see Section 2.2 for details).
-By combining the properties of the Student’s $t$ distribution and the YJ transformation, these models effectively account for heavy-tailedness and skewness, making them well-suited for complex real-world data.
-Although we illustrate these new model specifications using SEMs, the framework is applicable to any SAR model.
-Second, we propose efficient VB methods for estimating these new SEMs, with and without missing values in the response variable.
-For the missing data case, we specifically consider the missing not at random (MNAR) mechanism, as defined by Rubin, (1976), where the probability of a missing value depends on both observed and unobserved responses.
-In standard VB inference, the joint posterior distribution of missing values and model parameters is approximated using a Gaussian variational approximation (Wijayawardhana et al.,, 2025).
-However, it has been shown that the Gaussian variational approximation with a factor covariance structure (Ong et al.,, 2018), is not sufficiently flexible to jointly approximate the posterior distributions of the SAR model parameters and missing values.
-Therefore, in this study, we use the standard Gaussian variational approximation of Ong et al., (2018) for SEMs without missing values, while SEMs with missing values are estimated using an extended version of the hybrid VB (HVB) algorithm of Wijayawardhana et al., (2025).
-The rest of this paper is organised as follows.
-Section 2 presents the new SEMs.
-In Section 3, we present the VB methods to estimate the SEMs with missing data.
-In Section 4, simulation studies are conducted to evaluate the performance of the new SEMs and VB methods.
-Section 5 applies the SEMs to a real-world dataset.
-Section 6 discusses our major results and findings.
-The paper also has an online supplement with additional technical details and examples.
+We extend conventional Gaussian SAR models in two ways: by incorporating Student’s $t$-distributed errors to address heavy-tailedness, and by applying the Yeo-Johnson (YJ) transformation (Yeo and Johnson,, 2000) to the response variable to handle skewness (see Section 2.2).
+Unlike the traditional two-step approach, where skewness is first reduced using a transformation such as YJ or Box-Cox and then a model is fitted, our integrated framework combines these steps within the SAR structure.
+These innovations yield three novel SAR models that jointly capture skewness and heavy-tailedness, making them well-suited for complex real-world data.
+While we demonstrate these models using SEM, yielding three SEM variants, the framework is broadly applicable to other SAR models.
+Second, we propose efficient VB methods for estimating these new SEMs, both with and without missing values in the response variable.
+For SEMs with missing values, we use an extended version of the HVB algorithm of Wijayawardhana et al., (2025) to handle non-Gaussian response variables and missing data under the MNAR mechanism.
+While the proposed HVB algorithm is an approximation compared to the MCMC methods (Doğan and Taşpinar,, 2018; Seya et al.,, 2021) and Hamiltonian Monte Carlo (HMC) (Neal,, 2011), it offers substantial computational advantages.
+Our simulation study in Section 6 shows that the HVB algorithm is more than ten times faster than the HMC implementation in RStan (Stan Development Team,, 2023) for a dataset of $n=625$ observations with approximately $50\%$ missing responses.
+Furthermore, Sections 7 and 8 show that the proposed HVB approach scales efficiently to much larger datasets, for which HMC becomes computationally impractical.
+The rest of the paper is organised as follows.
+Section 2 introduces the new SEMs, and Section 3 discusses SEMs with missing responses.
+In Section 4, we present the VB methods for estimating SEMs with missing data.
+Section 5 discusses Bayesian model comparison.
+Sections 6 and 7 present two simulation studies, and Section 8 applies the SEMs to a real-world dataset.
+Section 9 summarises the main findings and outlines directions for future work.
+The paper also has an online supplement with additional technical details.
+The R code to reproduce all simulations and real-world analyses is available at https://github.com/AnjanaWijayawardhana/Non-Gaussian-SAR-with-Missing-Data-VB.
 2 Simultaneous Autoregressive Models
-This section introduces our proposed SAR models, focusing on the spatial error model (SEM), a widely used SAR-type model. We present three novel SEMs that extend the standard Gaussian SEM.
-Let $\textbf{y}^{\ast}=(y^{\ast}_{1},y^{\ast}_{2},...,y^{\ast}_{n})^{\top}$ be the vector of response variable observed at $n$ spatial locations $s_{1},\ldots,s_{n}$, X be the $n\times(r+1)$ design matrix containing the covariates, and W be the $n\times n$ spatial weight matrix.
+In this section, we use our proposed framework to extend conventional Gaussian SAR models to non-Gaussian settings.
+We showcase the extension using spatial error models (SEMs), a widely adopted form of SAR-type models.
+Let $\textbf{y}^{\ast}=(y^{\ast}_{1},y^{\ast}_{2},...,y^{\ast}_{n})^{\top}$ be the $n\times 1$ vector of response variable observed at $n$ spatial locations $\textbf{s}_{1},\ldots,\textbf{s}_{n}$, X be the $n\times(r+1)$ design matrix containing the covariates, and W be the $n\times n$ spatial weight matrix.
 The SEM is given by
 $\textbf{y}^{\ast}=\textbf{X}\boldsymbol{\beta}+(\textbf{I}_{n}-\rho\textbf{W})^{-1}\textbf{e},$ (1)
-where e is an $n\times 1$ vector of error terms, $\boldsymbol{\beta}=(\beta_{0},\beta_{1},\ldots,\beta_{r})^{\top}$ is an $(r+1)\times 1$ vector of fixed effects parameters, $\rho$ is the spatial autocorrelation parameter, which measures the strength and direction of spatial dependence (LeSage and Pace,, 2009), and $\textbf{I}_{n}$ is the $n\times n$ identity matrix.
+where e is the $n\times 1$ vector of error terms, $\boldsymbol{\beta}=(\beta_{0},\beta_{1},\ldots,\beta_{r})^{\top}$ is the $(r+1)\times 1$ vector of fixed effects parameters, $\rho$ is the spatial autocorrelation parameter, which measures the strength and direction of spatial dependence (LeSage and Pace,, 2009), and $\textbf{I}_{n}$ is the $n\times n$ identity matrix.
 Section 2.1 discusses SEMs with Gaussian and Student’s $t$ errors.
-Section 2.2 discusses two SEMs with Yeo-Johnson (Yeo and Johnson,, 2000) transformation.
-Section 2.3 discusses SEMs with missing data.
-2.1 Spatial error models with Gaussian and Student’s $t$ errors
-The standard SEM assumes Gaussian errors $\textbf{e}\sim N(0,\sigma_{\textbf{e}}^{2}\textbf{I}_{n})$, where $\sigma_{\textbf{e}}^{2}$ is the variance parameter.
-This assumption implies that the distribution of $\textbf{y}^{\ast}$ is also Gaussian, with the mean vector and the covariance matrix given in Table 1.
+Section 2.2 discusses two SEMs with the Yeo-Johnson transformation (Yeo and Johnson,, 2000).
+2.1 SEMs with Gaussian and Student’s $t$ errors
+The conventional SEM assumes Gaussian error terms $\textbf{e}\sim N(\textbf{0},\sigma_{\textbf{e}}^{2}\textbf{I}_{n})$ in the model presented in Equation (1), where $\sigma_{\textbf{e}}^{2}$ is the error variance parameter.
+This assumption implies that the distribution of $\textbf{y}^{\ast}$ is also Gaussian, with the mean vector $\textbf{X}\boldsymbol{\beta}$ and covariance matrix $\sigma^{2}_{\textbf{e}}(\textbf{A}^{\top}\textbf{A})^{-1}$, where $\textbf{A}=\textbf{I}_{n}-\rho\textbf{W}$.
 We call this model the SEM-Gau.
+We now discuss SEM-t, which assumes that the error terms e in Equation (1) follow a Student’s $t$-distribution.
 Student’s $t$-distribution has been widely used to model heavy-tailed data in popular statistical models, such as regression (Lange et al.,, 1989; Fernández and Steel,, 1999) and mixed-effects models (Pinheiro et al.,, 2001; Wang and Lin,, 2014).
 Huang et al., (2021) applied Student’s $t$ errors to spatial autoregressive scalar-on-function regression models.
-We now discuss SEM-t, which replaces the Gaussian errors in the SEM-Gau model with a Student’s t-distribution.
-Let $e_{i}\sim t_{\nu}(0,\sigma^{2}_{e})$, where $e_{i}$ is the $i^{th}$ element of the error vector e in Equation (1) and $t_{\nu}(0,\sigma^{2}_{e})$ is the Student’s t distribution with $\nu$ degrees of freedom ($\nu\geq 3$), mean zero, and scale parameter $\sigma^{2}_{e}$.
-We write Student’s t errors, $e_{i}$, $i=1,...,n$, as a scale mixture of normals (Chan et al.,, 2019),
+Let $e_{i}\sim t_{\nu}(0,\sigma^{2}_{e})$, where $e_{i}$ is the $i^{\textrm{th}}$ element of the error vector e in Equation (1) and $t_{\nu}(0,\sigma^{2}_{e})$ is the Student’s $t$ distribution with $\nu$ degrees of freedom ($\nu\geq 3$), mean zero, and scale parameter $\sigma^{2}_{e}$.
+These Student’s $t$ errors, $e_{i}$ for $i=1,\ldots,n$, can be expressed as a scale mixture of normals (Chan et al.,, 2019):
 $\begin{split}e_{i}&\mid\tau_{i},\sigma^{2}_{e}~\sim N(0,\sigma^{2}_{e}\tau_{i}),\\ \tau_{i}&\mid\nu\sim IG\left(\frac{\nu}{2},\frac{\nu}{2}\right),~~~~i=1,\ldots n\end{split}$ (2)
-where $IG(a,b)$ denotes the inverse gamma distribution with shape parameter $a$ and scale parameter $b$.
+where $IG(a,b)$ denotes the inverse gamma distribution with shape parameter $a$ and rate parameter $b$.
 The conditional distribution of errors presented in Equation (2), can be written in vector notation as
@@
 We now incorporate the error structure from Equation (3) into the SEM described in Equation (1).
-As a result, the distribution of $\textbf{y}^{\ast}$ given the latent vector $\boldsymbol{\tau}$ is also Gaussian, with the mean vector and the covariance matrix given in Table 1.
-We refer to this model as SEM-t.
-The log-likelihood function of $\textbf{y}^{\ast}$ for SEM-Gau and SEM-t is expressed as:
+As a result, the distribution of $\textbf{y}^{\ast}$ conditional on the latent vector $\boldsymbol{\tau}$ is also Gaussian, with the mean vector $\textbf{X}\boldsymbol{\beta}$ and covariance matrix $\sigma^{2}_{\textbf{e}}(\textbf{A}^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A})^{-1}$, where $\textbf{A}=\textbf{I}_{n}-\rho\textbf{W}$.
+This model is referred to as the SEM-t.
+The log-likelihood function of $\textbf{y}^{\ast}$ for both the SEM-Gau and SEM-t is given by:
 $\text{log}~p_{\textbf{y}^{\ast}}(\textbf{y}^{\ast}\mid\boldsymbol{\xi})=-\frac{n}{2}\textrm{log}(2\pi)-\frac{n}{2}\textrm{log}(\sigma^{2}_{\boldsymbol{e}})+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{1}{2\sigma^{2}_{\boldsymbol{e}}}\textbf{r}^{\top}\textbf{M}\textbf{r},$ (4)
-where for the SEM-Gau model, $\boldsymbol{\xi}=\boldsymbol{\phi}$, while for the SEM-t model, $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$.
+where for the SEM-Gau, $\boldsymbol{\xi}=\boldsymbol{\phi}$, while for the SEM-t, $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$.
 The definitions for r, M, and $\boldsymbol{\phi}$ are given in Table 1.
-2.2 Spatial error models with transformations
-This section extends the SEM-Gau and SEM-t defined in Section 2.1.
-We form a novel class of SEMs using an element-wise Yeo and Johnson (YJ) transformation (Yeo and Johnson,, 2000) for the response variable, $\textbf{y}^{\ast}$, in both the SEM-Gau and SEM-t models.
-We assume that the $i^{th}$ observed response variable $(y_{i})$ is modeled as,
-${y}_{i}=t_{{\gamma}}^{-1}(y^{\ast}_{i}),$
-where $t_{{\gamma}}^{-1}(.)$ is the inverse function of the YJ transformation. For $20<\gamma<$ 2, it is defined as:
-$y_{i}=t_{{\gamma}}^{-1}(y^{\ast}_{i})=\begin{cases}(y^{\ast}_{i}\gamma+1)^{1/\gamma}-1&\text{if }y^{\ast}_{i}\geq 0\\ 1-(-(2-\gamma)y^{\ast}_{i}+1)^{1/(2-\gamma)}&\text{if }y^{\ast}_{i}<0.\\ \end{cases}$ (5)
+Table 1: The definitions of r, M, and the parameters $\boldsymbol{\phi}$ for SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t.
+For SEM-Gau and SEM-t, the density of the response vector conditional on $\boldsymbol{\xi}$ follows a multivariate Gaussian distribution with mean $\textbf{X}\boldsymbol{\beta}$ and covariance matrix $\sigma^{2}_{\textbf{e}}\textbf{M}^{-1}$, where $\textbf{A}=\textbf{I}_{n}-\rho\textbf{W}$.
+Model $\boldsymbol{\phi}$ r M
+SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $\textbf{A}^{\top}\textbf{A}$
+SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $\textbf{A}^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$
+YJ-SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e},},\rho,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $\textbf{A}^{\top}\textbf{A}$
+YJ-SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $\textbf{A}^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$
+2.2 SEMs with transformations
+This section extends the SEM-Gau and SEM-t in Equation (1) by applying the YJ transformation element-wise to the response vector $\textbf{y}^{\ast}$.
+Let $y_{i}$ be the $i^{\textrm{th}}$ element of the response vector in the SEM with YJ transformation. We assume that $y_{i}$ is modelled as
+${y}_{i}=t_{{\gamma}}^{-1}(y^{\ast}_{i}),\quad i=1,...,n,$ (5)
+where $y^{\ast}_{i}$ is the $i^{\textrm{th}}$ element of the response vector in the SEM given in Equation (1), with either Gaussian or Student’s $t$-distributed errors, and $t_{{\gamma}}^{-1}(\cdot)$ is the inverse function of the YJ transformation.
+For $0<\gamma<2$, it is defined as:
+[…]
 As the YJ transformation makes the data more symmetric and less skewed, applying its inverse transformation moves the data away from symmetry.
 This implies that $\textbf{y}=t_{{\gamma}}^{-1}(\textbf{y}^{\ast})=(t^{-1}_{{\gamma}}(y^{\ast}_{1}),t^{-1}_{{\gamma}}(y^{\ast}_{2}),\dots,t^{-1}_{{\gamma}}(y^{\ast}_{n}))^{\top}$ becomes an asymmetric, non-Gaussian response variable.
-The SEM with the YJ transformation (YJ-SEM) is defined as:
-$\textbf{y}=t_{{\gamma}}^{-1}(\textbf{y}^{\ast})=t_{{\gamma}}^{-1}(\textbf{X}\boldsymbol{\beta}+(\textbf{I}_{n}-\rho\textbf{W})^{-1}\textbf{e}).\$ (6)
-The density of y for the YJ-SEM is:
+In vector notation, the SEM with the YJ transformation (YJ-SEM) is defined as:
+$\textbf{y}=t_{{\gamma}}^{-1}(\textbf{y}^{\ast})=t_{{\gamma}}^{-1}(\textbf{X}\boldsymbol{\beta}+(\textbf{I}_{n}-\rho\textbf{W})^{-1}\textbf{e}).\\$ (7)
+Using the standard technique for transforming random variables, it is straightforward to show that the density of y for the YJ-SEM is:
 $p_{\textbf{y}}{(\textbf{y}\mid\boldsymbol{\xi})}=p_{\textbf{y}^{\ast}}({t_{\boldsymbol{\gamma}}(\textbf{y})})\prod_{i=1}^{n}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right),~~~~i=1,\ldots n,$ (8)
-where $p_{\textbf{y}^{\ast}}(\cdot)$ is the density of $\textbf{y}^{\ast}$; see Section 2.1, and $\frac{dt_{\gamma}(y_{i})}{dy_{i}}$ is the derivative of the YJ transformation with respect to $y_{i}$; see Section S1 of the online supplement.
-The YJ-SEM with normally distributed errors (denoted as YJ-SEM-Gau) assumes that the error term in Equation (6) follows a normal distribution: $\mathbf{e}\sim N(0,\sigma_{\mathbf{e}}^{2}\mathbf{I}_{n})$.
-The YJ-SEM with Student’s t-distributed errors (denoted as YJ-SEM-t) assumes that the error term follows the structure specified in Equation (2).
-The log-likelihood function for y in YJ-SEM-Gau and YJ-SEM-t is:
+[…]
+The YJ-SEM with Gaussian errors (YJ-SEM-Gau) assumes that the error term in Equation (7) follows a normal distribution: $\textbf{e}\sim N(0,\sigma_{\textbf{e}}^{2}\textbf{I}_{n})$, where $\sigma_{\textbf{e}}^{2}$ is a variance parameter.
+The YJ-SEM with Student’s $t$-distributed errors (YJ-SEM-t) assumes that the error term follows the structure given in Equation (3).
+The log-likelihood function of y for YJ-SEM-Gau and YJ-SEM-t is expressed as:
 $\begin{split}\text{log}~p_{\textbf{y}}(\textbf{y}\mid\boldsymbol{\xi})&=-\frac{n}{2}\textrm{log}(2\pi)-\frac{n}{2}\textrm{log}(\sigma^{2}_{\boldsymbol{e}})+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{1}{2\sigma^{2}_{\boldsymbol{e}}}\textbf{r}^{\top}\textbf{M}\textbf{r}\\ &+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right),\end{split}$ (9)
-where for the YJ-SEM-Gau model, $\boldsymbol{\xi}=\boldsymbol{\phi}$, while for the YJ-SEM-t model, $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$.
+where for the YJ-SEM-Gau, $\boldsymbol{\xi}=\boldsymbol{\phi}$, while for the YJ-SEM-t, $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$.
 The definitions for r, M, and $\boldsymbol{\phi}$ are given in Table 1.
-Note that $y_{i}=y^{\ast}_{i}$ for $i=1,...,n$, corresponds to an identity transformation in both SEM-Gau and SEM-t.
-The three SEMs, SEM-Gau, SEM-t, and YJ-SEM-Gau, are special cases of the YJ-SEM-t model.
-The SEM-Gau model is a YJ-SEM-t model with the degrees of freedom parameter ($\nu$) set to $\infty$ and the YJ parameter ($\gamma$) set to 1, while the YJ-SEM-Gau model is a YJ-SEM-t model with $\nu=\infty$ and $\gamma\neq 1$.
-Finally, the SEM-t model is also a YJ-SEM-t model with $\gamma=1$.
-Table 1: The definitions of r, M, and the parameters $\boldsymbol{\phi}$ for the four spatial error models.
-Each SEM has a mean vector of $\mathbf{X}\boldsymbol{\beta}$ and a covariance matrix of the form $\boldsymbol{\Sigma}=\sigma^{2}_{\textbf{e}}\textbf{M}^{-1}$, with $\textbf{A}=\textbf{I}_{n}-\rho\textbf{W}$.
-Model $\boldsymbol{\phi}$ r M
-SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $^{\top}\textbf{A}$ A
-SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$ A
-YJ-SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e},},\rho,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $^{\top}\textbf{A}$ A
-YJ-SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$ A
-2.3 Spatial error models with missing responses
+Note that $y_{i}=y^{\ast}_{i}$ for i=1,,ni=1,...,n, corresponds to an identity transformation in both SEM-Gau and SEM-t.
+The three SEMs, SEM-Gau, SEM-t, and YJ-SEM-Gau, are special cases of the YJ-SEM-t.
+The SEM-Gau is a YJ-SEM-t where the degrees of freedom parameter ($\nu$) equals $\infty$ and the YJ parameter ($\gamma$) equals 1, while the YJ-SEM-Gau is a YJ-SEM-t with $\nu=\infty$ and $\gamma\neq 1$.
+Finally, the SEM-t is also a YJ-SEM-t with $\gamma=1$.
+3 SEMs with missing responses
+This section develops the formulation of SEMs in settings where response values are missing and introduces the underlying missing data mechanisms, focusing in particular on the missing not at random (MNAR) mechanism.
 Consider that the response vector y of an SEM presented in Sections 2.1 and 2.2 contains missing values.
-Let $\textbf{y}_{o}$ be the subset of y with $n_{o}$ observed units, and $\textbf{y}_{u}$ be the subset of y with $n_{u}$ unobserved units.
-The complete response vector is $\textbf{y}=({\textbf{y}}_{o}^{\top},{\textbf{y}}_{u}^{\top})^{\top}$.
-A missing data indicator vector m of length ${n}$ containing 1’s and 0’s is defined.
-If an element in y is missing, then the corresponding element in m is 1 and 0, otherwise.
-In the presence of missing values, the vector r and the matrices X and M are decomposed into distinct components as follows:
+Let $\textbf{y}_{o}$ denote the subset of y with $n_{o}$ observed units, and $\textbf{y}_{u}$ denote the subset with $n_{u}$ unobserved (missing) units.
+The complete response vector can then be written as $\textbf{y}=(\textbf{y}_{o}^{\top},\textbf{y}_{u}^{\top})^{\top}$.
+We define a binary missingness indicator vector $\textbf{m}=(m_{1},\ldots,m_{n})^{\top}$, where
+$m_{i}=\begin{cases}1,&\text{if }y_{i}\text{ is missing,}\\ 0,&\text{if }y_{i}\text{ is observed.}\end{cases}$
+In the presence of missing responses, the vector r and the matrices X and M, as defined in Table 1, are partitioned into corresponding observed and unobserved components as follows:
 $\textbf{r}=\begin{pmatrix}\textbf{r}_{o}\\ {\textbf{r}_{u}}\end{pmatrix},~\textbf{X}=\begin{pmatrix}\textbf{X}_{o}\\ {\textbf{X}_{u}}\end{pmatrix},~\textbf{M}=\begin{pmatrix}\textbf{M}_{oo}&\textbf{M}_{ou}\\ \textbf{M}_{uo}&\textbf{M}_{uu}\end{pmatrix}$, (10)
-where $\textbf{r}_{o}$ and $\textbf{r}_{u}$ are subvectors of r, corresponding to the observed and unobserved responses, respectively.
-Similarly, $\textbf{X}_{o}$ and $\textbf{X}_{u}$ represent the design matrices associated with the observed and unobserved responses.
-Additionally, the matrices $\textbf{M}_{oo}$, $\textbf{M}_{ou}$, $\textbf{M}_{uo}$, and $\textbf{M}_{uu}$ are submatrices of M.
-We assume missing responses follow a missing not at random (MNAR) mechanism (Rubin,, 1976).
-The missing data mechanism is defined by the conditional distribution of m given y, commonly referred to as the missing data model, denoted by $p(\textbf{m}|\textbf{y},\boldsymbol{\psi},\textbf{X}^{*})$, where $\textbf{X}^{*}$ is an $n\times(q+1)$ design matrix containing the covariates of the missing data model.
-These covariates may be a subset of those used in SEMs.
-Assuming the conditional independence of m given y, $\textbf{X}^{*}$, and $\boldsymbol{\psi}$, the density is the product of $p(m_{i}\mid y_{i},\textbf{x}^{*}_{i},\boldsymbol{\psi})$ for $i=1,\ldots,n$, where $m_{i}$ and $y_{i}$ are the $i^{th}$ elements of m and y, and ${\textbf{x}_{i}^{*}}^{\top}$ is the $i^{th}$ row of $\textbf{X}^{*}$.
-The parameter vector $\boldsymbol{\psi}=(\boldsymbol{\psi}_{\textbf{x}}^{\top},\psi_{\textbf{y}})^{\top}$ consists of the fixed effects vector associated with covariates $\textbf{X}^{*}$, denoted as $\boldsymbol{\psi}_{\textbf{x}}=(\psi_{0},\psi_{1},\psi_{2},\ldots,\psi_{q})^{\top}$, and the fixed effect corresponding to y, denoted as $\psi_{\textbf{y}}$.
-A logistic regression model is used as the missing data model and is given by:
-$p(\textbf{m}\mid\textbf{y},\textbf{X}^{*},\boldsymbol{\psi})=\prod_{i=1}^{n}\frac{e^{({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\psi_{{y}}){m_{i}}}}{1+e^{{\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+\textbf{y}_{i}\psi_{\textbf{y}}}}.$ (10)
+where $\textbf{r}_{o}$ and $\textbf{r}_{u}$ are subvectors of r corresponding to the observed and unobserved responses, respectively.
+Similarly, $\textbf{X}_{o}$ and $\textbf{X}_{u}$ represent the design matrices associated with the observed and unobserved responses, while $\textbf{M}_{oo}$, $\textbf{M}_{ou}$, $\textbf{M}_{uo}$, and $\textbf{M}_{uu}$ denote the corresponding submatrices of M.
+3.1 Missing data mechanism and MNAR assumption
+The missing data mechanism is characterised by the joint probability mass function of the missingness indicator vector $\textbf{m}=(m_{1},\ldots,m_{n})^{\top}$, where each element $m_{i}$ denotes whether the corresponding response $y_{i}$ is missing ($m_{i}=1$) or observed ($m_{i}=0$).
+Let $p(m_{i}\mid y_{i},\textbf{x}_{i}^{*},\boldsymbol{\psi})$ denote the probability that $y_{i}$ is missing, possibly depending on itself and on a set of covariates $\textbf{x}_{i}^{*}=(1,x_{i1},\ldots,x_{iq})^{\top}$.
+The parameter vector $\boldsymbol{\psi}=(\boldsymbol{\psi}_{\textbf{x}}^{\top},\psi_{\textbf{y}})^{\top}$ consists of the fixed effect parameter vector associated with covariates $\textbf{x}^{*}_{i}$, denoted as $\boldsymbol{\psi}_{\textbf{x}}=(\psi_{0},\psi_{1},\psi_{2},\ldots,\psi_{q})^{\top}$, and the fixed effect parameter corresponding to y, denoted as $\psi_{\textbf{y}}$.
+Three main types of missing-data mechanisms are commonly distinguished following Rubin, (1976): (i) missing completely at random (MCAR), where the probability of missingness is independent of both the observed and unobserved data and on observed covariates, that is, $p(m_{i}\mid y_{i},\mathbf{x}_{i}^{*},\boldsymbol{\psi})=p(m_{i}\mid\boldsymbol{\psi})$; (ii) missing at random (MAR), where the probability of missingness may depend on observed covariates but not on the observed and unobserved responses, so that $p(m_{i}\mid y_{i},\mathbf{x}_{i}^{*},\boldsymbol{\psi})=p(m_{i}\mid\mathbf{x}_{i}^{*},\boldsymbol{\psi})$; and (iii) missing not at random (MNAR), where the probability of missingness depends on the response variable itself and observed covariates.
+Under MNAR, the term $p(m_{i}\mid y_{i},\mathbf{x}_{i}^{*},\boldsymbol{\psi})$ cannot be simplified further.
+The conditional probability mass function of $m_{i}$ under MNAR, $p(m_{i}\mid y_{i},\textbf{x}_{i}^{*},\boldsymbol{\psi})$ can be modelled using any binary-response model, such as a logistic or probit regression.
+In this paper, we model the probability that the $i^{\text{th}}$ response is missing via a logistic regression:
+$p(m_{i}=1\mid y_{i},\textbf{x}_{i}^{*},\boldsymbol{\psi})=\frac{\exp(\textbf{x}_{i}^{*\top}\boldsymbol{\psi}_{\textbf{x}}+y_{i}\psi_{y})}{1+\exp(\textbf{x}_{i}^{*\top}\boldsymbol{\psi}_{\textbf{x}}+y_{i}\psi_{y})},$ (11)
+so that the joint probability mass function of vector m is
+$p(\textbf{m}\mid\textbf{y},\textbf{X}^{*},\boldsymbol{\psi})=\prod_{i=1}^{n}\frac{\exp\{(\textbf{x}_{i}^{*\top}\boldsymbol{\psi}_{\textbf{x}}+y_{i}\psi_{y})m_{i}\}}{1+\exp(\textbf{x}_{i}^{*\top}\boldsymbol{\psi}_{\textbf{x}}+y_{i}\psi_{y})},$ (12)
+where $\textbf{X}^{\ast}=(\textbf{x}^{\ast\top}_{1},\dots,\textbf{x}^{\ast\top}_{n})^{\top}$ denotes the design matrix for the missingness model, with dimension $n\times(q+1)$.
+3.2 Joint modelling of y and m
+In our study, the primary objective is the estimation of the parameters of the proposed SEMs, denoted by $\boldsymbol{\xi}$, which govern the distribution of the response vector y.
+However, under the MNAR mechanism, the missingness process is not ignorable, and valid inference requires modelling the joint distribution of $(\textbf{y},\textbf{m})$ or, equivalently, the joint estimation of $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$ (Wijayawardhana et al.,, 2025).
+The joint distribution of y and m is denoted by $p(\textbf{y},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})$.
+Using the selection model factorisation (Little and Rubin,, 2019), $p(\textbf{y},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})$ can be factorised as
+$p(\textbf{y},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})=p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\textbf{y}\mid\boldsymbol{\xi}),$ (13)
+where $p(\textbf{y}\mid\boldsymbol{\xi})$ represents the likelihood of the SEM, and $p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})$ denotes the joint probability mass function of missingness indicators defined in Equation (12).
+The log-likelihood of $\mathbf{y}$ for the different SEMs, along with their full model specifications, is provided in Section 2.
+It is important to note that the factorisation in Equation (13) is a standard and widely applicable decomposition for jointly modelling a missingness mechanism and a statistical model with missing data, and it applies to any SAR-type model; see Section S4 of the online supplement.
+This factorisation enables practical inference for SAR models with missing values under MNAR.
+In Section 4, we discuss the estimation of the joint posterior distribution of the model parameters ($\boldsymbol{\xi},\boldsymbol{\psi}$) and $\textbf{y}_{u}$ using the proposed variational Bayes method.
 4 Variational Bayes inference for SEMs with missing data
-Section 3.1 introduces the variational Bayes (VB) inference method.
-Section 3.2 discusses the efficient hybrid variational Bayes (HVB) methods for estimating the SEMs with missing data.
-Section 3.3 discusses the Deviance Information Criterion (DIC) to select the best SEMs.
-The VB method for the proposed SEMs in Section 2, when there are no missing data, is provided in Section S2 of the online supplement.
-3.1 Variational Bayes inference
-Variational Bayes (VB) inference approximates a Bayesian posterior distribution by formulating it as an optimisation problem, offering a more computationally efficient alternative to the computationally intensive Markov chain Monte Carlo (MCMC) methods for complex statistical models.
-This section presents the VB method for SEMs with missing data.
-We assume that some values in the response vector of an SEM are missing under the missing not at random (MNAR) mechanism.
-Under this assumption, Bayesian estimation of the SEM requires consideration of the joint distribution of the missing data indicator vector m and the response vector y (Wijayawardhana et al.,, 2025), which is denoted by $p(\textbf{y},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})$.
-Using the selection model factorisation (Little and Rubin,, 2019), $p(\textbf{y},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})$ is expressed as
-$p(\textbf{y},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})=p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\textbf{y}\mid\boldsymbol{\xi}).$ (11)
-where $p(\textbf{y}\mid\boldsymbol{\xi})$ is the likelihood of y.
-The log-likelihood for y for different SEMs is given in Section 2.
-Additionally, $p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})$ denotes the density of the missing data model, given in Equation (10).
-Consider Bayesian inference to estimate the joint posterior of $\boldsymbol{\xi}$, $\boldsymbol{\psi}$, and the missing values $\textbf{y}_{u}$.
-For the SEM-Gau and YJ-SEM-Gau models, $\boldsymbol{\xi}=\boldsymbol{\phi}$, while for the SEM-t and YJ-SEM-t models, $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$; see Table 1 for the definition of $\boldsymbol{\phi}$ for each SEM.
+Section 4.1 provides an overview of the variational Bayes (VB) inference framework for estimating SEMs with missing responses.
+Section 4.2 then introduces an efficient hybrid variational Bayes (HVB) algorithm.
+For completeness, Section S2 of the online supplement presents the VB algorithm for SEMs with fully observed data.
+4.1 Variational Bayes inference under missing responses
+Variational Bayes (VB) inference approximates a posterior distribution by formulating it as an optimisation problem, offering a more computationally efficient alternative to the computationally intensive Markov chain Monte Carlo (MCMC) methods for complex statistical models.
+The standard VB approach for SEMs with missing responses is described below.
+We assume that some values in the response vector of an SEM are missing under the MNAR mechanism, as described in Section 3.
+Under this assumption, Bayesian inference requires estimating the joint posterior distribution of the model parameters $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$, based on the joint likelihood defined in Equation (13).
+Since the subvector $\textbf{y}_{u}$ of $\textbf{y}=(\textbf{y}_{o}^{\top},\textbf{y}_{u}^{\top})^{\top}$ is unobserved, it must also be inferred.
+Consequently, Bayesian inference involves estimating the joint posterior distribution of $\boldsymbol{\xi}$, $\boldsymbol{\psi}$, and $\textbf{y}_{u}$.
+For the SEM-Gau and YJ-SEM-Gau, $\boldsymbol{\xi}=\boldsymbol{\phi}$, while for the SEM-t and YJ-SEM-t, $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$; see Table 1 for the definition of $\boldsymbol{\phi}$ for each SEM.
 Let the joint posterior distribution of $\boldsymbol{\xi}$, $\boldsymbol{\psi}$, and $\textbf{y}_{u}$ be denoted as $p(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m})$.
-We let $p(\boldsymbol{\xi})$ and $p(\boldsymbol{\psi})$ as the prior distributions of $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$, respectively.
+We denote the prior distributions of $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$ by $p(\boldsymbol{\xi})$ and $p(\boldsymbol{\psi})$, respectively.
 Using the selection model factorisation in Equation (13), the joint posterior distribution $p(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m})$} is given by
@@
 We define $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\xi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\xi})p(\boldsymbol{\psi})$.
-The term $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ for each SEM with missing data, along with the total number of parameters to be estimated ($s_{m}$) for each model are provided in Table 2.
-Table 2: The term $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\xi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\xi})p(\boldsymbol{\psi})$, and the total number of model parameters $s_{m}$ (including the length of the latent vector $\boldsymbol{\tau}$) for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models with missing values.
-The distribution $p(\boldsymbol{\tau}\mid\boldsymbol{\phi})=\prod_{i=1}^{n}p(\tau_{i}\mid\nu)$ represents the $n$ independent inverse gamma latent variables used in SEMs with Student-$t$ errors (see Equation (2)).
-Model $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ $s_{m}$
+The term $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ for each SEM with missing responses, along with the total number of parameters to be estimated ($s$) for each model, are provided in Table 2.
+Table 2: The term $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\xi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\xi})p(\boldsymbol{\psi})$, and the total number of model parameters $s$ (including the length of the latent vector $\boldsymbol{\tau}$ if exists) for the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t with missing responses.
+The density $p(\boldsymbol{\tau}\mid\boldsymbol{\phi})=\prod_{i=1}^{n}p(\tau_{i}\mid\nu)$} represents the $n$ independent inverse gamma latent variables used in SEMs with Student-$t$ errors (see Equation (2)).
+Model $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ $s$
 SEM-Gau $p({\textbf{y}}\mid\boldsymbol{\phi})p(\textbf{m}\mid{\textbf{y}},\boldsymbol{\psi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$ $r+q+5$
@@
 YJ-SEM-t $p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\textbf{m}\mid{\textbf{y}},\boldsymbol{\psi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$ $r+q+7+n$
+We consider the variational approximation $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$, indexed by the variational parameter $\boldsymbol{\lambda}$ to approximate the joint posterior $p(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m})$ in Equation (14).
+The VB approach approximates this posterior distribution by minimising the Kullback-Leibler (KL) divergence between $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ and $p(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m})$, which is defined as
 […]
-$\mathcal{L}(\boldsymbol{\lambda})=E_{q}\left[\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})-\text{log}~{}q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})\right],$ (13)
-where $E_{q}\left[\cdot\right]$ denotes the expectation with respect to $q_{\boldsymbol{\lambda}}$.
+Minimising the KL divergence in Equation (S5) is equivalent to maximising the evidence lower bound (ELBO) on the logarithm of the marginal likelihood, $\text{log}~p(\textbf{y}_{o},\textbf{m})$, denoted by $\mathcal{L}(\boldsymbol{\lambda})$, where $p(\textbf{y}_{o},\textbf{m})=\int~p(\textbf{y}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})p(\textbf{y}_{u}\mid\boldsymbol{\xi},\boldsymbol{\psi})p\left(\boldsymbol{\xi},\boldsymbol{\psi}\right)d\boldsymbol{\boldsymbol{\xi}}d\boldsymbol{\psi}d\textbf{y}_{u}$ (Blei et al.,, 2017).
+The ELBO can be written as an expectation with respect to the variational distribution and is given by
+$\mathcal{L}(\boldsymbol{\lambda})=\mathbb{E}_{{q}_{\boldsymbol{\lambda}}}\left[\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})-\text{log}~q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})\right],$ (16)
+where $\mathbb{E}_{{q}_{\boldsymbol{\lambda}}}\left[\cdot\right]$ denotes the expectation with respect to $q_{\boldsymbol{\lambda}}$.
 Table 2 provides expressions for $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ for different SEMs.
-To maximise ELBO given in Equation (S6) with respect to variational parameters, $\boldsymbol{\lambda}$, stochastic gradient ascent (SGA) methods are used (Nott et al.,2012Titsias and Lázaro-Gredilla,20142015).
+To maximise ELBO given in Equation (S6) with respect to variational parameters, $\boldsymbol{\lambda}$, the stochastic gradient ascent (SGA) method is used (Nott et al.,2012Titsias and Lázaro-Gredilla,20142015).
 The SGA method updates the initial value for $\boldsymbol{\lambda}$ (say $\boldsymbol{\lambda}^{(0)}$) according to the iterative scheme,
-[…]
-[…]
+$\boldsymbol{\lambda}^{(t+1)}=\boldsymbol{\lambda}^{(t)}+\mathbcal{a}^{(t)}\circ\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$, (17)
+where $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})}$ is an unbiased estimate of the true gradient $\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$ of the ELBO in Equation (S6), $\mathbcal{a}^{(t)}$ ($t=0,1,\ldots$) is a vector of adaptive step sizes, obtained using the ADADELTA method (Zeiler,, 2012), described in Section S3 of the online supplement to facilitate rapid convergence of the SGA algorithm.
+The symbol $\circ$ represents the element-wise product of two vectors.
+The updating of Equation (17) is done until a stopping criterion is satisfied.
+Further details on the SGA method are provided in Section S2 of the online supplement.
+In standard VB inference, the missing data vector, $\textbf{y}_{u}$, is treated as a set of parameters, and the joint posterior distribution of $\boldsymbol{\xi}$, $\boldsymbol{\psi}$, and $\textbf{y}_{u}$ is approximated using the joint variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$.
+However, Wijayawardhana et al., (2025) demonstrated that this joint approximation is often insufficiently flexible to capture the posterior distributions of model parameters and $\textbf{y}_{u}$ simultaneously, due to the complex dependencies among the spatially correlated elements of $\textbf{y}_{u}$.
+To address this limitation, Section 4.2 introduces an alternative hybrid algorithm that effectively combines VB optimisation with MCMC sampling.
 4.2 Hybrid variational Bayes
-To efficiently handle missing values $\textbf{y}_{u}$ in SEMs described in Section 2, we extend the hybrid VB method of Wijayawardhana et al., (2025) that uses the Gaussian variational approximation with a factor covariance structure to approximate the posterior distribution of the parameters, while employing an MCMC sampler to generate samples of $\textbf{y}_{u}$ from its conditional distribution, given $\textbf{y}_{o}$, m, $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$.
-We first factorise the variational distribution that approximates $\boldsymbol{\xi},\boldsymbol{\psi}$, and $\textbf{y}_{u}$, $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ as follows:
+This section extends the hybrid VB (HVB) algorithm of Wijayawardhana et al., (2025) to efficiently handle missing values $\textbf{y}_{u}$ in non-Gaussian SEMs proposed in Section 2.
+The approach employs a Gaussian variational approximation with a factor covariance structure (Ong et al.,, 2018) to approximate the posterior distribution of the parameters, while employing an MCMC sampler to generate samples of $\textbf{y}_{u}$ from its conditional distribution, given $\textbf{y}_{o}$, m, $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$.
+We first factorise the joint variational distribution that approximates $\boldsymbol{\xi},\boldsymbol{\psi}$, and $\textbf{y}_{u}$, $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ as follows:
 $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p({\textbf{y}}_{u}\mid{\textbf{y}}_{o},\textbf{m},\boldsymbol{\xi,\boldsymbol{\psi}})q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi}),$ (18)
-where $p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ represents the conditional distribution of $\textbf{y}_{u}$, given the observed data ($\textbf{y}_{o}$, m), and the model parameters ($\boldsymbol{\xi}$ and $\boldsymbol{\psi}$).
-The term $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ is the Gaussian variational approximation of Ong et al., (2018), used to approximate the posterior distribution of $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$.
-We rewrite the ELBO in Equation (S6) by substituting the factorised variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ from Equation (15).
-Using Bayes’ rule, it is straightforward to show
-$\mathcal{L}(\boldsymbol{\lambda})=E_{q}\left[\text{log}~{}p({\textbf{y}}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})+\text{log}~{}p(\boldsymbol{\xi},\boldsymbol{\psi})-\text{log}~{}q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\right]=\mathcal{L}^{0}(\boldsymbol{\lambda}),$ (16)
-where $\mathcal{L}^{0}(\boldsymbol{\lambda})$ is the ELBO resulting from approximating only the posterior distribution of parameters $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$, $p(\boldsymbol{\xi},\boldsymbol{\psi}\mid\textbf{y}_{o},\textbf{m})$, directly using the variational distribution $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$; see Wijayawardhana et al., (2025) for the complete proof.
-We assume that the variational distribution, $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$, follows a multivariate normal distribution with a factor covariance structure (Ong et al.,, 2018), $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\sim N\left((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top};\boldsymbol{\mu}_{s_{m}},\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2}\right)$, where $\boldsymbol{\mu}_{s_{m}}$ is an $s_{m}\times 1$ vector of variational means, $\textbf{B}_{s_{m}}$ is an $s_{m}\times p$ matrix with the upper triangular elements set to zero, and $\textbf{D}_{s_{m}}$ is an $s_{m}\times s_{m}$ diagonal matrix with positive diagonal elements $\textbf{d}_{s_{m}}=(d_{1},\dots,d_{s_{m}})^{\top}$.
-For all SEMs with missing values, the value of $s_{m}$ is provided in Table 2.
-The vector of variational parameters is given by $\boldsymbol{\lambda}_{s_{m}}=\left(\boldsymbol{\mu}_{s_{m}}^{\top},\text{vech}(\textbf{B}_{s_{m}})^{\top},\textbf{d}_{s_{m}}^{\top}\right)^{\top}$, where the ’vech’ operator vectorises a matrix by stacking its columns from left to right while removing all the elements above the diagonal (the superdiagonal elements) of the matrix.
-In the SGA algorithm (Equation (14)), we compute an unbiased estimate of the ELBO gradient, $\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\lambda)$, denoted as $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\lambda)}$, at each iteration using the reparameterisation trick (Kingma and Welling,, 2013).
-To do this, we first generate samples from the variational distribution $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ by drawing ${\boldsymbol{\delta}^{0}}=(\boldsymbol{\eta}^{0},\boldsymbol{\epsilon}^{0})\sim N(\mathbf{0},\mathbf{I}_{s_{m}+p})$, where $\boldsymbol{\eta}^{0}$ is a $p$-dimensional vector and $\boldsymbol{\epsilon}^{0}$ is an $s_{m}$-dimensional vector.
-Next, we compute $(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}=t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})=\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}$.
+where $p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ represents the conditional distribution of $\textbf{y}_{u}$, given the observed data ($\textbf{y}_{o}$, m), and the model parameters ($\boldsymbol{\xi}$, $\boldsymbol{\psi}$).
+The term $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ is the variational approximation used to approximate the posterior distribution of $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$.
+The ELBO in Equation (S6) can be rewritten by substituting the factorised variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ from Equation (18).
+By applying Bayes’ rule and simplifying, we obtain
+$\mathcal{L}(\boldsymbol{\lambda})=\mathbb{E}_{{q}_{\boldsymbol{\lambda}}}\left[\text{log}~p({\textbf{y}}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})+\text{log}~p(\boldsymbol{\xi},\boldsymbol{\psi})-\text{log}~q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\right]=\mathcal{L}^{0}(\boldsymbol{\lambda}),$ (19)
+where $\mathcal{L}^{0}(\boldsymbol{\lambda})$ is the ELBO resulting from approximating only the posterior distribution of parameters ($\boldsymbol{\xi}$, $\boldsymbol{\psi}$), that is, $p(\boldsymbol{\xi},\boldsymbol{\psi}\mid\textbf{y}_{o},\textbf{m})$, directly using the variational distribution $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$; see Wijayawardhana et al., (2025) for the complete proof.
+We assume that the variational distribution, $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$, follows a multivariate Gaussian distribution with a factor covariance structure (Ong et al.,, 2018), $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\sim N\left((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top};\boldsymbol{\mu},\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2}\right)$}, where $\boldsymbol{\mu}$ is an $s\times 1$ vector of variational means, B is an $s\times p$ matrix with the upper triangular elements set to zero ($p<<s$), and D is an $s\times s$ diagonal matrix with positive diagonal elements $\textbf{d}=(d_{1},\dots,d_{s})^{\top}$.
+For all SEMs with missing values, the value of $s$ is provided in Table 2.
+The vector of variational parameters is given by $\boldsymbol{\lambda}=\left(\boldsymbol{\mu}^{\top},\text{vech}(\textbf{B})^{\top},\textbf{d}^{\top}\right)^{\top}$, where the ’vech’ operator vectorises a matrix by stacking its columns from left to right while removing all the elements above the diagonal (the super-diagonal elements) of the matrix.
+We now describe how the SGA method is used in the proposed HVB algorithm.
+At each iteration, an unbiased estimate of the ELBO gradient, $\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$, denoted $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})}$, needs to be computed for the SGA update given in Equation (17).
+This is achieved using the reparameterisation trick (Kingma and Welling,, 2013).
+First, samples are generated from the variational distribution $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ by drawing ${\boldsymbol{\delta}^{0}}=(\boldsymbol{\eta}^{0},\boldsymbol{\epsilon}^{0})\sim N(\mathbf{0},\mathbf{I}_{s+p})$, where $\boldsymbol{\eta}^{0}$ is a $p$-dimensional vector and $\boldsymbol{\epsilon}^{0}$ is an $s$-dimensional vector.
+Next, we compute $(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}=t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})=\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0}$.
 The density of $\boldsymbol{\delta}^{0}$ is denoted by $f_{\boldsymbol{\delta}^{0}}(\boldsymbol{\delta}^{0})$.
-Let $\boldsymbol{\delta}=({\boldsymbol{\delta}^{0}}^{\top},\textbf{y}_{u}^{\top})^{\top}$, where the product density is expressed as $f_{\boldsymbol{\delta}}(\boldsymbol{\delta})=f_{\boldsymbol{\delta}^{0}}(\boldsymbol{\delta}^{0})p({\textbf{y}}_{u}\mid{\textbf{y}}_{o},\textbf{m},t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}}))$.
-Finally, the transformation from $\boldsymbol{\delta}$ to the parameter and augmented missing value space is given by $((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top},{\textbf{y}}_{u}^{\top})^{\top}=t(\boldsymbol{\delta},\boldsymbol{\lambda}_{s_{m}})=(t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top},\textbf{y}_{u}^{\top})^{\top}=((\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0})^{\top},\textbf{y}_{u}^{\top})^{\top}$.
+Let $\boldsymbol{\delta}=({\boldsymbol{\delta}^{0}}^{\top},\textbf{y}_{u}^{\top})^{\top}$, where the product density is expressed as $f_{\boldsymbol{\delta}}(\boldsymbol{\delta})=f_{\boldsymbol{\delta}^{0}}(\boldsymbol{\delta}^{0})p({\textbf{y}}_{u}\mid{\textbf{y}}_{o},\textbf{m},t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}))$}.
+Finally, the transformation from $\boldsymbol{\delta}$ to the parameter and augmented missing value space is given by $((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top},{\textbf{y}}_{u}^{\top})^{\top}=t(\boldsymbol{\delta},\boldsymbol{\lambda})=(t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})^{\top},\textbf{y}_{u}^{\top})^{\top}=((\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0})^{\top},\textbf{y}_{u}^{\top})^{\top}$.
 The reparameterisation gradient of the ELBO in Equation (19) is derived by differentiating under the integral sign as shown below:
-$\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})=\mathbb{E}_{f_{\boldsymbol{\delta}}}\left[\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top}}{d\boldsymbol{\lambda}_{s_{m}}}(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})-\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi}))\right],$ (17)
-where $\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})}{d\boldsymbol{\lambda_{s_{m}}}}$ is the derivative of the transformation $t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})=\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}$ with respect to the variational parameters $\boldsymbol{\lambda}_{s_{m}}=(\boldsymbol{\mu}_{s_{m}}^{\top},\text{vech}(\textbf{B}_{s_{m}})^{\top},\textbf{d}_{s_{m}}^{\top})^{\top}$; see Section S4.1 of the online supplement for the proof.
-The expressions for $\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top}}{d\boldsymbol{\lambda}_{s_{m}}}$ and $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ can be found in Section S4.1 of the online supplement.
-In addition, the expressions for $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ are provided in Section S4.3 for different SEMs.
+$\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})=\mathbb{E}_{f_{\boldsymbol{\delta}}}\left[\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})^{\top}}{d\boldsymbol{\lambda}}(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})-\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi}))\right],$ (20)
+where $\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})}{d\boldsymbol{\lambda}}$ is the derivative of the transformation $t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})=\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0}$ with respect to the variational parameters $\boldsymbol{\lambda}=(\boldsymbol{\mu}^{\top},\text{vech}(\textbf{B})^{\top},\textbf{d}^{\top})^{\top}$; see Section S5.1 of the online supplement for the proof.
+The expressions for $\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})^{\top}}{d\boldsymbol{\lambda}}$ and $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ can be found in Section S5.2 of the online supplement.
+In addition, the expressions for $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ are provided in Section S5.3 of the online supplement for different SEMs.
 Algorithm 1 outlines the HVB algorithm.
-For all four SEMs with missing data, the conditional distribution of missing data; $p(\mathbf{y}_{u}\mid\mathbf{y}_{o},\mathbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ is not available in closed form.
-To sample from $p(\mathbf{y}_{u}\mid\mathbf{y}_{o},\mathbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ in step 5 of Algorithm 1, we use the Markov chain Monte Carlo (MCMC) steps outlined in Algorithm 2.
+For all four SEMs with missing data, the conditional distribution of missing data, $p(\mathbf{y}_{u}\mid\mathbf{y}_{o},\mathbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ is not available in closed form.
+To sample from $p(\mathbf{y}_{u}\mid\mathbf{y}_{o},\mathbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ in step 5 of Algorithm 1, we use the MCMC steps outlined in Algorithm 2.
 Algorithm 1 Hybrid variational Bayes (HVB) algorithm.
-1:  Initialise $\boldsymbol{\lambda}_{s_{m}}^{(0)}=(\boldsymbol{\mu}_{s_{m}}^{\top(0)},\textrm{vech}{(\textbf{B}_{s_{m}})}^{\top(0)},\textbf{d}_{s_{m}}^{\top(0)})^{\top}$ and set $t=0$
+1: Initialise $\boldsymbol{\lambda}^{(0)}=(\boldsymbol{\mu}^{\top(0)},\textrm{vech}{(\textbf{B})}^{\top(0)},\textbf{d}^{\top(0)})^{\top}$ and set $t=0$
 2:repeat
-3:     Generate $({\boldsymbol{\eta}^{0}}^{(t)},{\boldsymbol{\epsilon}^{0}}^{(t)})\sim N(\textbf{0},\textbf{I}_{s_{m}+p})$
-4:     Generate $(\boldsymbol{\xi}^{(t)^{\top}},\boldsymbol{\psi}^{(t)^{\top}})^{\top}\sim q_{\boldsymbol{\lambda}_{s_{m}}^{(t)}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ using its reparameterised representation.
+3:  Generate $({\boldsymbol{\eta}^{0}}^{(t)},{\boldsymbol{\epsilon}^{0}}^{(t)})\sim N(\textbf{0},\textbf{I}_{s+p})$
+4:  Generate $(\boldsymbol{\xi}^{(t)^{\top}},\boldsymbol{\psi}^{(t)^{\top}})^{\top}\sim q_{\boldsymbol{\lambda}^{(t)}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$ using its reparameterised representation.
 5:  Generate $\textbf{y}_{u}^{(t)}~\sim p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi}^{(t)},\boldsymbol{\psi}^{(t)})$
-6:     Construct unbiased estimates $\widehat{{\nabla_{\boldsymbol{\mu}_{s_{m}}}\mathcal{L}(\boldsymbol{\lambda})}},\widehat{{\nabla_{\textrm{vech}{(\textbf{B}_{s_{m}})}}\mathcal{L}(\boldsymbol{\lambda})}},\text{and}\widehat{{\nabla_{\textbf{d}_{s_{m}}}\mathcal{L}(\boldsymbol{\lambda})}}$ using Equations  (S44), (S45) and  (S46) in Section S4.1 of the online supplement.
-[…]
-[…]
-[…]
-[…]
-11:     Set $\boldsymbol{\lambda}_{s_{m}}^{(t+1)}=(\boldsymbol{\mu}_{s_{m}}^{\top(t+1)},\textrm{vech}(\textbf{B}_{s_{m}})^{\top(t+1)},\textbf{d}_{s_{m}}^{\top(t+1)})^{\top}$, and $t=t+1$
+6:  Construct unbiased estimates $\widehat{{\nabla_{\boldsymbol{\mu}}\mathcal{L}(\boldsymbol{\lambda})}},\widehat{{\nabla_{\textrm{vech}{(\textbf{B})}}\mathcal{L}(\boldsymbol{\lambda})}},\text{and}\widehat{{\nabla_{\textbf{d}}\mathcal{L}(\boldsymbol{\lambda})}}$ using Equations  (S46), (S47) and  (S48) in Section S5.2 of the online supplement.
+7:  Set adaptive learning rates $\mathbcal{a}^{(t)}_{\boldsymbol{\mu}}$ , $\mathbcal{a}^{(t)}_{\text{vech}(\textbf{B})}$ and $\mathbcal{a}_{\textbf{d}}^{(t)}$, using ADADELTA described in Section S3 of the online supplement.
+8:  Set $\boldsymbol{\mu}^{(t+1)}=\boldsymbol{\mu}^{(t)}+\mathbcal{a}^{(t)}_{\boldsymbol{\mu}}\circ\widehat{\nabla_{\boldsymbol{\mu}}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$.
+9:  Set $\textrm{vech}(\textbf{B})^{(t+1)}=\textrm{vech}(\textbf{B})^{(t)}+\mathbcal{a}_{\text{vech}(\textbf{B})}^{(t)}\circ\widehat{\nabla_{\textrm{vech}(\textbf{B})}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$.
+10:  Set $\textbf{d}^{(t+1)}=\textbf{d}^{(t)}+\mathbcal{a}_{\textbf{d}}^{(t)}\circ\widehat{\nabla_{\textbf{d}}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$.
+11:  Set $\boldsymbol{\lambda}^{(t+1)}=(\boldsymbol{\mu}^{\top(t+1)},\textrm{vech}(\textbf{B})^{\top(t+1)},\textbf{d}^{\top(t+1)})^{\top}$, and $t=t+1$
 12:until some stopping rule is satisfied
@@
 1: Initialise missing values $\textbf{y}_{u,0}~\sim p(\textbf{y}_{u}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{o})$
-2:  for i=1, …, $N_{1}$ do
+2:for $i=1,\dots,N_{1}$ do
 3:  Sample $\widetilde{\textbf{y}}_{u}$ from the proposal distribution $p(\widetilde{\textbf{y}}_{u}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{o})$.
 4:  Sample $u$ from uniform distribution, $u~\sim\mathcal{U}(0,1)$
-5:     Calculate $a=\text{min}\left(1,\frac{p(\textbf{m}\mid\widetilde{\textbf{y}},\boldsymbol{\psi}^{(t)})}{p(\textbf{m}\mid\textbf{y}_{{i-1}},\boldsymbol{\psi}^{(t)})}\right)$, where $\widetilde{\textbf{y}}=(\textbf{y}_{o}^{\top},{\widetilde{\textbf{y}}_{u}}^{\top})$ and $\textbf{y}_{i-1}=(\textbf{y}^{\top}_{o},{\textbf{y}^{\top}_{u,{i-1}}})$
+5:  Calculate $a=\text{min}\left(1,\frac{p(\textbf{m}\mid\widetilde{\textbf{y}},\boldsymbol{\psi}^{(t)})}{p(\textbf{m}\mid\textbf{y}_{{i-1}},\boldsymbol{\psi}^{(t)})}\right)$, where $\widetilde{\textbf{y}}=(\textbf{y}_{o}^{\top},{\widetilde{\textbf{y}}_{u}}^{\top})^{\top}$ and $\textbf{y}_{i-1}=(\textbf{y}^{\top}_{o},{\textbf{y}^{\top}_{u,{i-1}}})^{\top}$
 6:  if $a>u$ then
@@
 We refer to the modified HVB algorithm with block-wise updates as HVB-AllB (which uses MCMC steps in Algorithm S2), and the original version without block-wise updates as HVB-NoB (which uses MCMC steps in Algorithm 2) in the following sections.
-3.3 Bayesian model comparison
-To assess model fit in both the simulation study (Section 4) and the real-world analysis (Section 5), we employ extended versions of the standard Deviance Information Criterion (DIC, Spiegelhalter et al., (2002)) proposed by Celeux et al., (2006).
-This section first outlines DIC methods for models without missing data, then for those with missing data.
+5 Bayesian model comparison
+To assess the model fit of different SEMs in the subsequent sections of this study, we employ extended versions of the standard Deviance Information Criterion (DIC; Spiegelhalter et al., (2002)) as proposed by Celeux et al., (2006).
+This section first outlines DIC methods for models without missing data, followed by those for models with missing data.
+5.1 Model comparison with full data
 For full data SEMs, both $\text{DIC}_{1}$ and $\text{DIC}_{2}$ are computed to evaluate model fit, with lower values indicating better fit. $\text{DIC}_{1}$ is defined as:
-$\text{DIC}_{1}=-4E_{\boldsymbol{\phi}}[\text{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})]+2~{}\text{log}~{}p(\textbf{y}\mid\overline{\boldsymbol{\phi}}),$ (18)
-where $\overline{\boldsymbol{\phi}}$ is the posterior mean of ${\boldsymbol{\phi}}$, and $p(\textbf{y}\mid\boldsymbol{\phi})$ denotes the density of y given the model parameters $\boldsymbol{\phi}$.
-Both the expectation and the posterior mean $\overline{\boldsymbol{\phi}}$ are computed using samples drawn from the approximate posterior distribution estimated by the proposed VB method.
-For all SEMs, the logarithm of the density $p(\textbf{y}\mid\boldsymbol{\phi})$ is provided in Section S5 of the online supplement.
+$\text{DIC}_{1}=-4\mathbb{E}_{\boldsymbol{\phi}}[\text{log}~p(\textbf{y}\mid\boldsymbol{\phi})]+2~\text{log}~p(\textbf{y}\mid\overline{\boldsymbol{\phi}}),$ (21)
+[…]
 $\text{DIC}_{2}$ is defined as:
-$\text{DIC}_{2}=-4E_{\boldsymbol{\phi}}[\text{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})]+2~{}\text{log}~{}p(\textbf{y}\mid\widehat{\boldsymbol{\phi}}),$ (19)
+[…]
 where $\widehat{\boldsymbol{\phi}}$ is the value of $\boldsymbol{\phi}$} that maximises the function $p(\textbf{y}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$, estimated using samples drawn from the approximate posterior obtained by the proposed VB method.
+The expectation is also computed using the samples from the approximate posterior.
+5.2 Model comparison with missing responses
+$\text{DIC}_{5}$ is used to assess model fit for SEMs with missing data and is defined as:
+[…]
+where $p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})$ is the likelihood of y and m, conditional on model parameters $\boldsymbol{\phi}$ and $\boldsymbol{\psi}$.
+The values $\widehat{\boldsymbol{\phi}}$, $\widehat{\boldsymbol{\psi}}$, and $\widehat{\textbf{y}}_{u}$ are the posterior samples of $\boldsymbol{\phi}$, $\boldsymbol{\psi}$ and $\textbf{y}_{u}$ that maximise the function $p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})p(\boldsymbol{\phi},\boldsymbol{\psi})$, where the samples are drawn from the approximate posterior distribution estimated by the proposed HVB method in Section 4.2.
 The expectation is also computed using these variational posterior samples.
-$\text{DIC}_{5}$ is used to assess model fit for SEMs with missing data and is defined as:
-$\text{DIC}_{5}=-4E_{\boldsymbol{\phi},\boldsymbol{\psi},\textbf{y}_{u}}[\text{log}~{}p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})]+2~{}\text{log}~{}p(\textbf{y}_{o},\widehat{\textbf{y}}_{u},\textbf{m}\mid\widehat{\boldsymbol{\phi}},\widehat{\boldsymbol{\psi}}),$ (20)
-where $p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})$ is the joint density of y and m, conditional on model parameters $\boldsymbol{\phi}$ and $\boldsymbol{\psi}$.
-The values $\widehat{\boldsymbol{\phi}}$, $\widehat{\boldsymbol{\psi}}$, and $\widehat{\textbf{y}}_{u}$ are the posterior sample values of $\boldsymbol{\phi}$, $\boldsymbol{\psi}$ and $\textbf{y}_{u}$ that maximise the density $p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})p(\boldsymbol{\phi},\boldsymbol{\psi})$, where the samples are drawn from the approximate posterior estimated by the proposed HVB method.
-The expectation is also computed using these variational posterior samples.
-Note that $p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})=p(\textbf{y}\mid\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})$; see the selection model factorisation in Equation (11).
-4 Simulation study
-The simulation study is divided into two subsections.
-Section 4.1 compares the posterior distributions obtained using variational Bayes (VB) methods with those of Hamiltonian Monte Carlo (HMC, Neal, (1996)).
-The posterior densities estimated using the HMC method are considered as ground truth to assess the accuracy of the VB methods.
-In Section 4.1, we utilise a small simulated dataset ($n=625$), as implementing HMC is computationally intensive for SEMs, particularly when handling large datasets with missing data (Wijayawardhana et al.,, 2025).
-Section 4.2 uses two moderately large simulated datasets, each considered with and without missing values, and each exhibiting distinct characteristics, to evaluate the robustness of the proposed models and VB methods.
-We now describe the prior distributions used in this study for both the VB and HMC methods, applied in the simulation and real-world examples presented in Sections 4.1, 4.2, and 5.
+Note that $p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})=p(\textbf{y}\mid\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})$; see the selection model factorisation in Equation (13) in Section 3.2.
+[…]
+6 Simulation study: Assessing the accuracy of VB methods
+This section compares the accuracy of the proposed VB methods (with full data and with missing data) against the Hamiltonian Monte Carlo (HMC) (Neal,, 2011) method.
+The HMC method is implemented using the RStan interface (Stan Development Team,, 2023), specifically employing the No-U-Turn Sampler (NUTS) algorithm (Hoffman and Gelman,, 2014).
+The posterior densities estimated using HMC are considered as ground truth to assess the accuracy of the VB methods.
+A relatively small simulated dataset ($n=625$) is used for this comparison, as HMC becomes computationally expensive for SAR models with larger datasets, particularly in the presence of missing data (Wijayawardhana et al.,, 2025).
+The simulation study presented in Section 7 and the real data application in Section 8 use moderately large datasets (n=4,378n=4{,}378).
+The prior distributions used for both the VB and HMC algorithms throughout this paper are described next.
 To map the parameters $\sigma^{2}_{\textbf{e}}$, $\rho$, $\nu$, and $\gamma$ onto the real line, we apply the following transformations: $\omega^{\prime}=\text{log}(\sigma^{2}_{\textbf{e}})$, $\rho^{\prime}=\text{log}(1+\rho)-\text{log}(1-\rho)$, $\nu^{\prime}=\text{log}(\nu-3)$, and $\gamma^{\prime}=\text{log}(\gamma)-\text{log}(2-\gamma)$.
-Table S3 gives the prior distributions for the transformed parameters.
+The prior distributions are then specified on the transformed parameter space.
+Table 3 summarises the priors and hyperparameters.
+We use diffuse (non-informative) priors for all model parameters so that posterior inference is driven primarily by the data.
+For example, the fixed effects $\boldsymbol{\beta}$ are assigned weakly informative normal priors, $\boldsymbol{\beta}\sim N(\textbf{0},100\textbf{I})$, where I is an identity matrix of appropriate dimension (see Table 3).
+For the relatively large datasets analysed in this paper, the choice of these non-informative hyperparameters has little impact on the parameter estimates or on the convergence behaviour of the VB and HMC algorithms.
+Table 3: Prior distributions of model parameters
 Parameter $\boldsymbol{\beta}$ $\omega^{\prime}$ $\rho^{\prime}$ $\nu^{\prime}$ $\gamma^{\prime}$ $\boldsymbol{\psi}$
@@
 Hyperparameters $\sigma^{2}_{\boldsymbol{\beta}}=10^{2}$ $\sigma^{2}_{\omega^{\prime}}=10^{2}$ $\sigma^{2}_{\rho^{\prime}}=10^{2}$ $\sigma^{2}_{\nu^{\prime}}=10^{2}$ $\sigma^{2}_{\gamma^{\prime}}=10^{2}$ $\sigma^{2}_{\boldsymbol{\psi}}=10^{2}$
-Table 3: Prior distributions of model parameters
-4.1 Assessing the accuracy of variational Bayes approximations
-This section compares the accuracy of the proposed VB approximation methods with the computationally intensive, exact HMC methods.
-The comparison focuses solely on the YJ-SEM-Gau model, both with and without missing data.
-To facilitate this comparison, a dataset with 625 observations is simulated from the YJ-SEM-Gau model with five covariates generated from standard normal distributions.
-The weight matrix, W is constructed based on a $25\times 25$ regular grid, where neighbouring units are defined using the rook neighbourhood criteria, as described in Lloyd, (2010).
-The model parameters are specified as follows: the fixed effects ($\boldsymbol{\beta}$) are randomly drawn from discrete uniform values between $-3$ and $3$ (excluding 00).
-The error variance parameter is set to $\sigma^{2}_{\mathbf{e}}=1$ and the spatial autocorrelation parameter is set to $\rho=0.8$.
-The Yeo Johnson (YJ) parameter is specified as $\gamma=1.25$, inducing moderate left skewness into the simulated response variable.
-See Figure S1 in Section S7.1 of the online supplement for the kernel density plot of the response variables.
-The VB algorithm, outlined in Algorithm S1 in Section S2 of the online supplement, and the HMC algorithm are used to fit the YJ-SEM-Gau model to the simulated dataset.
-The initial values for the VB algorithm are set as follows: The variational mean vector ($\boldsymbol{\mu}$) is initialised using the estimates of $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ obtained from fitting the SEM-Gau model via maximum likelihood estimation (MLE) method.
-The initial value for $\gamma$ is set to 1.
+The comparison in this section is based on a simulated dataset generated from the YJ-SEM-Gau.
+The dataset contains 625 observations with five covariates, all generated from standard normal distributions.
+The spatial weight matrix, W, is based on a $25\times 25$ regular lattice, where neighbouring units are defined using the rook neighbourhood (Lloyd,, 2010).
+The model parameters are specified as follows: the fixed effects ($\boldsymbol{\beta}$) are randomly drawn from a discrete uniform distribution over 3-3 to $3$ (excluding 0).
+The error variance is set to $\sigma^{2}_{\mathbf{e}}=1$ and the spatial autocorrelation parameter is $\rho=0.8$.
+The Yeo–Johnson parameter is set to $\gamma=1.25$, introducing moderate left skewness into the simulated response variable.
+See Figure S1 in Section S7 of the online supplement for the kernel density plot of the response variable.
+Section 6.1 compares the proposed VB method with HMC using this dataset without missing values, while Section 6.2 compares the proposed HVB method with HMC in the presence of missing responses by introducing missing values into this dataset.
+6.1 SEMs with full data
+We now assess the accuracy of the VB method for SEMs with full data.
+The VB algorithm, outlined in Algorithm S1 in Section S2 of the online supplement, and the HMC algorithm are used to fit the YJ-SEM-Gau to the simulated dataset.
+The initial values for the VB algorithm are set as follows: the variational mean vector ($\boldsymbol{\mu}$) is initialised using the estimates of $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ obtained from fitting the SEM-Gau via maximum likelihood (ML) estimation method.
+The variational mean corresponding to the parameter $\gamma$ is initialised to 1.
 For the variational covariance matrix parameters, the elements of B and the diagonal elements of D are all initialised to 0.01.
-We use $p=4$ factors.
-The results do not improve when we increase the number of factors.
-We used similar starting values for the HMC algorithms as well.
+A value of $p=4$ factors is used.
+The results do not improve when the number of factors is increased.
+The HMC algorithm is initialised in the same way, with its initial parameter values set to the initial variational means described above.
 Both the VB and HMC algorithms are run for 10,000 iterations, at which point convergence is achieved.
@@
 In contrast, the VB algorithm first estimates the variational parameters, $\boldsymbol{\lambda}$, by running the variational optimisation.
-Then, 10,000 posterior draws of the model parameters are generated from the variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$.
-Figure 1 compares the posterior densities of selected model parameters estimated using the VB method and the HMC algorithm implemented via RSTAN (Stan Development Team,, 2023) to fit the simulated dataset.
-For the fixed effect parameter (${\beta}_{1}$), the spatial autocorrelation parameter ($\rho$), the variance parameter ($\sigma^{2}_{\textbf{e}}$), and the YJ parameter $\gamma$, the posterior densities estimated by both methods are nearly identical and include the true values.
-Figure 1: Comparison of posterior densities of selected parameters in the YJ-SEM-Gau model, obtained using the HMC and VB methods on the simulated dataset without missing values.
+Then, 10,000 posterior draws of the model parameters, $\boldsymbol{\xi}=\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},{\sigma}^{2}_{e},\rho,\gamma)^{\top}$}, are generated from the variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$; see Section S2 of the online supplement.
+Figure 1 compares the posterior densities of selected model parameters estimated using the VB and HMC algorithms.
+For all model parameters, the posterior densities obtained from both methods are almost indistinguishable and successfully recover the true values.
+Figure 1: Comparison of posterior densities of selected parameters in the YJ-SEM-Gau, obtained using the HMC and VB methods on the simulated dataset without missing values.
 The vertical line represents the true parameter value.
-Next, we introduce missing values into the response variable y under the MNAR mechanism, using the logistic regression model described in Section 2.3.
-To do this, we generate an additional covariate $\textbf{x}^{*}$ from a log-normal distribution with a log-scale mean of 0 and a log-scale standard deviation of 1.
-This covariate is included in the missing data model, which is specified as a logistic regression model with $\textbf{x}^{*}$ and the response variable y as covariates (see Equation (10)).
-The regression coefficients are set as follows: $\psi_{0}=-1.0$ (intercept), $\psi_{\textbf{x}^{*}}=0.5$ (the coefficient of $\textbf{x}^{*}$), and $\psi_{\textbf{y}}=-0.1$ (the coefficient of y).
+6.2 SEMs with missing responses
+Now, missing values are introduced into the response variable y under the MNAR mechanism using the logistic model described in Equation (11), in order to compare the performance of the HVB algorithm presented in Section 4.2 with HMC.
+To serve as a covariate in the logistic model, an additional variable $\textbf{x}^{\ast}$ is generated from a log-normal distribution with log-scale mean 0 and log-scale standard deviation 1.
+This simulated covariate, and the response variable y from the YJ-SEM-Gau, are used as the predictors in the missing value model defined in Equation (11).
+The coefficients are set as follows: $\psi_{0}=-1.0$ (intercept), $\psi_{\textbf{x}^{\ast}}=0.5$ (coefficient of $\textbf{x}^{\ast}$), and $\psi_{\textbf{y}}=-0.1$ (coefficient of y).
 This configuration yields approximately $50\%$ missing values in y, corresponding to $n_{u}=320$.
-The HVB-NoB algorithm (without block-wise update) described in Algorithm 1 and the HMC algorithm (implemented via RSTAN) are used to estimate the YJ-SEM-Gau model.
-The tuning parameter, $N_{1}$ of the HVB-NoB algorithm, is set to 10.
-The initial values for the algorithm are set as follows: the variational mean vector ($\boldsymbol{\mu}$) is initialised using the estimates of $\boldsymbol{\beta}$, $\sigma^{2}_{\mathbf{e}}$, and $\rho$ obtained by fitting the SEM-Gau model via MLE method.
-The initial value of the transformation parameter $\gamma$ is set to 1.
-All fixed-effect parameters of the missing value model ($\boldsymbol{\psi}$) are initialised to 0.1.
-For the variational covariance matrix, all elements of $\mathbf{B}$ and the diagonal elements of $\mathbf{D}$ are initialised to 0.01.
-We use $p=4$ factors, as increasing the number of factors did not lead to improved results.
-Similar initial values are used for the HMC algorithm.
-Both the HVB-NoB and HMC algorithms are run for 10,000 iterations.
-Convergence diagnostics plots are provided in Section S9.1 of the online supplement.
-Now, we briefly explain how the posterior sample of model parameters, $\boldsymbol{\xi}$, $\boldsymbol{\psi}$, and missing values $\boldsymbol{y}_{u}$ are generated in the HVB-NoB method.
-Once the algorithm converges, the set of variational parameters, $\boldsymbol{\lambda}_{s_{m}}=\left(\boldsymbol{\mu}_{s_{m}}^{\top},\text{vech}(\textbf{B}_{s_{m}})^{\top},\textbf{d}_{s_{m}}^{\top}\right)^{\top}$, is obtained.
-Given these variational parameters, $10,000$ draws from the variational distribution, $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\sim N\left((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top};\boldsymbol{\mu}_{s_{m}},\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2}\right)$, are generated, which form the posterior sample of $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$.
-Based on this posterior sample of model parameters, the posterior sample of $\boldsymbol{y}_{u}^{(i)}$ is generated from the conditional distribution $p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi}^{(i)},\boldsymbol{\psi}^{(i)})$, where $\boldsymbol{\xi}^{(i)}$ and $\boldsymbol{\psi}^{(i)}$ are the $i^{th}$ posterior sample for $i=1,\dots,10,000$.
-Figure 2 compares the posterior densities of selected model parameters estimated using HVB-NoB and HMC for fitting the YJ-SEM-Gau model on the simulated dataset with missing data.
-For most model parameters, the posterior densities obtained from both methods are nearly identical.
-However, the posterior density of $\gamma$ estimated using the HVB-NoB method is slightly overestimated compared to that of the HMC method.
-Figure 2: Comparison of posterior densities of selected parameters in the YJ-SEM-Gau model, obtained using the HMC and HVB-NoB methods on the simulated dataset with missing values.
+[…]
+Now, we briefly explain how posterior samples of the model parameters, $\boldsymbol{\xi}$, $\boldsymbol{\psi}$, and the missing values $\boldsymbol{y}_{u}$, are generated in the HVB-NoB method.
+Once the algorithm converges, the set of variational parameters $\boldsymbol{\lambda}=\left(\boldsymbol{\mu}^{\top},\mathrm{vech}(\mathbf{B})^{\top},\mathbf{d}^{\top}\right)^{\top}$ is obtained.
+Given these variational parameters, 10,000 draws from the variational distribution
+$q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})=N\!\left((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top};\,\boldsymbol{\mu},\,\mathbf{B}\mathbf{B}^{\top}+\mathbf{D}^{2}\right)$
+are generated, forming the posterior samples of $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$. Conditional on these parameter samples, posterior samples of $\mathbf{y}_{u}^{(i)}$ are drawn from
+$p(\mathbf{y}_{u}\mid\mathbf{y}_{o},\mathbf{m},\boldsymbol{\xi}^{(i)},\boldsymbol{\psi}^{(i)}),$
+using the MCMC steps described in Algorithm 2, where $\boldsymbol{\xi}^{(i)}$ and $\boldsymbol{\psi}^{(i)}$ denote the $i^{\text{th}}$ posterior draws of the model parameters, for $i=1,\dots,10{,}000$.
+Figure 2 compares the posterior densities of selected model parameters from the YJ-SEM-Gau and the missing data model, estimated using HVB-NoB and HMC on the simulated dataset with missing values.
+For most parameters, the posterior densities produced by the two methods are nearly identical.
+The posterior density of $\gamma$ estimated using HVB-NoB is slightly different compared to HMC, which is expected since HVB is an approximate inference method, whereas HMC yields asymptotically exact posterior samples.
+Importantly, this minor loss in accuracy is offset by the substantial computational efficiency gained with HVB-NoB, as discussed in Section 6.3.
+Figure 2: Comparison of posterior densities for selected parameters of the YJ-SEM-Gau and missing data models, estimated using the HMC and HVB-NoB methods on the simulated dataset with missing values.
 The vertical line represents the true parameter value.
-Notably, the computation time per iteration of the VB and HVB-NoB algorithms is significantly lower than that of the HMC algorithm.
-The computation time per iteration for the VB algorithm is approximately 0.0361, while for the HMC algorithm, it is 0.4680 when estimating the YJ-SEM-Gau model without missing data.
-Similarly, the computation time per iteration for the HVB-NoB algorithm is approximately 0.0542 seconds, while for the HMC algorithm, it is 41.1658 seconds for estimating the YJ-SEM-Gau model with missing data.
-The high computational cost of the HMC algorithm in estimating SEMs (such as SEM-Gau) with missing data is also highlighted in the findings of Wijayawardhana et al., (2025).
-4.2 Evaluating the performance of proposed SEMs
-This section focuses on evaluating the robustness of the proposed SEMs and the VB methods using moderately large simulated datasets.
-We use a spatial weight matrix from the spData R package (Bivand et al.,, 2023), based on $25,357$ houses sold between 1993 and 1998 in Lucas County, Ohio, USA.
-Further details on this dataset are provided in Section 5.
-We simulate two datasets from the YJ-SEM-t model using the spatial weight matrix $\textbf{W}_{1998}$, which corresponds to 4,378 houses sold in 1998 in Lucas County, Ohio, USA.
-In the first dataset, we set the parameters $\gamma=0.5$ (right skew) and $\nu=4$ (heavy tails).
-In the second dataset, we set the parameters $\gamma=1$ (no skew) and $\nu=30$ (light tails).
-The other parameters are set as follows: 6 fixed effects ($\boldsymbol{\beta}$) are randomly drawn from the discrete uniform values between -3 and 3 (excluding 0), $\sigma^{2}_{\textbf{e}}=0.5$, $\rho=0.8$, and all covariates are generated from a standard normal distribution, $N(0,1)$.
-Figure 3 shows kernel density plots of the response variables for both datasets.
-Figure 3: Kernel density plots of the simulated data: The left panel shows the kernel density of the response variable from YJ-SEM-t with $\nu=4$ and $\gamma=0.5$, displaying right skewness and heavy tails (referred to as dataset 1).
-The right panel shows the kernel density of the response variable from YJ-SEM-t with $\nu=30$ and $\gamma=1$, which appears nearly symmetric with light tails (referred to as dataset 2).
-Section S6 of the online supplement fits the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models to simulated datasets 1 and 2 without missing data, and evaluates how effectively these models, along with their VB estimation methods, capture the characteristics of these datasets.
-Section 4.2.1 introduces missing values into both datasets and evaluates the performance of the models and estimation method (the HVB method) in the presence of missing values.
-4.2.1 SEMs with missing data
-This section evaluates the accuracy of the HVB method for estimating the SEMs with missing values in the response variable using simulated data.
-For each simulated dataset, missing values in the response variable y are introduced according to the MNAR mechanism using a logistic regression model defined in Section 2.3.
-To construct this model, an additional covariate $\mathbf{x}^{*}$ is generated from a log-normal distribution with with a log-scale mean of 0 and a log-scale standard deviation of 1, and the model is regressed on $\mathbf{x}^{*}$ and the response variable y.
-We set $\psi_{0}=1$ (intercept), $\psi_{\textbf{x}^{*}}=-1$ (the coefficient for $\textbf{x}^{\ast}$), and $\psi_{\textbf{y}}=-0.1$ (the coefficient for y).
-This configuration results in approximately $40\%$ missing values, leading to $n_{u}=1,752$.
-We fit SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models for the simulated datasets with missing values.
-As $n$ and $n_{u}$ are large, the HVB-AllB algorithm (Algorithm S2) described in Section S4.2 of the online supplement is used.
-The tuning parameters for the HVB-AllB algorithm is set as follows: the number of MCMC iterations ($N_{1}$) is set to 10, and the block size ($k^{*}$) is defined as $10\%$ of $n_{u}$, resulting in 10 or 11 blocks.
-The initial values for the variational parameters corresponding to $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, $\rho$, and $\gamma$ are set as described in Section 4.1, while $\nu$ is initialised with a value of 4.
-The initial values for the missing data ($\mathbf{y}_{u}$) are simulated from the conditional distribution $p(\mathbf{y}_{u}\mid\boldsymbol{\xi}^{(0)},\mathbf{y}_{o})$, where $\boldsymbol{\xi}^{(0)}$ represents the vector of initial parameter values $\boldsymbol{\xi}$.
-We ran the HVB-AllB algorithm for 10,000 iterations for the SEM-Gau and YJ-SEM-Gau models across both simulated datasets.
-For the SEM-t and YJ-SEM-t models, more iterations were required to achieve convergence due to the additional complexity introduced by estimating the latent variables $\boldsymbol{\tau}$ and the degrees of freedom parameter $\nu$.
-Specifically, for simulated dataset 1, SEM-t was run for 20,000 iterations and YJ-SEM-t for 10,000 iterations.
-For simulated dataset 2, SEM-t required 30,000 iterations, while YJ-SEM-t was run for 20,000 iterations.
-Table 4 presents the posterior means and 95% credible intervals for some of the model parameters of the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models obtained using the HVB-AllB method applied to the simulated dataset 1 with missing values, along with the computational cost for one VB iteration.
-The table also includes the $\text{DIC}_{5}$ values for each model, calculated using the formula in Equation (20).
-Corresponding posterior density plots of the parameters are given in Section S7.2.2 of the online supplement.
-We start by comparing the estimated values of the fixed effects (specifically $\beta_{0}$ and ${\beta}_{1}$ for the SEM and $\psi_{0}$ and ${\psi}_{1}$ for the missing data model), the variance parameter $\sigma_{\textbf{e}}^{2}$, and the spatial autocorrelation parameter $\rho$, as these parameters are common to all four models.
-The posterior means for $\beta_{0}$, $\beta_{1}$, $\psi_{0}$, $\psi_{1}$, and $\rho$ are nearly identical for both YJ-SEM-Gau and YJ-SEM-t, closely matching the true values.
-In contrast, the estimates from SEM-Gau and SEM-t show significant deviations from the true values.
-The estimated posterior means of $\sigma_{\textbf{e}}^{2}$ from SEM-Gau, SEM-t, and YJ-SEM-Gau differ considerably from the true value, while the posterior mean from YJ-SEM-t is much closer to the true value.
-The models incorporating the YJ transformation (YJ-SEM-Gau and YJ-SEM-t) successfully recover the parameter $\gamma$.
-The estimates of $\nu$ obtained from the SEM-t and YJ-SEM-t models are slightly inaccurate.
-Furthermore, the posterior means of the estimated missing values from YJ-SEM-t and YJ-SEM-Gau align more closely with the true missing values compared to those from SEM-Gau and SEM-t; see Figure S7 in Section S7.2.2 of the online supplement.
-For the simulated dataset 1 with missing values in the response variable, the YJ-SEM-t model provides the best fit, as indicated by the lowest $\text{DIC}_{5}$ value, among all four models.
-This result is expected, as the dataset was generated from the YJ-SEM-t model, which incorporates both heavy tails and skewness.
-The second-best performer is YJ-SEM-Gau, which has the next lowest DIC value.
-Furthermore, similar to the case without missing values discussed in Section S6 of the online supplement, the relatively small difference in the $\text{DIC}_{5}$ values between YJ-SEM-t and YJ-SEM-Gau suggests that YJ-SEM-Gau achieves a model fit nearly comparable to that of YJ-SEM-t for this dataset.
-The left panel of Figure 4 shows the kernel densities of the true missing values ($\mathbf{y}_{u}$) and the kernel density of posterior means of the missing values at locations $s_{1},s_{2},\ldots,s_{n}$, estimated from different SEMs for the simulated dataset 1.
-The densities from both YJ-SEM-Gau and YJ-SEM-t are nearly identical and closely align with the density of the true missing values.
-The right panel of Figure 4 presents the posterior density of the maximum missing value ($\textbf{y}_{u_{\textrm{max}}}$) predicted by each SEM.
-The true value of the maximum missing value lies closer to the posterior densities produced by YJ-SEM-Gau and YJ-SEM-t than to those from SEM-Gau and SEM-t.
-These outcomes are as expected, since YJ-SEM-Gau and YJ-SEM-t yield similar parameter estimates and DIC values (see Table 4).
-Figure S8 in Section S7.2.2 of the online supplement provides similar density plots for the simulated dataset 2.
-Figure 4: Left panel: The kernel density of the true missing values ($\textbf{y}_{u}$) and the kernel density of posterior means of the missing values ($\textbf{y}_{u}$) obtained using the HVB-AllB method for different SEMs for the simulated dataset 1 with missing values.
-Right panel: The posterior density of the maximum missing value ($\textbf{y}_{u_{\textrm{max}}}$) obtained by different SEMs for the simulated dataset 1 with missing values.
-The true value of the maximum missing value is indicated by the vertical line.
+6.3 Summary and computation time comparison
+Overall, the simulation results demonstrate that the proposed VB methods, both with full data and with missing data, provide accurate approximations to the posterior distributions obtained via HMC.
+For all key model parameters, the VB-based posterior densities closely match those from HMC, with only a slight overestimation observed for the transformation parameter $\gamma$ in the missing data setting.
+Notably, the computation times per iteration of the VB and HVB-NoB algorithms are significantly lower than those of the HMC algorithm.
+The computation time per iteration for the VB algorithm is approximately 0.0361 seconds, while for the HMC algorithm, it is 0.4680 seconds when estimating the YJ-SEM-Gau without missing data.
+The computation time per iteration for the HVB-NoB algorithm is approximately 0.0542 seconds, while for the HMC algorithm, it is 41.1658 seconds for estimating the YJ-SEM-Gau with missing data.
+These results demonstrate that the proposed VB and HVB-NoB methods provide accurate parameter estimates while offering substantial computational savings relative to HMC.
+7 Simulation study: Assessing the accuracy and robustness of proposed SEMs
+This section evaluates the robustness and accuracy of the proposed SEMs and the VB methods using moderately large simulated datasets.
+The objective of this simulation study is to assess how well the proposed SEMs capture key characteristics of non-Gaussian spatial data, such as skewness and heavy-tailed distributions, both with full data and in the presence of missing values.
+We generate datasets using a spatial weight matrix from the spData R package (Bivand et al.,, 2023), based on $25,357$ houses sold between 1993 and 1998 in Lucas County, Ohio, USA (see Section 8 for further details).
+We simulate two datasets from the YJ-SEM-t using the spatial weight matrix $\textbf{W}_{1998}$, corresponding to 4,378 houses sold in 1998:
+  1. 1.
+    Dataset 1: $\gamma=0.5$ (right skew) and $\nu=4$ (heavy tails).
+  2. 2.
+    Dataset 2: $\gamma=1$ (no skew) and $\nu=30$ (light tails).
+The remaining parameters for both datasets are set as follows: six fixed effect parameters ($\boldsymbol{\beta}$) are randomly drawn from a discrete uniform distribution between -3 and 3 (excluding 0), $\sigma^{2}_{\textbf{e}}=0.5$, $\rho=0.8$, and all covariates are generated from a standard normal distribution, $N(0,1)$.
+Figure 3 displays kernel density plots of the response variables for both datasets.
+Figure 3: Kernel density plots of the simulated data: The left panel displays the kernel density of the response variable from the YJ-SEM-t with $\nu=4$ and $\gamma=0.5$, displaying right skewness and heavy tails (dataset 1).
+The right panel displays the kernel density of the response variable from the YJ-SEM-t with $\nu=30$ and $\gamma=1$, which appears nearly symmetric with light tails (dataset 2).
+Section S8 of the online supplement fits the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t to simulated datasets 1 and 2 with complete data and evaluates how effectively these models, along with their VB estimation methods, capture the underlying characteristics of the datasets.
+In Section 7.1, missing values are introduced, and the performance of the SEMs along with the HVB estimation method (described in Section 4.2) is assessed under missing data.
+7.1 SEMs with missing responses
+This section evaluates the robustness and accuracy of the proposed SEMs and the HVB method in the presence of missing responses, using simulated datasets 1 and 2.
+First, missing values are introduced in the response variable y of each simulated dataset under the MNAR mechanism using the logistic regression model in Equation (11).
+To construct this model, an additional covariate $\mathbf{x}^{*}$ is generated from a log-normal distribution with a log-scale mean of 0 and a log-scale standard deviation of 1, and the model is regressed on $\mathbf{x}^{*}$ and the response variable y.
+For the dataset 1, the logistic regression parameters are set to $\psi_{0}=1$ (intercept), $\psi_{\textbf{x}^{*}}=-1$ (the coefficient for $\textbf{x}^{\ast}$), and $\psi_{\textbf{y}}=-0.1$ (the coefficient for y).
+This configuration results in approximately $40\%$ missing responses, leading to $n_{u}=1,729$.
+For the dataset 2, the intercept is set to $\psi_{0}=0.5$ (with the other coefficients unchanged), leading to approximately 40% missing responses ($n_{u}=1,774$).
+We fit SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t to these simulated datasets with missing values.
+As $n$ and $n_{u}$ are large, the HVB-AllB algorithm, which employs the MCMC steps detailed in Algorithm S2 in Section S5.4 of the online supplement, is utilised.
+The tuning parameters for the HVB-AllB algorithm are set as follows: the number of MCMC iterations ($N_{1}$) is set to 10, and the block size ($k^{*}$) is defined as $10\%$ of $n_{u}$, resulting in 11 blocks.
+The initial values for the variational mean vector $\boldsymbol{\mu}$ corresponding to $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, $\rho$, $\gamma$, as well as the initial values for the missing responses $\mathbf{y}_{u}$, are specified using the procedure described in Section 6.2.
+For models with Student’s $t$ errors, the degrees of freedom parameter $\nu$ and the set of latent variables $\boldsymbol{\tau}=(\tau_{1},\dots,\tau_{n})^{\top}$ must also be initialised.
+The variational mean corresponding to $\nu$ is initialised at 4.
+The variational means corresponding to the latent variables are initialised by sampling from the prior distribution $p(\tau_{i}\mid\nu^{(0)})\sim\text{IG}(\frac{\nu^{(0)}}{2},\frac{\nu^{(0)}}{2})$ for $i=1,\dots,n$, where $\nu^{(0)}=4$ denotes the initial value of $\nu$.
+For the variational covariance matrix, all elements of B and the diagonal elements of D are initialised to 0.01.
+We run the HVB-AllB algorithms for 10,000 iterations for the SEM-Gau and YJ-SEM-Gau across both simulated datasets.
+For the SEM-t and YJ-SEM-t, convergence generally requires more iterations due to the additional complexity introduced by the latent variables $\boldsymbol{\tau}$ and the parameter $\nu$.
+However, the number of required iterations varies across datasets and models.
+For simulated dataset 1, SEM-t is run for 20,000 iterations, whereas YJ-SEM-t achieves convergence within 10,000 iterations.
+For simulated dataset 2, SEM-t requires 30,000 iterations, while YJ-SEM-t is run for 20,000 iterations.
+Convergence diagnostic plots are provided in Section S11.2 of the online supplement.
+7.1.1 Results for the simulated dataset 1
+[…]
+[…]
+Table 4: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 1 with missing data, along with the computation time (CT) in seconds for one VB iteration.
+The table also presents the $\text{DIC}_{5}$ values for each model.
+Parameters labelled as ’NA’ indicate that they are not applicable to the corresponding model.
 SEM-Gau SEM-t YJ-SEM-Gau YJ-SEM-t
@@
 $\psi_{0}=1$ 0.8251 (0.7048, 0.9441) 0.9171 (0.8126, 1.0203) 0.9337 (0.8193, 1.0442) 0.9515 (0.8525, 1.0522)
-${\psi}_{\textbf{y}}=-0.1$0.0793 (-0.0907, -0.0679) -0.0997 (-0.1178, -0.0812) -0.0794 (-0.0924, -0.0663) -0.0790 (-0.0891, -0.0686)
+${\psi}_{\textbf{y}}=-0.1$ -0.0793 (-0.0907, -0.0679) -0.0997 (-0.1178, -0.0812) -0.0794 (-0.0924, -0.0663) -0.0790 (-0.0891, -0.0686)
 $\text{DIC}_{5}$ 36009.5 36302.7 18278.55 17855.59
 CT 0.5894 1.2250 0.6000 1.339
-Table 4: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 1 with missing data, along with the computation time (CT) in seconds for one HVB iteration.
+[…]
+The left four panels of Figure 4 compare the kernel density of the true missing values ($\textbf{y}_{u}$) with the kernel densities of the posterior means of the missing values estimated under different SEMs for the simulated dataset 1.
+The densities from both YJ-SEM-Gau and YJ-SEM-t are nearly identical and closely align with the density of the true missing values.
+The right panel of Figure 4 presents the posterior density of the maximum missing value, $\textrm{max}(\textbf{y}_{u})$, estimated under each SEM.
+The true value of the maximum missing value lies closer to the posterior densities produced by YJ-SEM-Gau and YJ-SEM-t than to those produced by SEM-Gau and SEM-t.
+This behaviour is expected, as YJ-SEM-Gau and YJ-SEM-t provide parameter estimates that are closer to the true values and yield more comparable $\text{DIC}_{5}$ values (see Table 4).
+Figure 4: Left panel: The kernel density of the true missing values ($\textbf{y}_{u}$) and the kernel densities of posterior means of the missing values obtained using the HVB-AllB method for different SEMs for the simulated dataset 1 with missing values.
+Right panel: The posterior density of the maximum missing value, $\textrm{max}(\textbf{y}_{u})$, obtained by different SEMs for the simulated dataset 1 with missing values.
+The true value of the maximum missing value is indicated by the vertical line.
+7.1.2 Results for the simulated dataset 2
+[…]
+Table 5: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 2 with missing data, along with the computation time (CT) in seconds for one VB iteration.
 The table also presents the $\text{DIC}_{5}$ values for each model.
-Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
-Table 5 reports the posterior means and $95\%$ credible intervals for key parameters of the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models, obtained using the HVB-AllB method on the simulated dataset 2 with missing values, along with the computational cost for one VB iteration.
-The table also presents the corresponding $\text{DIC}_{5}$ values for each model.
-Posterior density plots for these parameters are provided in Section S7.2.2 of the online supplement.
-The estimated posterior means for $\beta_{0}$, $\beta_{1}$, $\psi_{0}$, and $\sigma^{2}_{\textbf{e}}$ across all four models closely match their true values.
-Similarly, the estimated posterior mean of $\gamma$ for the YJ-SEM-Gau and YJ-SEM-t models is also close to the true value.
-The parameter $\nu$ is accurately estimated in both the SEM-t and YJ-SEM-t models.
-In addition, all four models produce posterior means for the estimated missing values that closely match the true values; see Figure S8 in Section S7.2.2 of the online supplement.
+Parameters labelled as ’NA’ indicate that they are not applicable to the corresponding model.
 SEM-Gau SEM-t YJ-SEM-Gau YJ-SEM-t
-${\beta}_{0}=-3$3.0131 (-3.0495, -2.9764) -3.0150 (-3.0497, -2.9792) -2.9511 (-2.9820, -2.9199) -2.9467 (-2.9767, -2.9161)
+${\beta}_{0}=-3$ -3.0131 (-3.0495, -2.9764) -3.0150 (-3.0497, -2.9792) -2.9511 (-2.9820, -2.9199) -2.9467 (-2.9767, -2.9161)
 ${\beta}_{1}=2$ 1.9942 (1.9582, 2.0309) 1.9884 (1.9605, 2.0165) 1.9803 (1.9621, 1.9979) 1.9834 (1.9568, 2.0096)
@@
 $\psi_{0}=0.5$ 0.4949 (0.3847, 0.6044) 0.5036 (0.4219, 0.5841) 0.5477 (0.4319, 0.6619) 0.5467 (0.4742, 0.6204)
-${\psi}_{\textbf{y}}=-0.1$0.0956 (-0.1210, -0.0706) -0.1090 (-0.1320, -0.0864) -0.0965 (-0.1106, -0.0827) -0.0972 (-0.1137, -0.0809)
+${\psi}_{\textbf{y}}=-0.1$ -0.0956 (-0.1210, -0.0706) -0.1090 (-0.1320, -0.0864) -0.0965 (-0.1106, -0.0827) -0.0972 (-0.1137, -0.0809)
 $\text{DIC}_{5}$ 14713.89 14221.78 14695.32 14164.86
 CT 0.5755 1.2166 0.6048 1.3486
-Table 5: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 2 with missing data, along with the computation time (CT) in seconds for one HVB iteration.
-The table also presents the $\text{DIC}_{5}$ values for each model.
-Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
-Based on the $\text{DIC}_{5}$ values, the YJ-SEM-t model offers the best fit for the simulated dataset 2 with missing values.
-In contrast, for the simulated dataset 2 without missing values, the SEM-Gau model provides the best fit; see Section S6 of the online supplement.
-For the simulated dataset 2 with missing data, the SEM-t model has the second-lowest $\text{DIC}_{5}$, indicating it is the second-best performing model.
-As shown in Tables 4 and 5, the HVB-AllB algorithms for SEM-t and YJ-SEM-t, which incorporate Student’s $t$ errors, require significantly more computation time than those for SEM-Gau and YJ-SEM-Gau.
-This increased computational cost arises from the greater complexity of VB (HVB-AllB) optimisation in models with $t$ errors.
-Specifically, these models introduce an additional length $n$ latent variable vector and include the degrees of freedom parameter $\nu$, both of which increase the dimensionality and complexity of the VB optimisation.
-Our simulations demonstrate that the standard Gaussian VB method for full data and the hybrid VB (HVB-NoB and HVB-AllB) methods for missing data offer accurate yet computationally efficient approximations for estimating both Gaussian and non-Gaussian SEMs, outperforming HMC in terms of computational cost.
-The YJ-SEM-Gau and YJ-SEM-t models are particularly effective for modeling data with pronounced skewness and heavy tails.
-In most scenarios, the DIC values show that YJ-SEM-Gau performs nearly as well as YJ-SEM-t.
-Although estimating YJ-SEM-t incurs significantly higher computational cost compared to SEM-Gau (the standard SEM with Gaussian errors), YJ-SEM-Gau has a similar computational cost to SEM-Gau while offering substantial improvements in modeling non-Gaussian data.
+Based on the $\text{DIC}_{5}$ values, the YJ-SEM-t provides the best fit for simulated dataset 2 with missing values, while the SEM-t has the second-lowest $\text{DIC}_{5}$, indicating it is the second-best performing model.
+[…]
+7.2 Summary and computation time comparison
+This simulation study evaluates how well the proposed SEMs capture key characteristics of non-Gaussian spatial data, both with fully observed data and in the presence of missing values, using two simulated datasets: dataset 1 exhibits skewness and heavy tails, while dataset 2 is closer to Gaussian.
+This section summarises the findings of the simulation study with missing data, while Section S8.3 of the online supplement summarises the corresponding results for the full data.
+As shown in Tables 4 and 5, the HVB-AllB algorithm applied to the SEM-t and YJ-SEM-t, both of which assume Student’s $t$-distributed errors, requires more computation time per iteration compared to SEM-Gau and YJ-SEM-Gau.
+This is due to the added complexity of VB optimisation: these models introduce an additional latent variable vector of length $n$ and include the degrees of freedom parameter $\nu$, increasing the dimensionality of the optimisation.
+In addition, SEM-t and YJ-SEM-t typically require more VB iterations to converge.
+Convergence diagnostics plots in Section S11.2 of the online supplement illustrate this behaviour.
+Consequently, HVB-AllB for SEM-t and YJ-SEM-t is generally more time-consuming than for SEM-Gau and YJ-SEM-Gau, both because each iteration is slower and more iterations are needed to achieve convergence.
+In terms of performance, YJ-SEM-Gau and YJ-SEM-t are particularly effective for modelling skewed and heavy-tailed data.
+For simulated dataset 1, YJ-SEM-t attains the lowest $\textrm{DIC}_{5}$ value, indicating the best fit, while YJ-SEM-Gau achieves the second-lowest value, which is close to that of YJ-SEM-t.
+For simulated dataset 2, which is closer to Gaussian, all four models recover the true parameter values accurately and yield nearly identical $\textrm{DIC}_{5}$ values, with YJ-SEM-t again achieving the lowest.
+Although estimating YJ-SEM-t incurs substantially higher computational cost compared to SEM-Gau (the conventional SEM with Gaussian errors), YJ-SEM-Gau maintains a computation time comparable to SEM-Gau while providing considerable improvements when modelling non-Gaussian data.
+Similar patterns are observed in the analysis of the fully observed datasets, as presented in Section S8 of the online supplement.
 8 Real data application
-We apply the proposed SEMs and VB methods to analyse a dataset of 25,357 single-family homes sold between 1993 and 1998, along with a spatial weight matrix of size $357\times 25,$ 357 × 25 , 357 from Lucas County, Ohio, USA.
+We apply the proposed SEMs and VB methods to a dataset of single-family homes sold in Lucas County, Ohio, USA.
 This dataset is available in the R package spData (Bivand et al.,, 2023).
@@
 House prices are typically positive and often exhibit skewness.
-To transform house prices into the entire real line and reduce skewness, researchers commonly apply logarithmic transformations before modeling (LeSage and Pace,, 2004; Wijayawardhana et al.,, 2024).
+To transform house prices into the entire real line and reduce skewness, researchers commonly apply logarithmic transformations before modelling (LeSage and Pace,, 2004; Wijayawardhana et al.,, 2024).
 Accordingly, we apply the natural logarithmic transformation to the house prices in the Lucas-1998-HP dataset.
-Despite this transformation, the resulting variable still exhibits left skewness (see Figure S9 in Section S8 of the online supplement).
-To evaluate the impact of missing data on Gaussian and non-Gaussian SEMs, we use the Lucas-1998-HP dataset, a complete real-world dataset.
-We first fit the SEMs to the full dataset using the VB algorithm.
-Then, we intentionally introduce missing values into the response variable and refit the SEMs using the proposed HVB-AllB method.
-This setup allows for a direct comparison of parameter estimates obtained with and without missing data.
-For the Lucas-1998-HP dataset, we fit all four models described in Section 2.
-The set of independent variables includes various powers of house age (age, $\text{age}^{2}$, and $\text{age}^{3}$), the natural logarithm of the lot size in square feet ($\ln(\text{lotsize})$), the number of rooms (rooms), the natural logarithm of the total living area in square feet ($\ln(\text{LTA})$), and the number of bedrooms.
-The natural logarithm of house price in hundreds of thousands is selected as the dependent variable.
-Results for the complete dataset (without missing values) are presented in Section 5.1, while Section 5.2 summarises the findings with missing data.
+However, even after the transformation, the resulting log-prices remain left-skewed; see Figure S9 in Section S10 of the online supplement.
+First, the proposed VB algorithm (Algorithm S1 in Section S2 of the online supplement) is used to estimate all four SEMs using the full Lucas-1998-HP dataset.
+Next, missing values are introduced into the response variable, and the SEMs are re-estimated using the proposed HVB-AllB algorithm (Algorithm 1, together with the MCMC steps in Algorithm S2 in Section S5.4 of the online supplement).
+This design enables a direct comparison between estimates obtained from the full dataset and those obtained in the presence of missing values, with the full-data estimates treated as the ground truth.
+The set of predictor variables includes various powers of house age (age, $\text{age}^{2}$, and $\text{age}^{3}$), the natural logarithm of the lot size in square feet ($\ln(\text{lotsize})$), the number of rooms (rooms), the natural logarithm of the total living area in square feet ($\ln(\text{LTA})$), and the number of bedrooms (beds).
+The natural logarithm of house price in hundreds of thousands of US dollars is selected as the response variable.
+Results for the complete dataset (without missing values) are presented in Section 8.1, while Section 8.2 provides the findings with missing data.
 8.1 Full data SEMs
-This section presents the results of analysing the Lucas-1998-HP dataset by fitting all four SEMs introduced in Section 2, using the VB algorithms outlined in Section S2 of the online supplement.
-The starting values for the VB algorithms are obtained as follows: The variational mean vector ($\boldsymbol{\mu}$) is set to the estimates for $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ obtained from fitting the SEM-Gau model via maximum likelihood estimation (MLE).
-The initial value for $\nu$ is set to 4, and the initial value for $\gamma$ is set to 1.
-For the variational covariance matrix parameters, the elements of B and the diagonal elements of D are all initialised to 0.01.
-We use $p=4$ factors.
-As in the simulation study with full data described in Section S6 of the online supplement, models with Student-$t$ errors required more iterations to achieve convergence.
-The VB algorithms were run for 20,000 iterations for the SEM-Gau and YJ-SEM-Gau models and 75,000 iterations for the SEM-t and YJ-SEM-t models, at which point convergence was attained.
-Convergence was assessed visually by examining the plots of variational means over iterations, which are provided in Section S9.2 of the online supplement.
-The estimated posterior means and $95\%$ credible intervals for selected model parameters across different models are summarised in Table 6, along with the corresponding DIC values.
-According to the $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values, the YJ-SEM-t model is the most appropriate SEM for the Lucas-1998-HP dataset, followed by the YJ-SEM-Gau model as the second most suitable.
+This section presents the results of fitting all four SEMs to the full Lucas-1998-HP dataset.
+The initial values for the VB algorithms are set following the same approach used in the simulation studies in Sections 6 and 7, and we use $p=4$ factors.
+Consistent with the simulation results with full data (see Section S8 of the online supplement), the models with Student’s $t$ errors require more iterations to achieve convergence.
+Accordingly, the VB algorithms are run for 20,000 iterations for SEM-Gau and YJ-SEM-Gau, and 75,000 iterations for SEM-t and YJ-SEM-t.
+Convergence is assessed visually using the trajectories of the variational means, with the diagnostic plots provided in Figure S20 in Section S11.3 of the online supplement.
+The estimated posterior means and $95\%$ credible intervals for selected model parameters across different SEMs, along with their corresponding DIC values, are summarised in Table 6.
+The posterior density plots for these parameters are provided in Figure S10 in Section S10 of the online supplement.
+According to the $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values, the YJ-SEM-t is the most appropriate SEM for the Lucas-1998-HP dataset, followed by the YJ-SEM-Gau as the second most suitable.
+Table 6: Posterior means and 95% credible intervals for selected model parameters of different SEMs for the Lucas-1998-HP dataset (without missing values).
+The table also includes DIC values for each model.
+Parameters labelled as ’NA’ indicate that they are not applicable to the corresponding model.
 SEM-Gau SEM-t YJ-SEM-Gau YJ-SEM-t
-$intercept$0.4336 (-0.4469, -0.4207) -0.3865 (-0.3979, -0.3752) -0.3020 (-0.3278, -0.2765) -0.2849 (-0.3087, -0.2612)
+interceptintercept -0.4336 (-0.4469, -0.4207) -0.3865 (-0.3979, -0.3752) -0.3020 (-0.3278, -0.2765) -0.2849 (-0.3087, -0.2612)
 $\beta_{\text{age}}$ 0.1650 (0.0557, 0.2742) -0.4197 (-0.5236, -0.3162) -0.1643 (-0.2926, -0.0321) -0.3201 (-0.3984, -0.2405)
@@
 $\text{DIC}_{2}$ 4469.655 5031.204 4332.07 3766.714
-Table 6: Posterior means and 95% credible intervals for selected parameters of different SEMs for the Lucas-1998-HP dataset (no missing values).
-The table also includes DIC values for each model.
-Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
 We now compare estimates of key parameters across different SEMs.
 A crucial aspect of SEM estimation is identifying the underlying spatial correlation, represented by the parameter $\rho$.
-For the SEM-Gau, YJ-SEM-Gau, and YJ-SEM-t models, the estimated values of $\rho$ are relatively close, with posterior mean values of 0.6291, 0.6019, and 0.5990, respectively.
+For the SEM-Gau, YJ-SEM-Gau, and YJ-SEM-t, the estimated values of $\rho$ are relatively close, with posterior mean values of 0.6291, 0.6019, and 0.5990, respectively.
 This suggests a moderate spatial correlation in the Lucas-1998-HP dataset.
-However, for the SEM-t model, the estimated $\rho$ is noticeably lower, with a posterior mean value of 0.5475, compared to the other three models.
+However, for the SEM-t, the estimated $\rho$ is noticeably lower, with a posterior mean value of 0.5475, compared to the other three models.
 Similar to linear regression models, the fixed effects $\boldsymbol{\beta}$ in SEMs represent the influence of covariates on the response variable, which in this case is the logarithm of the house price.
-An interesting disparity arises in the estimated posterior mean of the fixed effect corresponding to the covariate ’age’ ($\beta_{\text{age}}$).
-For the SEM-Gau model, $\beta_{\text{age}}$ is positive, while for the other three models, $\beta_{\text{age}}$ is negative.
-This suggests that for the SEM-Gau model, the age of the house is associated with an increase in house prices, while for the other models, the age of the house is linked to a decrease in house prices.
-The estimated posterior means of the fixed effect for the number of rooms ($\beta_{\text{rooms}}$) are 0.0068, 0.0188, 0.0311, and 0.0238 for the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models, respectively.
-These estimates suggest that the effect of the number of rooms on house prices varies across models.
-Notably, the YJ-SEM-Gau model shows the strongest positive influence, while the SEM-Gau model exhibits the weakest.
-Additionally, the SEM-t, YJ-SEM-Gau, and YJ-SEM-t models provide further insights into the characteristics of the dataset.
-The posterior mean of $\nu=3.0249$ for SEM-t and $\nu=7.9482$ for YJ-SEM-t suggest that the log-transformed house price distribution exhibits heavy tails.
-Furthermore, the posterior means of $\gamma$, $1.5264$ for YJ-SEM-Gau and $1.5529$ for YJ-SEM-t, suggest left skewness in the response variable, effectively capturing the asymmetry in the distribution of log-transformed house prices.
+An interesting disparity arises in the estimated posterior mean of the fixed effect for the covariate ’age’ ($\beta_{\text{age}}$).
+For the SEM-Gau, $\beta_{\text{age}}$ is positive, while for the other three models, $\beta_{\text{age}}$ is negative.
+This discrepancy reflects a key limitation of SEM-Gau: it is misspecified for this dataset because it cannot adequately capture important features of the response variable, such as skewness and heavy tails.
+As a result, SEM-Gau produces misleading estimates for the effect of house age, suggesting an increase in price rather than the decrease indicated by the more flexible models.
+Similarly, the estimated posterior means of the fixed effect for the number of rooms ($\beta_{\text{rooms}}$) are 0.0068, 0.0188, 0.0311, and 0.0238 for SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t, respectively, showing that the effect of the number of rooms on house prices varies across models.
+Notably, YJ-SEM-Gau shows the strongest positive influence, while SEM-Gau exhibits the weakest, further reflecting the consequences of misspecification.
+Additionally, the SEM-t, YJ-SEM-Gau, and YJ-SEM-t provide further insights into the characteristics of the dataset.
+The posterior means of $\nu=3.0249$ for SEM-t and $7.9482$ for YJ-SEM-t suggest that the log-transformed house price distribution exhibits heavy tails.
+Furthermore, the posterior means of $\gamma$, $1.5264$ for YJ-SEM-Gau and $1.5529$ for YJ-SEM-t, indicate left skewness in the response variable, effectively capturing its asymmetry.
 These distributional features are visually evident in the kernel density plot of the response variable (see Figure S9 in Section S10 of the online supplement).
-5.2 SEMs with missing data
-This section discusses the analysis of the Lucas-1998-HP dataset with missing data in the response variable.
-We use the age of the house and the response variable of SEMs (logarithm of the house price) as the covariates in the missing data model.
-The missing data model has three parameters, $\boldsymbol{\psi}=(\psi_{0},\psi_{\textbf{x}^{*}},\psi_{\textbf{y}})$.
-We set the values $\psi_{0}=0.1$, $\psi_{\textbf{x}^{*}}=0.1$, and $\psi_{\textbf{y}}=-0.1$.
-With these settings, approximately 50% of the responses are missing, resulting in $n_{u}=2,348$.
-Since both $n$ and $n_{u}$ are large, we utilise the HVB-AllB algorithm.
-We set the number of MCMC iterations $N_{1}$ to 10 and the block size to 10% of $n_{u}$, resulting in 11 blocks.
-The starting values for the HVB-AllB algorithms are obtained as follows: the variational mean vector ($\boldsymbol{\mu}$) is set to the estimates of $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ obtained from fitting the SEM-Gau model with only the observed data, $\textbf{y}_{o}$, using MLE.
-The initial value for $\nu$ is set to 4, and the initial value for $\gamma$ is set to 1.
-For the parameters of the missing data generating model ($\psi_{0}$, $\psi_{\textbf{x}^{*}}$, and $\psi_{\textbf{y}}$), all coefficients are set to an initial value of 0.01.
-For the variational covariance matrix parameters, the elements of B and the diagonal elements of D are all initialised to 0.01.
-The initial values for the missing data ($\mathbf{y}_{u}$) are simulated from the conditional distribution $p(\mathbf{y}_{u}\mid\boldsymbol{\xi}^{(0)},\mathbf{y}_{o})$, where $\boldsymbol{\xi}^{(0)}$ represents the vector of initial parameter values for $\boldsymbol{\xi}$.
-We use $p=4$ factors.
-Similar to the simulation study with missing data, models with Student’s $t$ errors required a relatively large number of iterations to converge.
-Therefore, we ran the HVB-AllB algorithm for 10,000 iterations for the SEM-Gau and YJ-SEM-Gau models, 40,000 iterations for the SEM-t model, and 20,000 iterations for the YJ-SEM-t model.
-Convergence plots can be found in Section S9.2 of the online supplement.
-The estimated posterior means and 95% credible intervals for selected model parameters across different models are summarised in Table 7, along with the corresponding $\text{DIC}_{5}$ values.
-According to the $\text{DIC}_{5}$ values, the YJ-SEM-t model is the most appropriate SEM for the Lucas-1998-HP dataset with missing data, followed by the YJ-SEM-Gau model as the second best model.
+8.2 SEMs with missing responses
+This section analyses the Lucas-1998-HP dataset with missing values in the response variable (logarithm of the house price).
+Starting from the complete dataset, missing responses are generated using the logistic regression model in Equation (11), with the age of the house and the response variable y (from the SEMs) serving as covariates.
+The missing data model parameters are set to $\psi_{0}=0.1$, $\psi_{\textbf{x}^{*}}=0.1$, and $\psi_{\textbf{y}}=-0.1$.
+This configuration results in approximately 50% of the responses being missing, yielding $n_{u}=2{,}348$.
+[…]
+[…]
+Table 7: Posterior means and 95% credible intervals for selected parameters of different SEMs for the Lucas-1998-HP dataset with missing data.
+The table also includes $\text{DIC}_{5}$ values for each model.
+Parameters labelled as ’NA’ indicate that they are not applicable to the corresponding model.
 SEM-Gau SEM-t YJ-SEM-Gau YJ-SEM-t
-$intercept$0.4189 (-0.4413, -0.3960) -0.3721 (-0.3886, -0.3562) -0.2856 (-0.3008, -0.2703) -0.2752 (-0.2956, -0.2544)
+interceptintercept -0.4189 (-0.4413, -0.3960) -0.3721 (-0.3886, -0.3562) -0.2856 (-0.3008, -0.2703) -0.2752 (-0.2956, -0.2544)
 $\beta_{\text{age}}$ 0.1935 (0.0335, 0.3510) -0.4640 (-0.5990, -0.3324) -0.1473 (-0.2598, -0.0345) -0.3768 (-0.4870, -0.2687)
@@
 $\text{DIC}_{5}$ 8074.382 7734.356 7309.394 6657.172
-Table 7: Posterior means and 95% credible intervals for selected parameters of different SEMs for the Lucas-1998-HP dataset with missing data.
-The table also includes $\text{DIC}_{5}$ values for each model.
-Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
 The estimated parameters from different SEMs are now briefly discussed.
-For the spatial autocorrelation parameter, $\rho$, the posterior mean estimates across all SEMs are slightly lower than those obtained without missing values, shown in Table 6.
+For the spatial autocorrelation parameter, $\rho$, the posterior mean estimates across all SEMs are slightly lower than those obtained using the full dataset without missing values, except under the SEM-t (see Table 6).
 However, these estimates still indicate a moderate spatial correlation in the Lucas-1998-HP dataset, consistent with the conclusions drawn from SEMs fitted without missing values in Section 8.1.
 The posterior mean estimates for the fixed effects ($\boldsymbol{\beta}$) in the SEMs lead to the same conclusions as those drawn from the Lucas-1998-HP dataset without missing values.
-For example, the posterior mean estimate of $\beta_{\text{age}}$ in the SEM-Gau model suggests a positive effect of age on house prices, whereas the estimates from the other three models indicate a negative effect.
-Moreover, the estimated intercepts in the missing data model, ${\psi}_{0}$, are close to the true value of $0.1$ for all models except SEM-t.
-The posterior mean estimates of $\nu$ for the SEM-t and YJ-SEM-t models indicate moderate to strong heavy-tailed behavior.
-Meanwhile, the estimated $\gamma$ values for the YJ-SEM-Gau and YJ-SEM-t models suggest left skewness in the data.
+For example, the posterior mean estimate of $\beta_{\text{age}}$ in the SEM-Gau suggests a positive effect of age on house prices, whereas the estimates from the other three models indicate a negative effect.
+Moreover, the estimated posterior means of intercept in the missing data model, ${\psi}_{0}$, are close to the true value of $0.1$ for all models except SEM-t.
+The posterior mean estimates of $\nu$ for the SEM-t and YJ-SEM-t indicate moderate to strong heavy-tailed behaviour.
+Similarly, the estimated posterior mean values of $\gamma$ for the YJ-SEM-Gau and YJ-SEM-t suggest left skewness in the response.
 These conclusions are consistent with the findings from the analysis without missing data discussed in Section 8.1.
@@
 Our article introduces three novel simultaneous autoregressive (SAR) models for non-Gaussian spatial data.
-Specifically, we extend the conventional spatial error model (SEM-Gau), a widely used SAR model, to accommodate skewed and heavy tailed features of real datasets.
+Specifically, we extend the conventional spatial error model (SEM-Gau), a widely used SAR model that assumes Gaussian errors, to accommodate skewed and heavy-tailed features of real datasets.
 The novel SEMs are SEM-t, YJ-SEM-Gau, and YJ-SEM-t.
-The SEM-t assumes that the error terms follow a Student’s t distribution, allowing heavier tails than a Gaussian distribution.
-The YJ-SEM-Gau retains Gaussian errors but applies the Yeo-Johnson (YJ) transformation to handle asymmetric response variables.
-The YJ-SEM-t combines these approaches by assuming Student’s t errors while incorporating the YJ transformation to account for skewness in the response variable.
-These extensions improve the robustness of SAR models in analysing complex spatial datasets.
+The SEM-t assumes that the error terms follow a Student’s $t$ distribution, allowing for heavier tails than the Gaussian case.
+The YJ-SEM-Gau retains Gaussian errors but applies the Yeo–Johnson (YJ) transformation to handle asymmetric response variables.
+The YJ-SEM-t combines these approaches by assuming Student’s $t$ errors while incorporating the YJ transformation.
+These extensions improve the robustness of SAR models for analysing non-Gaussian spatial datasets.
 Additionally, we develop efficient variational Bayes (VB) methods to estimate these models with and without missing data.
-For cases with missing data, our approach assumes a missing not at random (MNAR) mechanism.
+In cases with missing data, we assume that the responses are missing under the missing not at random (MNAR) mechanism.
 Standard VB methods are inadequate for accurately estimating these SEMs in the presence of missing data; therefore, we employ hybrid VB (HVB) algorithms to address this challenge.
-The posterior distributions obtained from the proposed VB and HVB approximation methods are compared with those of the exact Hamiltonian Monte Carlo (HMC) method.
-The empirical results show that: (1) The VB and HVB methods yield posterior density estimates for most model parameters that are similar to those obtained using the HMC method when estimating SEMs with and without missing values; (2) For models with Student’s $t$ errors (SEM-t and YJ-SEM-t), the posterior approximation of the degrees of freedom parameter ($\nu$) from the VB and HVB methods exhibits slight deviations compared to the true value; (3) For both VB and HVB algorithms, models with Student’s t errors require more computation time per iteration than models with Gaussian errors (SEM-Gau and YJ-SEM-Gau); (4) In general, models with Gaussian errors tend to converge in fewer iterations compared to those with Student’s t errors; (5) YJ-SEM-t is able to model complex response variable with skewness and heavy tails and is the best model for the Lucas-1998-HP dataset, both with and without missing data.
+The posterior distributions obtained from the proposed VB and HVB methods are compared with those of the exact Hamiltonian Monte Carlo (HMC) method, demonstrating both accuracy and substantial computational gains.
+Although the proposed model extensions are illustrated using the SEM, both the modelling framework and the associated VB/HVB inference methods are general and can be readily adapted to other SAR-type models.
+The empirical results show that: (1) The VB and HVB methods yield posterior density estimates for most model parameters that are similar to those obtained using the HMC method when estimating SEMs with and without missing values; (2) For models with Student’s $t$ errors (SEM-t and YJ-SEM-t), the posterior approximation of the degrees of freedom parameter ($\nu$) from the VB and HVB methods exhibits slight deviations compared to the true value; (3) For both VB and HVB algorithms, models with Student’s $t$ errors require more computation time per iteration than models with Gaussian errors (SEM-Gau and YJ-SEM-Gau); (4) In general, models with Gaussian errors tend to converge in fewer iterations compared to those with Student’s $t$ errors; (5) YJ-SEM-t is able to model complex response variables with skewness and heavy tails, and it provides the best fit to the Lucas-1998-HP dataset, both with and without missing data.
+Future work could explore more flexible transformation families, such as Tukey’s g-and-h transformation (Tukey,, 1977; Martinez and Iglewicz,, 1984), which jointly accommodate skewness and heavy tails.
+In addition, while the proposed non-Gaussian SEMs are developed for univariate spatial responses, extending the framework to multivariate and spatio-temporal settings would enable joint modelling of multiple non-Gaussian variables and dynamic processes.
+These extensions would substantially broaden the applicability of the models to a wider range of real-world spatial and spatio-temporal datasets.
 S1 Yeo and Johnson transformation and its derivatives
-This section provides the Yeo and Johnson (YJ) transformation (Yeo and Johnson,, 2000) and its derivatives, which are used to construct the YJ-SEM-Gau and YJ-SEM-t models discussed in Section 2.2 of the main paper.
-Let the random variable $y_{i}$ be asymmetric. The Yeo–Johnson (YJ) transformation, denoted $t_{\gamma}(y_{i})$, is used to make $y_{i}$ more symmetric, and is defined by:
+This section provides the Yeo and Johnson (YJ) transformation (Yeo and Johnson,, 2000) and its derivatives, which are used to construct the YJ-SEM-Gau and YJ-SEM-t discussed in Section 2.2 of the main paper.
+Let the random variable $y_{i}$ be asymmetric. The YJ transformation, denoted $t_{\gamma}(y_{i})$, is often applied to reduce asymmetry and make $y_{i}$ more symmetric. It is defined as:
 $y_{i}^{*}=t_{{\gamma}}(y_{i})=\begin{cases}\frac{(y_{i}+1)^{\gamma}-1}{\gamma}&\text{if }y_{i}\geq 0,\\ -\frac{(-y_{i}+1)^{2-\gamma}-1}{2-\gamma}&\text{if }y_{i}<0,\end{cases}$ (S1)
-where $20<\gamma<$ 2.
-The density function of SEMs with YJ transformations includes the derivative of the YJ transformation with respect to $y_{i}$, as shown in Equation (7) of the main paper. This derivative is equal to
+where $0<\gamma<2$.
+The density function of SEMs with YJ transformations includes the derivative of the YJ transformation with respect to $y_{i}$, as shown in Equation (8) of the main paper. This derivative is given by:
 $\frac{dt_{{\gamma}}(y_{i})}{dy_{i}}=\begin{cases}{(y_{i}+1)^{\gamma-1}}&\text{if }y_{i}\geq 0,\\ (-y_{i}+1)^{1-\gamma}&\text{if }y_{i}<0.\end{cases}$ (S2)
 S2 Variational Bayes approximation for SEMs with full data
-Table S1: The definitions of r, M, and the parameters $\boldsymbol{\phi}$ for the four spatial error models.
-Each SEM has a mean vector of $\mathbf{X}\boldsymbol{\beta}$ and a covariance matrix of the form $\boldsymbol{\Sigma}=\sigma^{2}_{\textbf{e}}\textbf{M}^{-1}$, with $\textbf{A}=\textbf{I}_{n}-\rho\textbf{W}$.
+In this section, we present the variational Bayes (VB) algorithm for estimating the SEMs without missing values introduced in Section 2 of the main paper.
+Table S1: The definitions of r, M, and the parameters $\boldsymbol{\phi}$ for SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t.
+For SEM-Gau and SEM-t, the density of the response vector conditional on $\boldsymbol{\xi}$ follows a multivariate Gaussian distribution with mean $\textbf{X}\boldsymbol{\beta}$ and covariance matrix $\sigma^{2}_{\textbf{e}}\textbf{M}^{-1}$, where $\textbf{A}=\textbf{I}_{n}-\rho\textbf{W}$.
 Model $\boldsymbol{\phi}$ r M
-SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $^{\top}\textbf{A}$ A
-SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$ A
-YJ-SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e},},\rho,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $^{\top}\textbf{A}$ A
-YJ-SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$ A
+SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $\textbf{A}^{\top}\textbf{A}$
+SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $\textbf{A}^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$
+YJ-SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e},},\rho,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $\textbf{A}^{\top}\textbf{A}$
+YJ-SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $\textbf{A}^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$
 Consider Bayesian inference for the parameter vector of SEMs, $\boldsymbol{\xi}$.
@@
 $p(\boldsymbol{\xi}\mid\textbf{y})\propto p(\textbf{y}\mid\boldsymbol{\xi})p(\boldsymbol{\xi}),$ (S3)
-where the term $p(\textbf{y}\mid\boldsymbol{\xi})$ represents the likelihood of y, and the log-likelihood of y for different SEMs are given in Section 2 of the main paper.
-We define $p(\textbf{y}\mid\boldsymbol{\xi})p(\boldsymbol{\xi})=h(\boldsymbol{\xi})$.
+where the term $p(\textbf{y}\mid\boldsymbol{\xi})$ represents the likelihood, and the log-likelihood for different SEMs are given in Section 2 of the main paper.
+We define $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\xi})p(\boldsymbol{\xi})$.
 The terms $h(\boldsymbol{\xi})$ for each SEM, along with the total length of the parameter vector to be estimated in each model (i.e., the length of the vector $\boldsymbol{\xi}$), are provided in Table S2.
-Table S2: The term $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\xi})p(\boldsymbol{\xi})$, and the total number of model parameters $s$ (including the length of the latent vector $\boldsymbol{\tau}$ if exists) for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models.
-The distribution $p(\boldsymbol{\tau}\mid\boldsymbol{\phi})=\prod_{i=1}^{n}p(\tau_{i}\mid\nu)$ represents the $n$ independent inverse gamma latent variables used in SEMs with Student-$t$ errors (see Equation (2) on the main paper).
+Table S2: The term $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\xi})p(\boldsymbol{\xi})$, and the total number of model parameters $s$ (including the length of the latent vector $\boldsymbol{\tau}$ if exists) for the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t.
+The density $p(\boldsymbol{\tau}\mid\boldsymbol{\phi})=\prod_{i=1}^{n}p(\tau_{i}\mid\nu)$ represents the $n$ independent inverse gamma latent variables used in SEMs with Student-$t$ errors (see Equation (2) of the main paper).
 Model $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\xi})p(\boldsymbol{\xi})$ $s$
@@
 YJ-SEM-t $p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$ $r+5+n$
-Now we explain how variational Bayes (VB) inference is used to approximate the posterior distribution in Equation (S3).
+Now we explain how VB inference is used to approximate the posterior distribution in Equation (S3).
 We use the variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$, indexed by the variational parameter $\boldsymbol{\lambda}$ to approximate the posterior $p(\boldsymbol{\xi}\mid\textbf{y})$.
@@
 […]
-[…]
+Minimising KL divergence between $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$ and $p(\boldsymbol{\xi}\mid\textbf{y})$ is equivalent to maximising evidence lower bound (ELBO) on the marginal log-likelihood, $\text{log}~p(\textbf{y})$, denoted by $\mathcal{L}(\boldsymbol{\lambda})$, with $p(\textbf{y})=\int~p(\textbf{y}\mid\boldsymbol{\xi})p\left(\boldsymbol{\xi}\right)d\boldsymbol{\xi}$ (Blei et al.,, 2017).
+The ELBO is
 […]
@@
 The ELBO in Equation (S5) can be written as an expectation with respect to $q_{\boldsymbol{\lambda}}$,
-$\mathcal{L}(\boldsymbol{\lambda})=E_{q}\left[\text{log}~{}h(\boldsymbol{\xi})-\text{log}~{}q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\right],$ (S6)
-where $E_{q}\left[\cdot\right]$ denotes the expectation with respect to $q_{\boldsymbol{\lambda}}$.
-We employ the Gaussian variational approximation, where we select $q_{\boldsymbol{\lambda}}$ to be a multivariate normal distribution.
+$\mathcal{L}(\boldsymbol{\lambda})=\mathbb{E}_{q}\left[\text{log}~h(\boldsymbol{\xi})-\text{log}~q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\right],$ (S6)
+where $\mathbb{E}_{q}\left[\cdot\right]$ denotes the expectation with respect to $q_{\boldsymbol{\lambda}}$.
+We employ the Gaussian variational approximation, where we select $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$ to be a multivariate Gaussian distribution.
 As a result, the variational parameters $\boldsymbol{\lambda}$ consist of both the mean vector and the distinct elements of the covariance matrix.
-We impose a factorised covariance structure (Ong et al.,, 2018) on the covariance matrix of $q_{\boldsymbol{\lambda}}$, which reduces the number of distinct elements in the covariance matrix.
-Under a factor covariance structure for the covariance matrix, the variational distribution is parameterised as $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\sim N(\boldsymbol{\xi};\boldsymbol{\mu},\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})$, where $\boldsymbol{\mu}$ is the $s\times 1$ mean vector, B is an $s\times p$ full rank matrix with $p<<s$, and D is an $s\times s$ diagonal matrix having diagonal elements $\textbf{d}=(d_{1},\ldots,d_{s})$.
+We impose a factor covariance structure (Ong et al.,, 2018) on the covariance matrix of $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$, which reduces the number of distinct elements in the covariance matrix.
+Under a factor covariance structure for the covariance matrix, the variational distribution is parameterised as $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\sim N(\boldsymbol{\xi};\boldsymbol{\mu},\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})$, where $\boldsymbol{\mu}$ is the $s\times 1$ mean vector, B is an $s\times p$ full rank matrix with $p<<s$, and D is an $s\times s$ diagonal matrix with positive diagonal elements $\textbf{d}=(d_{1},\ldots,d_{s})$.
 For all SEMs, the total number of parameters, $s$ is provided in Table S2.
@@
 The SGA method updates $\boldsymbol{\lambda}$ iteratively, starting from an initial value $\boldsymbol{\lambda}^{(0)}$, according to the following scheme:
-[…]
-[…]
-Minimising the variance of the gradient estimator, $\widehat{{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})}}$, is crucial for both stability and fast convergence of the SGA algorithm, which we address using the reparameterisation trick (Kingma and Welling,, 2013; Rezende et al.,, 2014).
-We now briefly explain how the reparameterisation trick, combined with a factor covariance structure for the Gaussian variational distribution, is used to obtain efficient gradient estimates in this article.
+$\boldsymbol{\lambda}^{(t+1)}=\boldsymbol{\lambda}^{(t)}+\mathbcal{a}^{(t)}\circ\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$, (S7)
+and continues until a stopping criterion is satisfied, where $\mathbcal{a}^{(t)}$ is a vector of element-wise learning rates at iteration $t$.
+The individual learning rates, denoted by $\mathcal{a}^{(t)}_{i}$, are typically chosen to satisfy the Robbins–Monro conditions: $\sum_{t}\mathcal{a}^{(t)}_{i}$$\infty$ and $\sum_{t}(\mathcal{a}^{(t)}_{i})^{2}<\infty$, which ensure convergence of the sequence $\boldsymbol{\lambda}^{(t)}$ to a local optimum as $t\rightarrow\infty$, under regularity conditions (Robbins and Monro,, 1951; Bottou,, 2010).
+The symbol $\circ$ denotes the element-wise product of two vectors.
+To speed up convergence, we use the ADADELTA algorithm for adaptive learning rates (Zeiler,, 2012), detailed in Section S3.
+Further, minimising the variance of the gradient estimator, $\widehat{{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})}}$, is crucial for both stability and fast convergence of the SGA algorithm, which we address using the reparameterisation trick (Kingma and Welling,, 2013; Rezende et al.,, 2014).
+We now briefly explain how the reparameterisation trick, combined with a factor covariance structure for the Gaussian variational distribution, is used to obtain efficient gradient estimates in this paper.
 More detailed information can be found in Ong et al., (2018).
@@
 Then, the expectation in Equation (S6) is expressed with respect to the distribution $f_{\boldsymbol{\zeta}}$ as
-$\begin{split}\mathcal{L}(\boldsymbol{\lambda})&=E_{q}\left[\text{log}~{}h(\boldsymbol{\xi})-\text{log}~{}q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\right]\\ &=E_{f_{\boldsymbol{\zeta}}}\left[\text{log}~{}h(u(\boldsymbol{\zeta,\boldsymbol{\lambda}}))-\text{log}~{}q_{\boldsymbol{\lambda}}(u(\boldsymbol{\zeta,\boldsymbol{\lambda}}))\right],\end{split}$ (S8)
+$\begin{split}\mathcal{L}(\boldsymbol{\lambda})&=\mathbb{E}_{q}\left[\text{log}~h(\boldsymbol{\xi})-\text{log}~q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\right]\\ &=\mathbb{E}_{f_{\boldsymbol{\zeta}}}\left[\text{log}~h(u(\boldsymbol{\zeta,\boldsymbol{\lambda}}))-\text{log}~q_{\boldsymbol{\lambda}}(u(\boldsymbol{\zeta,\boldsymbol{\lambda}}))\right],\end{split}$ (S8)
 and differentiating $\mathcal{L}(\boldsymbol{\lambda})$ under the integral sign, we obtain
-$\begin{split}\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\zeta}}}\left[\nabla_{\boldsymbol{\lambda}}~{}\text{log}~{}h(u(\boldsymbol{\zeta,\boldsymbol{\lambda}}))-\nabla_{\boldsymbol{\lambda}}~{}\text{log}~{}q_{\boldsymbol{\lambda}}(u(\boldsymbol{\zeta,\boldsymbol{\lambda}}))\right],\\ &=E_{f_{\boldsymbol{\zeta}}}\left[\frac{du(\boldsymbol{\zeta,\boldsymbol{\lambda}})^{\top}}{d\boldsymbol{\lambda}}\{\nabla_{\boldsymbol{\xi}}\text{log}~{}h(\boldsymbol{\xi})-\nabla_{\boldsymbol{\xi}}\text{log}~{}q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})\}\right],\\ \end{split}$ (S9)
+[…]
 where $\frac{du(\boldsymbol{\zeta,\boldsymbol{\lambda}})}{d\boldsymbol{\lambda}}$ is the derivative of the transformation $u(\boldsymbol{\zeta,\boldsymbol{\lambda}})=\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon}$ with respect to the variational parameters $\boldsymbol{\lambda}=(\boldsymbol{\mu}^{\top},\text{vech}(\textbf{B})^{\top},\textbf{d}^{\top})^{\top}$, where the ”vech” operator vectorises a matrix by stacking its columns from left to right while removing all the elements above the diagonal (the superdiagonal elements) of the matrix.
@@
 Algorithm S1 outlines the VB algorithm for estimating SEMs with full data.
-The unbiased estimate of the gradient, $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})}$, presented in step 4 of Algorithm S1 can be constructed using the reparameterisation trick described earlier, with a single sample drawn from $f_{\boldsymbol{\zeta}}(\boldsymbol{\zeta})$.
+The unbiased estimate of the gradient, $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})}$, in step 4 of Algorithm S1 can be constructed using the reparameterisation trick described earlier, with a single sample drawn from $f_{\boldsymbol{\zeta}}(\boldsymbol{\zeta})$.
 Algorithm S1 Variational Bayes algorithm
@@
 4:  Construct unbiased estimates $\widehat{{\nabla_{\boldsymbol{\mu}}\mathcal{L}(\boldsymbol{\lambda})}},\widehat{{\nabla_{\textrm{vech}(\textbf{B})}\mathcal{L}(\boldsymbol{\lambda})}},\text{and}~\widehat{{\nabla_{\textbf{d}}\mathcal{L}(\boldsymbol{\lambda})}}$ using Equations  (S11), (S12) and  (S13) in Section S2.1 at $\boldsymbol{\lambda}^{(t)}$.
-[…]
-[…]
-[…]
-[…]
+5:  Set adaptive learning rates for the variational means, $\mathbcal{a}^{(t)}_{\boldsymbol{\mu}}$ and the variational parameters $\textrm{vech}(\textbf{B})$ and d, $\mathbcal{a}_{\textrm{vech}{(\textbf{B})}}^{(t)}$ and $\mathbcal{a}_{\textbf{d}}^{(t)}$, respectively, using ADADELTA described in Section S3 of the online supplement.
+6:  Set $\boldsymbol{\mu}^{(t+1)}=\boldsymbol{\mu}^{(t)}+\mathbcal{a}^{(t)}_{\boldsymbol{\mu}}\circ\widehat{\nabla_{\boldsymbol{\mu}}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$.
+7:  Set $\textrm{vech}(\textbf{B})^{(t+1)}=\textrm{vech}(\textbf{B})^{(t)}+\mathbcal{a}_{\text{vech}(\textbf{B})}^{(t)}\circ\widehat{\nabla_{\textrm{vech}(\textbf{B})}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$.
+8:  Set $\textbf{d}^{(t+1)}=\textbf{d}^{(t)}+\mathbcal{a}_{\textbf{d}}^{(t)}\circ\widehat{\nabla_{\textbf{d}}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$.
 9:  Set $\boldsymbol{\lambda}^{(t+1)}=(\boldsymbol{\mu}^{\top(t+1)},\textrm{vech}(\textbf{B})^{\top(t+1)},\textbf{d}^{\top(t+1)})$, and $t=t+1$
 10:until some stopping rule is satisfied
-S2.1 Expressions for $\frac{du(\boldsymbol{\zeta},\boldsymbol{\lambda})}{d\boldsymbol{\lambda}}$
-In the reparameterisation gradient given in Equation (S9), $\frac{du(\boldsymbol{\zeta,\boldsymbol{\lambda}})}{d\boldsymbol{\lambda}}$ is the derivative of the transformation $u(\boldsymbol{\zeta,\boldsymbol{\lambda}})=\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon}$ with respect to the variational parameters $\boldsymbol{\lambda}=(\boldsymbol{\mu}^{\top},\text{vech}(\textbf{B})^{\top},\textbf{d}^{\top})^{\top}$.
+S2.1 Derivation of $\frac{du(\boldsymbol{\zeta},\boldsymbol{\lambda})}{d\boldsymbol{\lambda}}$, $\nabla_{\boldsymbol{\xi}}\log q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$ and $\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$
+In the reparameterisation gradient of the ELBO given in Equation (S9), $\frac{du(\boldsymbol{\zeta,\boldsymbol{\lambda}})}{d\boldsymbol{\lambda}}$ is the derivative of the transformation $u(\boldsymbol{\zeta,\boldsymbol{\lambda}})=\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon}$ with respect to the variational parameters $\boldsymbol{\lambda}=(\boldsymbol{\mu}^{\top},\text{vech}(\textbf{B})^{\top},\textbf{d}^{\top})^{\top}$.
 We write that $u(\boldsymbol{\zeta,\boldsymbol{\lambda}})=\boldsymbol{\mu}+(\boldsymbol{\eta}^{\top}\otimes\textbf{I}_{s})\text{vech}(\textbf{B})+\textbf{d}\circ\boldsymbol{\epsilon}$}, where $\otimes$ represents the Kronecker product, and $\textbf{I}_{s}$ is the identity matrix of size $s$.
 It can be shown that $\nabla_{\boldsymbol{\xi}}\text{log}~q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})=-(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\boldsymbol{\xi}-\boldsymbol{\mu})$},
-$\frac{du(\boldsymbol{\zeta},\boldsymbol{\lambda})}{d\boldsymbol{\mu}}=\textbf{I}_{s}\\ ~{}~{}~{}~{}~{}\text{and}~{}~{}~{}~{}~{}\frac{du(\boldsymbol{\zeta},\boldsymbol{\lambda})}{d\text{vech}(\textbf{B})}=\boldsymbol{\eta}^{\top}\otimes\textbf{I}_{s}.$ (S10)
-The derivatives of the lower bound with respect to variational parameters are:
-$\begin{split}\nabla_{\boldsymbol{\mu}}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\zeta}}}[\nabla_{\boldsymbol{\xi}}~{}\text{log}~{}h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})],\end{split}$ (S11)
-$\begin{split}\nabla_{\text{vech}(\textbf{B})}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\zeta}}}[\nabla_{\boldsymbol{\xi}}\text{log}~{}h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\eta}^{\top}\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\eta}^{\top}],\end{split}$ (S12)
+$\frac{\partial u(\boldsymbol{\zeta},\boldsymbol{\lambda})}{\partial\boldsymbol{\mu}}=\textbf{I}_{s}\\ ,~~~~~~~~~~\frac{\partial u(\boldsymbol{\zeta},\boldsymbol{\lambda})}{\partial\text{vech}(\textbf{B})}=\boldsymbol{\eta}^{\top}\otimes\textbf{I}_{s}\\ ,~~~\textrm{and}~~\frac{\partial u(\boldsymbol{\zeta},\boldsymbol{\lambda})}{\partial\textbf{d}}=\boldsymbol{\epsilon}\\ .$ (S10)
+The derivatives of the ELBO with respect to variational parameters are:
+$\begin{split}\nabla_{\boldsymbol{\mu}}\mathcal{L}(\boldsymbol{\lambda})&=\mathbb{E}_{f_{\boldsymbol{\zeta}}}[\nabla_{\boldsymbol{\xi}}~\text{log}~h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})],\end{split}$ (S11)
+$\begin{split}\nabla_{\text{vech}(\textbf{B})}\mathcal{L}(\boldsymbol{\lambda})&=\mathbb{E}_{f_{\boldsymbol{\zeta}}}[\nabla_{\boldsymbol{\xi}}\text{log}~h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\eta}^{\top}\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\eta}^{\top}],\end{split}$ (S12)
 and
-$\begin{split}\nabla_{\textbf{d}}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\zeta}}}[\text{diag}(\nabla_{\boldsymbol{\xi}}\text{log}~{}h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\epsilon}^{\top}\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\epsilon}^{\top})],\end{split}$ (S13)
+$\begin{split}\nabla_{\textbf{d}}\mathcal{L}(\boldsymbol{\lambda})&=\mathbb{E}_{f_{\boldsymbol{\zeta}}}[\text{diag}(\nabla_{\boldsymbol{\xi}}\text{log}~h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\epsilon}^{\top}\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})\boldsymbol{\epsilon}^{\top})],\end{split}$ (S13)
 where $\text{diag}(\cdot)$ is the vector of diagonal elements extracted from a square matrix.
 The analytical expressions for $\nabla_{\boldsymbol{\xi}}\text{log}~h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})=\nabla_{\boldsymbol{\xi}}\text{log}~h(\boldsymbol{\xi})$ in Equations (S11)-(S13) for different SEMs are provided in Section S2.2.
-S2.2 Expressions for $\nabla_{\boldsymbol{\xi}}\text{log}~{}h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})=\nabla_{\boldsymbol{\xi}}\text{log}~{}h(\boldsymbol{\xi})$
+S2.2 Derivation of $\nabla_{\boldsymbol{\xi}}\text{log}~h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon})=\nabla_{\boldsymbol{\xi}}\text{log}~h(\boldsymbol{\xi})$
 The term $h(\boldsymbol{\xi})$ represents the likelihood multiplied by the prior distribution (see Section S2).
 For different SEMs, the expressions for $h(\boldsymbol{\xi})$ are given in Table S2.
-Table S3 shows the prior distributions used for the model parameters in this study.
-In this section, we compute the derivatives of $\log h(\boldsymbol{\xi})$ with respect to the model parameter vector $\boldsymbol{\xi}$, denoted as $\nabla_{\boldsymbol{\xi}}\log h(\boldsymbol{\xi})$.
-These derivatives are necessary for computing the reparameterisation gradient, as outlined in Equation (S9), for each SEM.
+In this section, we present the derivatives (gradients) of $\log h(\boldsymbol{\xi})$ with respect to the model parameter vector $\boldsymbol{\xi}$, denoted as $\nabla_{\boldsymbol{\xi}}\log h(\boldsymbol{\xi})$.
+These gradients are required to construct unbiased estimates of the reparameterisation gradient.
+To map the parameters $\sigma^{2}_{\textbf{e}}$, $\rho$, $\nu$, and $\gamma$ onto the real line, we apply the following transformations: $\omega^{\prime}=\text{log}(\sigma^{2}_{\textbf{e}})$, $\rho^{\prime}=\text{log}(1+\rho)-\text{log}(1-\rho)$, $\nu^{\prime}=\text{log}(\nu-3)$, and $\gamma^{\prime}=\text{log}(\gamma)-\text{log}(2-\gamma)$.
+Prior distributions are then specified on the transformed parameter space.
+Table S3 shows the prior distributions used for the model parameters.
+Table S3: Prior distributions of model parameters
 Parameter $\boldsymbol{\beta}$ $\omega^{\prime}$ $\rho^{\prime}$ $\nu^{\prime}$ $\gamma^{\prime}$ $\boldsymbol{\psi}$
@@
 Hyperparameters $\sigma^{2}_{\boldsymbol{\beta}}=10^{2}$ $\sigma^{2}_{\omega^{\prime}}=10^{2}$ $\sigma^{2}_{\rho^{\prime}}=10^{2}$ $\sigma^{2}_{\nu^{\prime}}=10^{2}$ $\sigma^{2}_{\gamma^{\prime}}=10^{2}$ $\sigma^{2}_{\boldsymbol{\psi}}=10^{2}$
-Table S3: Prior distributions of model parameters
-S2.2.1 Derivatives for SEM-Gau
-For SEM-Gau, $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}$.
+S2.2.1 Gradients for SEM-Gau
+For SEM-Gau, we set $\boldsymbol{\xi}=\boldsymbol{\phi}$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\mathbf{e}},\rho)^{\top}$, so that $h(\boldsymbol{\xi})=p(\mathbf{y}\mid\boldsymbol{\phi})\,p(\boldsymbol{\phi})$.
 The logarithm of $h(\boldsymbol{\xi})$ is $\text{log}~p(\textbf{y}\mid\boldsymbol{\phi})+\text{log}~p(\boldsymbol{\phi})$.
-Note that, for $\sigma^{2}_{\textbf{e}}$, and $\rho$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3.
+Note that, for $\sigma^{2}_{\textbf{e}}$, and $\rho$, we utilise the transformations described in Section S2.2, and place the priors on the transformed parameters as in Table S3.
 This leads to
@@
 where $\text{tr}\{\cdot\}$ denotes the trace operator, which computes the sum of the diagonal elements of a matrix,
-$\begin{split}\left(\frac{\partial\textbf{A}^{\top}\textbf{A}}{\partial\rho}\right)&=-(\textbf{W}^{\top}+\textbf{W})+2\rho\textbf{W}^{\top}\textbf{W},\\ \left(\frac{\partial\rho}{\partial\rho^{\prime}}\right)&=\frac{2e^{\rho^{\prime}}}{(1+e^{\rho^{\prime}})^{2}}.\\ \end{split}$
+[…]
 S2.2.2 Gradients for SEM-t
-For SEM-t, $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu)^{\top}$.
-The logarithm of $h(\boldsymbol{\xi})$ is $p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\text{log}~{}p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\text{log}~{}p(\boldsymbol{\phi})$.
-Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$ and $\nu$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3.
+For the SEM-t, we set $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu)^{\top}$, so that $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$}.
+The logarithm of $h(\boldsymbol{\xi})$ is $\text{log}~p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\text{log}~p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\text{log}~p(\boldsymbol{\phi})$.
+Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$ and $\nu$, we utilise the transformations described in Section S2.2, and place the priors on the transformed parameters as in Table S3.
 This leads to
@@
 and the derivative of $\text{log}~h(\boldsymbol{\xi})$ with respect to $\nu^{\prime}$ is
-$\begin{split}\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\nu^{\prime}}&=n\left(\frac{\partial~{}\textrm{log}~{}\Gamma\left(\frac{e^{\nu^{\prime}+3}}{2}\right)}{\partial\nu^{\prime}}\right)-\frac{e^{\nu^{\prime}}}{2}\sum_{i=1}^{n}\left(\tau_{i}^{\prime}+e^{-\tau_{i}^{\prime}}\right)-\frac{ne^{\nu^{\prime}}}{2}\textrm{log}\left(\frac{2}{e^{\nu^{\prime}}+3}\right)+\frac{ne^{\nu^{\prime}}}{2}\\ &-\frac{\nu^{\prime}}{\sigma^{2}_{\nu^{\prime}}},\\ \end{split}$ (S22)
+[…]
 where $\frac{\partial~\textrm{log}~\Gamma\left(\frac{e^{\nu^{\prime}+3}}{2}\right)}{\partial\nu^{\prime}}$ is the derivative of the log-gamma function, which is computed numerically using R.
-The derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $i^{th}$ element of the vector $\boldsymbol{\tau}^{\prime}$, $\tau_{i}^{\prime}$ is
+The derivative of $\text{log}~h(\boldsymbol{\xi})$ with respect to $i^{\textrm{th}}$ element of the vector $\boldsymbol{\tau}^{\prime}$, $\tau_{i}^{\prime}$ is
 $\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\tau_{i}^{\prime}}=-\frac{1}{2}+\frac{1}{2}e^{-\omega^{\prime}}({\textbf{r}_{\textbf{A}}}_{i})^{2}e^{\tau_{i}^{\prime}}+\left(\frac{e^{\nu^{\prime}+3}}{2}\right)\left(-1+e^{-\tau_{i}^{\prime}}\right),$ (S23)
-where ${\textbf{r}_{\textbf{A}}}_{i}$ is the $i^{th}$ element of the vector Ar.
+where ${\textbf{r}_{\textbf{A}}}_{i}$ is the $i^{\textrm{th}}$ element of the vector Ar.
 S2.2.3 Gradients for YJ-SEM-Gau
-For YJ-SEM-Gau, $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\gamma)^{\top}$.
+For the YJ-SEM-Gau, we set $\boldsymbol{\xi}=\boldsymbol{\phi}$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\gamma)^{\top}$, so that $h(\boldsymbol{\xi})=p(\mathbf{y}\mid\boldsymbol{\phi})\,p(\boldsymbol{\phi})$}.
 The logarithm of $h(\boldsymbol{\xi})$ is $\text{log}~p(\textbf{y}\mid\boldsymbol{\phi})+\text{log}~p(\boldsymbol{\phi})$.
-Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$ and $\gamma$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3.
+Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$ and $\gamma$, we utilise the transformations described in Section S2.2, and place the priors on the transformed parameters as in Table S3.
 This leads to
@@
 $\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\gamma^{\prime}}=\left(-e^{-\omega^{\prime}}\textbf{r}^{\top}\textbf{M}\left(\frac{\partial t_{\gamma}(\textbf{y})}{\partial\gamma}\right)+\sum_{i=0}^{n}\frac{\partial\textrm{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right)}{\partial\gamma}\right)\left(\frac{\partial\gamma}{\partial\gamma^{\prime}}\right)-\frac{\gamma^{\prime}}{\sigma^{2}_{\gamma^{\prime}}},$ (S28)
-where $\frac{\partial t_{\gamma}(\textbf{y})}{\partial\gamma}=\left(\frac{\partial t_{\gamma}(y_{1})}{\partial\gamma},\frac{\partial t_{\gamma}({y}_{2})}{\partial\gamma},\ldots,\frac{\partial t_{\gamma}({y}_{n})}{\partial\gamma}\right)^{\top}$, with the $i^{th}$ elements of the vector is equal to
+where $\frac{\partial t_{\gamma}(\textbf{y})}{\partial\gamma}=\left(\frac{\partial t_{\gamma}(y_{1})}{\partial\gamma},\frac{\partial t_{\gamma}({y}_{2})}{\partial\gamma},\ldots,\frac{\partial t_{\gamma}({y}_{n})}{\partial\gamma}\right)^{\top}$, with the $i^{\textrm{th}}$ elements of this vector is equal to
 $\frac{\partial t_{\gamma}({y}_{i})}{\partial\gamma}=\begin{cases}\frac{(y_{i}+1)^{\gamma}(\gamma\textrm{log}(y_{i}+1)-1)+1}{\gamma^{2}}&\text{if }y_{i}\geq 0,\\ \frac{(2-\gamma)(-y_{i}+1)^{(2-\gamma)}\textrm{log}(-y_{i}+1)-(-y_{i}+1)^{2-\gamma}+1}{(2-\gamma)^{2}}&\text{if }y_{i}<0,\end{cases}$
@@
 S2.2.4 Gradients for YJ-SEM-t
-For YJ-SEM-t, $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$.
+For the YJ-SEM-t, we set $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$, so that $h(\boldsymbol{\xi})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$.
 The logarithm of $h(\boldsymbol{\xi})$ is $\text{log}~p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\text{log}~p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\text{log}~p(\boldsymbol{\phi})$.
-Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, $\nu$, and $\gamma$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3.
+Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, $\nu$, and $\gamma$, we utilise the transformations described in Section S2.2, and place the priors on the transformed parameters as in Table S3.
 This leads to
@@
 the derivative of $\text{log}~h(\boldsymbol{\xi})$ with respect to $\nu^{\prime}$ is
-$\begin{split}\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\nu^{\prime}}&=n\left(\frac{\partial~{}\textrm{log}~{}\Gamma\left(\frac{e^{\nu^{\prime}+3}}{2}\right)}{\partial\nu^{\prime}}\right)-\frac{e^{\nu^{\prime}}}{2}\sum_{i=1}^{n}\left(\tau_{i}^{\prime}+e^{-\tau_{i}^{\prime}}\right)-\frac{ne^{\nu^{\prime}}}{2}\textrm{log}\left(\frac{2}{e^{\nu^{\prime}}+3}\right)+\frac{ne^{\nu^{\prime}}}{2}\\ &-\frac{\nu^{\prime}}{\sigma^{2}_{\nu^{\prime}}},\\ \end{split}$ (S33)
+[…]
 and the derivative of $\text{log}~h(\boldsymbol{\xi})$ with respect to $\gamma^{\prime}$ is
 $\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\gamma^{\prime}}=\left(-e^{-\omega^{\prime}}\textbf{r}^{\top}\textbf{M}\left(\frac{\partial t_{\gamma}(\textbf{y})}{\partial\gamma}\right)+\sum_{i=0}^{n}\frac{\partial\textrm{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right)}{\partial\gamma}\right)\left(\frac{\partial\gamma}{\partial\gamma^{\prime}}\right)-\frac{\gamma^{\prime}}{\sigma^{2}_{\gamma^{\prime}}}.$ (S34)
-The derivative of $\text{log}~{}h(\boldsymbol{\xi})$ with respect to $i^{th}$ element of the vector $\boldsymbol{\tau}^{\prime}$, $\tau_{i}^{\prime}$ is
+The derivative of $\text{log}~h(\boldsymbol{\xi})$ with respect to $i^{\textrm{th}}$ element of the vector $\boldsymbol{\tau}^{\prime}$, $\tau_{i}^{\prime}$ is
 $\frac{\partial\text{log}h(\boldsymbol{\xi})}{\partial\tau_{i}^{\prime}}=-\frac{1}{2}+\frac{1}{2}e^{-\omega^{\prime}}({\textbf{r}_{\textbf{A}}}_{i})^{2}e^{\tau_{i}^{\prime}}+\left(\frac{e^{\nu^{\prime}+3}}{2}\right)\left(-1+e^{-\tau_{i}^{\prime}}\right),$ (S35)
-where ${\textbf{r}_{\textbf{A}}}_{i}$ is the $i^{th}$ element of the vector Ar.
+where ${\textbf{r}_{\textbf{A}}}_{i}$ is the $i^{\textrm{th}}$ element of the vector Ar.
 S3 Calculate adaptive learning rates using ADADELTA
-The adaptive learning rates (step sizes) for the VB algorithms are calculated using the ADADELTA algorithm (Zeiler,, 2012).
-The ADADELTA algorithm is now briefly described.
-Different step sizes are used for each element in variational parameters $\boldsymbol{\lambda}$.
-The update for the $i^{th}$ element of $\boldsymbol{\lambda}$ is
-$\boldsymbol{\lambda}^{(t+1)}_{i}=\boldsymbol{\lambda}^{(t)}_{i}+\Delta\boldsymbol{\lambda}^{(t)}_{i},$ (S36)
-where, the step size $\Delta\boldsymbol{\lambda}^{(t)}_{i}$ is $a_{i}^{(t)}g_{\boldsymbol{\lambda}_{i}}^{(t)}$.
-The term $g_{\boldsymbol{\lambda}_{i}}^{(t)}$ denotes the $i^{th}$ component of $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$ and $a_{i}^{(t)}$ is defined as:
-$\mathcal{a}_{i}^{(t)}=\sqrt{\frac{E\left(\boldsymbol{\Delta}_{\boldsymbol{\lambda}_{i}}^{2}\right)^{(t-1)}+\alpha}{E\left(g^{2}_{\boldsymbol{\lambda}_{i}}\right)^{(t)}+\alpha}},$ (S37)
-where $\alpha$ is a small positive constant, $E\left(\boldsymbol{\Delta}_{\boldsymbol{\lambda_{i}}}^{2}\right)^{(t)}$ and $E\left(g^{2}_{\boldsymbol{\lambda}_{i}}\right)^{(t)}$ are decayed moving average estimates of ${\boldsymbol{\Delta}{\boldsymbol{\lambda_{i}}}^{(t)}}^{2}$ and ${g_{\boldsymbol{\lambda}_{i}}^{(t)}}^{2}$, defined by
-$E\left(\boldsymbol{\Delta}_{\boldsymbol{\lambda}_{i}}^{2}\right)^{(t)}=\upsilon E\left(\boldsymbol{\Delta}_{\boldsymbol{\lambda}_{i}}^{2}\right)^{(t-1)}+(1-\upsilon){\boldsymbol{\Delta}{\boldsymbol{\lambda_{i}}}^{(t)}}^{2},$ (S38)
+To ensure stable and efficient convergence of all VB algorithms presented in this paper, we use the ADADELTA algorithm (Zeiler,, 2012) to compute adaptive learning rates $\mathbcal{a}^{(t)}$.
+ADADELTA assigns different step sizes to each element of the variational parameter vector $\boldsymbol{\lambda}$.
+A brief description of the ADADELTA algorithm is provided below.
+The update for the $i^{\textrm{th}}$ element of $\boldsymbol{\lambda}$, $\lambda_{i}$ is
+${\lambda}^{(t+1)}_{i}={\lambda}^{(t)}_{i}+\Delta{\lambda}^{(t)}_{i}$, (S36)
+where, the step size $\Delta{\lambda}^{(t)}_{i}$ is $a_{i}^{(t)}g_{{\lambda}_{i}}^{(t)}$.
+The term $g_{{\lambda}_{i}}^{(t)}$ denotes the $i^{th}$ component of $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$ and $a_{i}^{(t)}$ is defined as:
+$\mathcal{a}_{i}^{(t)}=\sqrt{\frac{\mathbb{E}\left({\Delta}_{{\lambda}_{i}}^{2}\right)^{(t-1)}+\alpha}{\mathbb{E}\left(g^{2}_{{\lambda}_{i}}\right)^{(t)}+\alpha}},$ (S37)
+where $\alpha$ is a small positive constant, $\mathbb{E}\left({\Delta}_{{\lambda_{i}}}^{2}\right)^{(t)}$ and $\mathbb{E}\left(g^{2}_{{\lambda}_{i}}\right)^{(t)}$ are decayed moving average estimates of ${{\Delta}{{\lambda_{i}}}^{(t)}}^{2}$ and ${g_{{\lambda}_{i}}^{(t)}}^{2}$, defined by
+$\mathbb{E}\left({\Delta}_{{\lambda}_{i}}^{2}\right)^{(t)}=\upsilon\mathbb{E}\left({\Delta}_{{\lambda}_{i}}^{2}\right)^{(t-1)}+(1-\upsilon){{\Delta}{{\lambda_{i}}}^{(t)}}^{2},$ (S38)
 and
-$E\left(g^{2}_{\boldsymbol{\lambda}_{i}}\right)^{(t)}=\upsilon E\left(g^{2}_{\boldsymbol{\lambda}_{i}}\right)^{(t-1)}+(1-\upsilon){g_{\boldsymbol{\lambda}_{i}}^{(t)}}^{2},$ (S39)
+$\mathbb{E}\left(g^{2}_{{\lambda}_{i}}\right)^{(t)}=\upsilon\mathbb{E}\left(g^{2}_{{\lambda}_{i}}\right)^{(t-1)}+(1-\upsilon){g_{{\lambda}_{i}}^{(t)}}^{2},$ (S39)
 where the variable $\upsilon$ is a decay constant.
-We use the default tuning parameter choices $\alpha=10^{-6}$ and $\upsilon=0.95$, and initialize $E\left(\boldsymbol{\Delta}_{\boldsymbol{\lambda}_{i}}^{2}\right)^{(0)}=E\left(g^{2}_{\boldsymbol{\lambda}_{i}}\right)^{(0)}=0$.
-S4 Additional Derivation of HVB algorithms for estimating SEMs with missing data
-S4.1 Derivation of the reparameterisation gradient for HVB algorithm
-Since $\mathcal{L}(\boldsymbol{\lambda})=E_{q}\left[\text{log}~{}p(\textbf{y}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})+\text{log}~{}p(\boldsymbol{\xi},\boldsymbol{\psi})-\text{log}~{}q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\right]=\mathcal{L}^{0}(\boldsymbol{\lambda})$ as shown in Equation (16) of the main paper.
-Then, the reparameterisation gradient of $\mathcal{L}$ is the same as that of $\mathcal{L}^{0}$, given by
-$\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$ $=E_{f_{\boldsymbol{\delta}^{0}}}\Bigg{[}\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top}}{d\boldsymbol{\lambda_{s_{m}}}}\Big{(}\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}p(\boldsymbol{\xi},\boldsymbol{\psi})$ (S40)
-$\qquad+\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}p(\textbf{y}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})-\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\theta})\Big{)}\Bigg{]},$
-where, the random vector $\boldsymbol{\delta}^{0}$ has density $f_{\boldsymbol{\delta}^{0}}$, which follows a standard normal, and does not depend on $\boldsymbol{\lambda_{s_{m}}}$, and $t^{0}$ is the one-to-one vector-valued transformation from $\boldsymbol{\delta}^{0}=({\boldsymbol{\eta}^{0}}^{\top},{\boldsymbol{\epsilon}^{0}}^{\top})^{\top}$ to the parameter vector, such that $(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}=t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})=\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}$.
+We use the default tuning parameter choices $\alpha=10^{-6}$ and $\upsilon=0.95$, and initialise $\mathbb{E}\left({\Delta}_{{\lambda}_{i}}^{2}\right)^{(0)}=\mathbb{E}\left(g^{2}_{{\lambda}_{i}}\right)^{(0)}=0$.
+S4 Selection model factorisation for SAR models
+This section outlines how the selection model factorisation for SEMs, discussed in Section 3.2 of the main paper, can be straightforwardly extended to other simultaneous autoregressive (SAR)–type models.
+A wide range of SAR-type models exists, including the spatial error model (SEM), the spatial autoregressive model (SAM), and the spatial Durbin model (SDM), the latter of which incorporates spatially lagged covariates (Anselin,, 1988; Cressie,, 1993).
+The selection model factorisation is a highly general decomposition and can be applied to any of these SAR-type models with only minor modifications.
+For illustration, consider the SDM:
+$\textbf{y}=\rho\textbf{W}\textbf{y}+\textbf{X}\boldsymbol{\beta}_{1}+\textbf{X}_{s}\boldsymbol{\beta}_{2}+\textbf{e},$ (S40)
+where $\boldsymbol{\beta}_{1}=(\beta_{10},\dots,\beta_{1r})^{\top}$ is the $(r+1)\times 1$ vector of fixed effects associated with the covariates in the design matrix X.
+The $n\times r$ matrix $\textbf{X}_{s}=\textbf{W}\textbf{X}^{o}$ represents the spatially lagged covariates, where $\textbf{X}^{o}$ contains the covariates excluding the intercept column.
+The vector $\boldsymbol{\beta}_{2}=(\beta_{21},\dots,\beta_{2r})^{\top}$ is the $r\times 1$ vector of fixed effects corresponding to the spatially lagged covariates, and $\textbf{e}\sim N(0,\sigma^{2}_{\textbf{e}}\textbf{I}_{n})$.
+The parameter vector for the SDM with Gaussian errors is therefore $\boldsymbol{\phi}=(\boldsymbol{\beta}_{1}^{\top},\boldsymbol{\beta}_{2}^{\top},\rho,\sigma^{2}_{\textbf{e}})^{\top}$}.
+In the Gaussian case, $\boldsymbol{\xi}=\boldsymbol{\phi}$.
+If we instead assume Student’s-$t$ errors, $e_{i}\sim t_{\nu}(0,\sigma^{2}_{\textbf{e}})$, then the parameter vector becomes $\boldsymbol{\phi}=(\boldsymbol{\beta}_{1}^{\top},\boldsymbol{\beta}_{2}^{\top},\rho,\sigma^{2}_{\textbf{e}},\nu)^{\top}$}, and $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$, which is consistent with the extension of the SEMs to Student’s-$t$ errors (SEM-t) described in Section 2.1 of the main paper.
+Importantly, the same selection model factorisation used in Equation (13) of the main paper applies directly to both the Gaussian SDM and its Student’s-$t$ extension, with the only difference being the form of the parameter vector $\boldsymbol{\xi}$.
+Specifically, the joint density of the missingness indicators m and the response vector y is factorised as
+$p(\textbf{y},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})=p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})\,p(\textbf{y}\mid\boldsymbol{\xi}).$ (S41)
+Thus, the methodological framework developed in the main paper, namely, extending SEMs to non-Gaussian settings, jointly modelling the missingness indicators and the spatial process via the selection model factorisation, and performing inference using the HVB algorithm, is fully general and can be readily adapted to other SAR-type models with alternative spatial dependence structures.
+S5 Additional derivations for the HVB algorithm
+This section provides additional proofs and derivations for the proposed HVB algorithm used to estimate non-Gaussian SEMs with missing data, introduced in Section 4.2 of the main paper.
+S5.1 Derivation of the reparameterisation gradient
+Since $\mathcal{L}(\boldsymbol{\lambda})=\mathbb{E}_{q_{\boldsymbol{\lambda}}}\left[\text{log}~p(\textbf{y}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})+\text{log}~p(\boldsymbol{\xi},\boldsymbol{\psi})-\text{log}~q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\right]=\mathcal{L}^{0}(\boldsymbol{\lambda})$, as shown in Equation (19) of the main paper, the reparameterisation gradient of $\mathcal{L}$ is the same as that of $\mathcal{L}^{0}$ and given by
+$\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$ $=\mathbb{E}_{f_{\boldsymbol{\delta}^{0}}}\Bigg[\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})^{\top}}{d\boldsymbol{\lambda}}\Big(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~p(\boldsymbol{\xi},\boldsymbol{\psi})$ (S42)
+$\qquad+\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~p(\textbf{y}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})-\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\Big)\Bigg],$
+where, the random vector $\boldsymbol{\delta}^{0}$ has density $f_{\boldsymbol{\delta}^{0}}$, which follows a standard normal, and does not depend on $\boldsymbol{\lambda}$, and $t^{0}$ is the one-to-one vector-valued transformation from $\boldsymbol{\delta}^{0}=({\boldsymbol{\eta}^{0}}^{\top},{\boldsymbol{\epsilon}^{0}}^{\top})^{\top}$} to the parameter vector, such that $(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}=t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})=\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0}$}.
 Then, the Fisher’s identity is given by
@@
 see Poyiadjis et al., (2011).
-Substituting this expression into Equation (S40), and writing $E_{f_{\boldsymbol{\delta}}}(\cdot)$ for expectation with respect to $f_{\boldsymbol{\delta}}(\boldsymbol{\delta})=f_{\boldsymbol{\delta}^{0}}(\boldsymbol{\delta}^{0})p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ and because $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}_{o},\textbf{m}\mid\textbf{y}_{u},\boldsymbol{\xi},\boldsymbol{\psi})p(\textbf{y}_{u}\mid\boldsymbol{\xi})p(\boldsymbol{\xi},\boldsymbol{\psi})$, we get
-$\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$ $={E}_{f_{\boldsymbol{\delta}}}\Bigg{[}\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top}}{d\boldsymbol{\lambda}_{s_{m}}}\Big{(}\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log p(\boldsymbol{\xi},\boldsymbol{\psi})$ (S42)
+Substituting this expression into Equation (S42), and writing $\mathbb{E}_{f_{\boldsymbol{\delta}}}[\cdot]$ for expectation with respect to $f_{\boldsymbol{\delta}}(\boldsymbol{\delta})=f_{\boldsymbol{\delta}^{0}}(\boldsymbol{\delta}^{0})p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ and because $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}_{o},\textbf{m}\mid\textbf{y}_{u},\boldsymbol{\xi},\boldsymbol{\psi})p(\textbf{y}_{u}\mid\boldsymbol{\xi})p(\boldsymbol{\xi},\boldsymbol{\psi})$, we get
+$\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$ $=\mathbb{E}_{f_{\boldsymbol{\delta}}}\Bigg[\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})^{\top}}{d\boldsymbol{\lambda}}\Big(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log p(\boldsymbol{\xi},\boldsymbol{\psi})$ (S44)
 $\qquad+\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log p(\textbf{y}_{u}\mid\boldsymbol{\xi})+\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log p(\textbf{y}_{o},\textbf{m}\mid\textbf{y}_{u},\boldsymbol{\xi},\boldsymbol{\psi})$
 $\qquad-\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\Big)\Bigg]$
-$={E}_{f_{\boldsymbol{\delta}}}\left[\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top}}{d\boldsymbol{\lambda}_{s_{m}}}\left(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})-\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\right)\right].$
-The term $\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})}{d\boldsymbol{\lambda_{s_{m}}}}$ in Equation (S42) is the derivative of the transformation $t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})=\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}$ with respect to the variational parameters $\boldsymbol{\lambda}_{s_{m}}=(\boldsymbol{\mu}_{s_{m}}^{\top},\text{vech}(\textbf{B}_{s_{m}})^{\top},\textbf{d}_{s_{m}}^{\top})^{\top}$.
-We can express that $t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})=\boldsymbol{\mu}_{s_{m}}+(\boldsymbol{\eta}^{0}\otimes\textbf{I}_{s_{m}})\text{vech}(\textbf{B}_{s_{m}})+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}$, where $\textbf{I}_{s_{m}}$ is the identity matrix of size $s_{m}$, and it can be further shown that $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}q^{0}_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi})=-(\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2})^{-1}((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}-\boldsymbol{\mu}_{s_{m}})$,
-$\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})}{d\boldsymbol{\mu}_{s_{m}}}=\textbf{I}_{s_{m}}\\ ~{}~{}~{}~{}~{}\text{and}~{}~{}~{}~{}~{}\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})}{d\text{vech}(\textbf{B}_{s_{m}})}={\boldsymbol{\eta}^{0}}^{\top}\otimes\textbf{I}_{s_{m}}.$ (S43)
-Then the derivatives of $\mathcal{L}(\boldsymbol{\lambda})$ with respect to $\boldsymbol{\mu}_{s_{m}},\textrm{vech}(\textbf{B}_{s_{m}})$, and $\textbf{d}_{s_{m}}$ are:
-$\begin{split}\nabla_{\boldsymbol{\mu}_{s_{m}}}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\delta}}}[\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi^{\top})^{\top}}}~{}\text{log}~{}h(\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u})\\ &+(\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2})^{-1}(\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0})],\end{split}$ (S44)
-$\begin{split}\nabla_{\textrm{vech}(\textbf{B}_{s_{m}})}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\delta}}}[\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi^{\top})^{\top}}}~{}\text{log}~{}h(\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u}){\boldsymbol{\eta}^{0}}^{\top}\\ &+(\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2})^{-1}(\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}){\boldsymbol{\eta}^{0}}^{\top}],\end{split}$ (S45)
-$\begin{split}\nabla_{\textbf{d}_{s_{m}}}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\delta}}}[\text{diag}(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi^{\top})^{\top}}}\text{log}~{}h(\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u}){\boldsymbol{\epsilon}^{0}}^{\top}\\ &+(\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2})^{-1}(\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}){\boldsymbol{\epsilon}^{0}}^{\top})].\end{split}$ (S46)
-The expressions for $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u})=\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ in Equations (S44)-(S46) can be found in Section S4.3.
-The expectations in these gradients are estimated using a single sample of $\boldsymbol{\delta}=({\boldsymbol{\delta}^{0}}^{\top},\textbf{y}_{u}^{\top})^{\top}$, which is drawn from $f_{\boldsymbol{\delta}^{0}}$, and $p(\textbf{y}_{u}\mid\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{o},\textbf{m})=p(\textbf{y}_{u}\mid t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}}),\textbf{y}_{o},\textbf{m})$; see  Wijayawardhana et al., (2025) for further details.
-S4.2 HVB-AllB algorithm
-This section presents the HVB-AllB algorithm for estimating SEMs with missing values, which is particularly suited for cases where both $n$ and $n_{u}$ are large.
+$=\mathbb{E}_{f_{\boldsymbol{\delta}}}\left[\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})^{\top}}{d\boldsymbol{\lambda}}\left(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})-\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\right)\right].$
+S5.2 Derivation of $\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})}{d\boldsymbol{\lambda}}$, $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$, and $\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$ for HVB
+In the reparameterisation gradient of the ELBO given in Equation (S44), $\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})}{d\boldsymbol{\lambda}}$ denotes the derivative of the transformation $t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})=\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0}$} with respect to the variational parameters $\boldsymbol{\lambda}=(\boldsymbol{\mu}^{\top},\text{vech}(\textbf{B})^{\top},\textbf{d}^{\top})^{\top}$}.
+We can express that $t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})=\boldsymbol{\mu}+(\boldsymbol{\eta}^{0}\otimes\textbf{I}_{s})\text{vech}(\textbf{B})+\textbf{d}\circ\boldsymbol{\epsilon}^{0}$}, where $\textbf{I}_{s}$ is the identity matrix of size $s$.
+It can be further shown that $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~q^{0}_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi})=-(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}-\boldsymbol{\mu})$,
+$\frac{\partial t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})}{\partial\boldsymbol{\mu}}=\textbf{I}_{s}\\ ,~~~~~~~~~~\frac{\partial t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})}{\partial\text{vech}(\textbf{B})}={\boldsymbol{\eta}^{0}}^{\top}\otimes\textbf{I}_{s}\\ ,~~~~~\text{and}~~~~~\frac{\partial t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})}{\partial\textbf{d}}=\boldsymbol{\epsilon}^{0}.$ (S45)
+Then the derivatives of $\mathcal{L}(\boldsymbol{\lambda})$ with respect to $\boldsymbol{\mu},\textrm{vech}(\textbf{B})$, and d are:
+$\begin{split}\nabla_{\boldsymbol{\mu}}\mathcal{L}(\boldsymbol{\lambda})&=\mathbb{E}_{f_{\boldsymbol{\delta}}}[\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi^{\top})^{\top}}}~\text{log}~h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u})\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0})],\end{split}$ (S46)
+$\begin{split}\nabla_{\textrm{vech}(\textbf{B})}\mathcal{L}(\boldsymbol{\lambda})&=\mathbb{E}_{f_{\boldsymbol{\delta}}}[\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi^{\top})^{\top}}}~\text{log}~h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u}){\boldsymbol{\eta}^{0}}^{\top}\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0}){\boldsymbol{\eta}^{0}}^{\top}],\end{split}$ (S47)
+$\begin{split}\nabla_{\textbf{d}}\mathcal{L}(\boldsymbol{\lambda})&=\mathbb{E}_{f_{\boldsymbol{\delta}}}[\text{diag}(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi^{\top})^{\top}}}\text{log}~h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u}){\boldsymbol{\epsilon}^{0}}^{\top}\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0}){\boldsymbol{\epsilon}^{0}}^{\top})].\end{split}$ (S48)
+The expressions for $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u})=\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$} in Equations (S46)-(S48) can be found in Section S5.3.
+The expectations in these gradients are estimated using a single sample of $\boldsymbol{\delta}=({\boldsymbol{\delta}^{0}}^{\top},\textbf{y}_{u}^{\top})^{\top}$, which is drawn from $f_{\boldsymbol{\delta}^{0}}$, and $p(\textbf{y}_{u}\mid\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{o},\textbf{m})=p(\textbf{y}_{u}\mid t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}),\textbf{y}_{o},\textbf{m})$; see  Wijayawardhana et al., (2025) for further details.
+S5.3 Expressions for $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$
+This section presents the derivatives (gradients) $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$, which are required to compute the gradients of the reparameterised ELBO, defined in Equation (20) of the main paper.
+S5.3.1 Gradients for SEM-Gau
+For the SEM-Gau with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$}, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}.$}
+The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~p(\textbf{y}\mid\boldsymbol{\phi})+\textrm{log}~p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~p(\boldsymbol{\phi})+\textrm{log}~p(\boldsymbol{\psi})$}.
+Note that, for $\sigma^{2}_{\textbf{e}}$, and $\rho$, we utilise the transformations described in Section S2.2, and place the priors on the transformed parameters as in Table S3.
+This leads to
+$\begin{split}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}\\ &+\sum_{i=1}^{n}m_{i}({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{\boldsymbol{\psi}^{\top}\boldsymbol{\psi}}{2\sigma^{2}_{\boldsymbol{\psi}}},\end{split}$ (S49)
+where the values of $\sigma^{2}_{\boldsymbol{\beta}}$, $\sigma^{2}_{\omega^{\prime}}$, $\sigma^{2}_{\rho^{\prime}}$, and $\sigma^{2}_{\boldsymbol{\psi}}$ are each set to 100, as detailed in Table S3, while r and M are provided in Table S1.
+The derivatives of $\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ in Equation (S49) with respect to $\boldsymbol{\beta}$, $\omega^{\prime}$}, and $\rho^{\prime}$} are similar to that of without missing data given in Equations (S15),  (S16), and (S17), respectively.
+The derivative of $\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ with respect to $\boldsymbol{\psi}$ is
+$\frac{\partial\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})}{\partial\boldsymbol{\psi}}=\sum_{i=1}^{n}\left(m_{i}-\frac{e^{{\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y}}}{1+e^{{\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y}}}\right)\textbf{z}_{i}-\frac{\boldsymbol{\psi}^{\top}}{\sigma^{2}_{\boldsymbol{\psi}}},$ (S50)
+where $\textbf{z}_{i}=({\textbf{x}^{*}_{i}}^{\top},y_{i})$, $\textbf{x}^{*}_{i}$$^{\top}$ is the $i^{\textrm{th}}$ row of matrix $\textbf{X}^{*}$, and $y_{i}$ is the $i^{\textrm{th}}$ element of the vector y; see Section 3 of the main paper.
+S5.3.2 Gradients for SEM-t
+For the SEM-t with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$}, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\nu,\rho)^{\top}$.
+The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\textrm{log}~p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\textrm{log}~p(\boldsymbol{\phi})+\textrm{log}~p(\boldsymbol{\psi})$.
+Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, and $\nu$, we utilise the transformations described in Section S2.2, and place the priors on the transformed parameters as in Table S3.
+This leads to
+$\begin{split}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}\\ &+\sum_{i=1}^{n}m_{i}({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &+\sum_{i=1}^{n}\textrm{log}~p(\tau_{i}\mid\nu)+\textrm{log}\left|\frac{\partial\tau_{i}}{\partial\tau_{i}^{\prime}}\right|-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\nu^{\prime}}^{2}}{2\sigma^{2}_{\nu^{\prime}}},\end{split}$ (S51)
+where the values of $\sigma^{2}_{\boldsymbol{\beta}}$, $\sigma^{2}_{\omega^{\prime}}$, $\sigma^{2}_{\rho^{\prime}}$, $\sigma^{2}_{\nu^{\prime}}$, and $\sigma^{2}_{\boldsymbol{\psi}}$ are each set to 100, as detailed in Table S3, while r and M are provided in Table S1.
+The derivatives of $\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ in Equation (S51) with respect to $\boldsymbol{\beta}$, $\omega^{\prime}$}, $\rho^{\prime}$}, $\nu^{\prime}$} and $\tau_{i}^{\prime}$ are similar to that of without missing data given in Equations (S19),  (S20), (S21), (S22), and (S23), respectively.
+The derivative of $\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ with respect to $\boldsymbol{\psi}$ is similar to that of SEM-Gau given in Equation (S50).
+S5.3.3 Gradients for YJ-SEM-Gau
+For the YJ-SEM-Gau with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$}, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\gamma)^{\top}$.
+The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~p(\textbf{y}\mid\boldsymbol{\phi})+\textrm{log}~p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~p(\boldsymbol{\phi})+\textrm{log}~p(\boldsymbol{\psi})$}.
+Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, and $\gamma$, we utilise the transformations described in in Section S2.2, and place the priors on the transformed parameters as in Table S3.
+This leads to
+$\begin{split}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right)\\ &+\sum_{i=1}^{n}m_{i}({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\gamma^{\prime}}^{2}}{2\sigma^{2}_{\gamma^{\prime}}}-\frac{\boldsymbol{\psi}^{\top}\boldsymbol{\psi}}{2\sigma^{2}_{\boldsymbol{\psi}}},\end{split}$ (S52)
+where the values of $\sigma^{2}_{\boldsymbol{\beta}}$, $\sigma^{2}_{\omega^{\prime}}$, $\sigma^{2}_{\rho^{\prime}}$, $\sigma^{2}_{\nu^{\prime}}$, $\sigma^{2}_{\gamma^{\prime}}$, and $\sigma^{2}_{\boldsymbol{\psi}}$ are each set to 100, as detailed in Table S3, while r and M are provided in Table S1.
+The derivatives of $\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ in Equation (S52) with respect to $\boldsymbol{\beta}$, $\omega^{\prime}$}, $\rho^{\prime}$}, and $\gamma^{\prime}$} are similar to that of without missing data given in Equations (S25),  (S26), (S27), and (S28), respectively.
+The derivative of $\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ with respect to $\boldsymbol{\psi}$ is similar to that of SEM-Gau given in Equation (S50).
+S5.3.4 Gradients for YJ-SEM-t
+For the YJ-SEM-t with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$}, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$.
+The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\textrm{log}~p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\textrm{log}~p(\boldsymbol{\phi})+\textrm{log}~p(\boldsymbol{\psi})$.
+Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, $\nu$, and $\gamma$, we utilise the transformations described in Section S2.2, and place the priors on the transformed parameters as in Table S3.
+This leads to
+$\begin{split}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right)\\ &+\sum_{i=1}^{n}m_{i}({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &+\sum_{i=1}^{n}\textrm{log}~p(\tau_{i}\mid\nu)+\textrm{log}\left|\frac{\partial\tau_{i}}{\partial\tau_{i}^{\prime}}\right|-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\nu^{\prime}}^{2}}{2\sigma^{2}_{\nu^{\prime}}}-\frac{{\gamma^{\prime}}^{2}}{2\sigma^{2}_{\gamma^{\prime}}},\end{split}$ (S53)
+where the values of $\sigma^{2}_{\boldsymbol{\beta}}$, $\sigma^{2}_{\omega^{\prime}}$, $\sigma^{2}_{\rho^{\prime}}$, $\sigma^{2}_{\nu^{\prime}}$, $\sigma^{2}_{\gamma^{\prime}}$, and $\sigma^{2}_{\boldsymbol{\psi}}$ are each set to 100, as detailed in Table S3, while r and M are provided in Table S1.
+The derivatives of $\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ in Equation S53 with respect to $\boldsymbol{\beta}$, $\omega^{\prime}$}, $\rho^{\prime}$}, $\nu^{\prime}$}, $\gamma^{\prime}$} and $\tau_{i}^{\prime}$ are similar to that of without missing data given in Equations (S30),  (S31), (S32), (S33), (S34), and (S35), respectively.
+The derivative of $\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ with respect to $\boldsymbol{\psi}$ is similar to that of SEM-Gau given in Equation (S50).
+S5.4 HVB-AllB algorithm
+This section presents the MCMC steps used in the HVB-AllB algorithm for estimating SEMs with missing values, which is particularly suited for cases where both $n$ and $n_{u}$ are large.
 The MCMC steps in Algorithm 2 of the main paper generates samples from the conditional distribution $p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$.
@@
 We start with partitioning the unobserved responses vector into $k$ blocks, such that $\textbf{y}_{u}=(\textbf{y}_{u_{1}}^{\top},\dots,\textbf{y}_{u_{k}}^{\top})^{\top}$.
-Using proposals from $p(\textbf{y}_{u_{j}}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{o},\textbf{y}_{u}^{(-j)})$, the Algorithm S2 outlines the MCMC steps for sampling the missing values one block at a time, for $j=1,..,k$, where $\textbf{y}_{u_{j}}$ is the updated block and $\textbf{y}_{u}^{(-j)}=(\textbf{y}_{u_{1}}^{\top},\dots,\textbf{y}_{u_{j-1}}^{\top},\textbf{y}_{u_{j+1}}^{\top},\dots,\textbf{y}_{u_{k}}^{\top})^{\top}$ is the remaining blocks.
+Using proposals from $p(\textbf{y}_{u_{j}}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{o},\textbf{y}_{u}^{(-j)})$, Algorithm S2 outlines the MCMC steps for sampling the missing values one block at a time, for $j=1,..,k$, where $\textbf{y}_{u_{j}}$ is the updated block and $\textbf{y}_{u}^{(-j)}=(\textbf{y}_{u_{1}}^{\top},\dots,\textbf{y}_{u_{j-1}}^{\top},\textbf{y}_{u_{j+1}}^{\top},\dots,\textbf{y}_{u_{k}}^{\top})^{\top}$ is the remaining blocks.
 The complete response vector y can now be written as $\textbf{y}=({\textbf{y}^{\top}_{s_{j}}},\textbf{y}_{u_{j}}^{\top})^{\top}$, where ${\textbf{y}_{s_{j}}}=(\textbf{y}_{o}^{\top},{\textbf{y}_{u}^{(-j)}}^{\top})^{\top}$.
@@
 Similarly, $\textbf{X}_{s_{j}}$ and $\textbf{X}_{u_{j}}$ are represent the sub-matrices of X, and $\textbf{M}_{{s_{j}}{s_{j}}}$, $\textbf{M}_{{s_{j}}{u_{j}}}$, $\textbf{M}_{{u_{j}}{s_{j}}}$, and $\textbf{M}_{{u_{j}}{u_{j}}}$ are sub-matrices of M.
+Further, r and M for different SEMs are provided in Table S1.
 Algorithm S2 MCMC steps within the $t^{th}$ iteration of the HVB algorithm under MNAR. The missing values are updated one block at a time.
 1: Initialise missing values $\textbf{y}_{u,0}=(\textbf{y}^{\top}_{u_{1},0},\dots,\textbf{y}^{\top}_{u_{k},0})^{\top}~\sim p(\textbf{y}_{u}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{o})$
-2:  for i=1, …, $N_{1}$ do
-3:     for j=1, …, $k$ do
+2:for $i=1,\dots,N_{1}$ do
+3:  for $j=1,\dots,k$ do
 4:   Sample $\widetilde{\textbf{y}}_{u_{j}}$ from the proposal distribution $p(\widetilde{\textbf{y}}_{u_{j}}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{o},\textbf{y}_{u,{i-1}}^{(-j)})$, where $\textbf{y}_{u,i-1}^{(-j)}=(\textbf{y}^{\top}_{u_{1},i},\textbf{y}^{\top}_{u_{2},i},\ldots,\textbf{y}^{\top}_{u_{j-1},i},\textbf{y}^{\top}_{u_{j+1},i-1},\ldots,,\textbf{y}^{\top}_{u_{k},i-1})^{\top}$}.
 5:   Sample $u$ from uniform distribution, $u~\sim\mathcal{U}(0,1)$
-6:        Calculate $a=\text{min}\left(1,\frac{p(\textbf{m}\mid\widetilde{\textbf{y}},\boldsymbol{\psi}^{(t)})}{p(\textbf{m}\mid\textbf{y}_{i-1},\boldsymbol{\psi}^{(t)})}\right)$, where $\widetilde{\textbf{y}}=(\textbf{y}_{o}^{\top},{\textbf{y}_{u}^{(-j)}}^{\top},{\widetilde{\textbf{y}}_{u_{j}}}^{\top})^{\top}$ and $\textbf{y}_{i-1}=(\textbf{y}_{o}^{\top},{\textbf{y}_{u}^{(-j)}}^{\top},{\textbf{y}_{u_{j},{i-1}}}^{\top})^{\top}$
+6:   Calculate $a=\text{min}\left(1,\frac{p(\textbf{m}\mid\widetilde{\textbf{y}},\boldsymbol{\psi}^{(t)})}{p(\textbf{m}\mid\textbf{y}_{i-1},\boldsymbol{\psi}^{(t)})}\right)$, where $\widetilde{\textbf{y}}=(\textbf{y}_{o}^{\top},{\textbf{y}_{u,i-1}^{(-j)}}^{\top},{\widetilde{\textbf{y}}_{u_{j}}}^{\top})^{\top}$ and $\textbf{y}_{i-1}=(\textbf{y}_{o}^{\top},{\textbf{y}_{u,i-1}^{(-j)}}^{\top},{\textbf{y}_{u_{j},{i-1}}}^{\top})^{\top}$}
 7:   if $a>u$ then
-8:           $\textbf{y}_{u_{j},i}=\textbf{y}_{u_{j}}^{*}$
+8:    $\textbf{y}_{u_{j},i}=\widetilde{\textbf{y}}_{u_{j}}$
 9:   else
@@
 15: Output $\textbf{y}^{(t)}_{u}=\textbf{y}_{u,N_{1}}$
-In step 4 of the Algorithm S2, proposals from the conditional distributions, $p(\widetilde{\textbf{y}}_{u_{j}}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{s_{j}})$ needs to be generated.
+In step 4 of Algorithm S2, proposals from the conditional distributions, $p(\widetilde{\textbf{y}}_{u_{j}}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{o},\textbf{y}_{u,{i-1}}^{(-j)})={p(\widetilde{\textbf{y}}_{u_{j}}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{s_{j}})}$} needs to be generated.
 For the SEM-Gau, and SEM-t, this distribution follows a multivariate Gaussian distribution with the mean $\textbf{X}_{u_{j}}\boldsymbol{\beta}-\textbf{M}_{{u_{j}}{u_{j}}}^{-1}\textbf{M}_{{u_{j}}{s_{j}}}\textbf{r}_{s_{j}}$ and the covariance matrix $\sigma^{2}_{\textbf{e}}\textbf{M}_{{u_{j}}{u_{j}}}^{-1}$.
+Expressions of r and M for SEM-Gau and SEM-t are provided in Table S1, and details on the partitioning of r, X, and M are given in Equation (S54).
 For YJ-SEM-Gau and YJ-SEM-t, the proposals are generated in two steps.
@@
 Then, we apply the inverse Yeo-Johnson (YJ) transformation to obtain the final proposal: $\widetilde{\textbf{y}}_{u_{j}}=t_{\gamma}^{-1}(\widetilde{\textbf{y}}_{u_{j}}^{*})$.
-S4.3 Expressions for $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$
-This section presents the derivatives $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$, which are required to compute the gradients of the reparameterised ELBO, defined in Equation (17) of the main paper for different SEMs with missing data.
-S4.3.1 Derivatives for SEM-Gau
-For SEM-Gau with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}$.
-The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})+\textrm{log}~{}p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~{}p(\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\psi})$.
-Note that, for $\sigma^{2}_{\textbf{e}}$, and $\rho$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3.
-This leads to
-$\begin{split}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}\\ &+\sum_{i=1}^{n}m_{i}(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{\boldsymbol{\psi}^{\top}\boldsymbol{\psi}}{2\sigma^{2}_{\boldsymbol{\psi}}},\end{split}$ (S48)
-where the values of $\sigma^{2}_{\boldsymbol{\beta}}$, $\sigma^{2}_{\omega^{\prime}}$, $\sigma^{2}_{\rho^{\prime}}$, and $\sigma^{2}_{\boldsymbol{\psi}}$ are each set to 100, as detailed in Table S3, while r and M are provided in Table S1.
-The derivatives of $\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ in Equation (S48) with respect to $\boldsymbol{\beta}$, ${\omega^{\prime}}$, and ${\rho^{\prime}}$ are similar to that of without missing data given in Equations (S15),  (S16), and (S17), respectively.
-The derivative of $\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ with respect to $\boldsymbol{\psi}$ is
-$\frac{\partial\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})}{\partial\boldsymbol{\psi}}=\sum_{i=1}^{n}\left(m_{i}-\frac{e^{{\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y}}}{1+e^{{\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y}}}\right)\textbf{z}_{i}-\frac{\boldsymbol{\psi}^{\top}}{\sigma^{2}_{\boldsymbol{\psi}}},$ (S49)
-where $\textbf{z}_{i}=({\textbf{x}^{*}_{i}},y_{i})$ is the vector containing the $i^{th}$ row vector of matrix $\textbf{X}^{*}$, and the $i^{th}$ element of the vector y, see Section 2.3 of the main paper.
-S4.3.2 Derivatives for SEM-t
-For SEM-t with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\nu,\rho)^{\top}$.
-The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~{}p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\textrm{log}~{}p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~{}p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\psi})$.
-Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, and $\nu$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3.
-This leads to
-$\begin{split}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}\\ &+\sum_{i=1}^{n}m_{i}(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &+\sum_{i=1}^{n}\textrm{log}~{}p(\tau_{i}\mid\nu)+\textrm{log}\left|\frac{\partial\tau_{i}}{\partial\tau_{i}^{\prime}}\right|-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\nu^{\prime}}^{2}}{2\sigma^{2}_{\nu^{\prime}}},\end{split}$ (S50)
-where the values of $\sigma^{2}_{\boldsymbol{\beta}}$, $\sigma^{2}_{\omega^{\prime}}$, $\sigma^{2}_{\rho^{\prime}}$, $\sigma^{2}_{\nu^{\prime}}$, and $\sigma^{2}_{\boldsymbol{\psi}}$ are each set to 100, as detailed in Table S3, while r and M are provided in Table S1.
-The derivatives of $\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ in Equation (S50) with respect to $\boldsymbol{\beta}$, ${\omega^{\prime}}$, ${\rho^{\prime}}$, ${\nu^{\prime}}$ and $\tau_{i}^{\prime}$ are similar to that of without missing data given in Equations (S19),  (S20), (S21), (S22), and (S23), respectively.
-The derivative of $\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ with respect to $\boldsymbol{\psi}$ is similar to that of SEM-Gau given in Equation (S49).
-S4.3.3 Derivatives for YJ-SEM-Gau
-For YJ-SEM-Gau with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\gamma)^{\top}$.
-The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})+\textrm{log}~{}p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~{}p(\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\psi})$.
-Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, and $\gamma$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3.
-This leads to
-$\begin{split}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right)\\ &+\sum_{i=1}^{n}m_{i}(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\gamma^{\prime}}^{2}}{2\sigma^{2}_{\gamma^{\prime}}}-\frac{\boldsymbol{\psi}^{\top}\boldsymbol{\psi}}{2\sigma^{2}_{\boldsymbol{\psi}}},\end{split}$ (S51)
-where the values of $\sigma^{2}_{\boldsymbol{\beta}}$, $\sigma^{2}_{\omega^{\prime}}$, $\sigma^{2}_{\rho^{\prime}}$, $\sigma^{2}_{\nu^{\prime}}$, $\sigma^{2}_{\gamma^{\prime}}$, and $\sigma^{2}_{\boldsymbol{\psi}}$ are each set to 100, as detailed in Table S3, while r and M are provided in Table S1.
-The derivatives of $\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ in Equation (S51) with respect to $\boldsymbol{\beta}$, ${\omega^{\prime}}$, ${\rho^{\prime}}$, and ${\gamma^{\prime}}$ are similar to that of without missing data given in Equations (S25),  (S26), (S27), and (S28), respectively.
-The derivative of $\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ with respect to $\boldsymbol{\psi}$ is similar to that of SEM-Gau given in Equation (S49).
-S4.3.4 Derivatives for YJ-SEM-t
-For YJ-SEM-t with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$.
-The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~{}p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\textrm{log}~{}p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~{}p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\psi})$.
-Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, $\nu$, and $\gamma$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3.
-This leads to
-$\begin{split}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right)\\ &+\sum_{i=1}^{n}m_{i}(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &+\sum_{i=1}^{n}\textrm{log}~{}p(\tau_{i}\mid\nu)+\textrm{log}\left|\frac{\partial\tau_{i}}{\partial\tau_{i}^{\prime}}\right|-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\nu^{\prime}}^{2}}{2\sigma^{2}_{\nu^{\prime}}}-\frac{{\gamma^{\prime}}^{2}}{2\sigma^{2}_{\gamma^{\prime}}},\end{split}$ (S52)
-where the values of $\sigma^{2}_{\boldsymbol{\beta}}$, $\sigma^{2}_{\omega^{\prime}}$, $\sigma^{2}_{\rho^{\prime}}$, $\sigma^{2}_{\nu^{\prime}}$, $\sigma^{2}_{\gamma^{\prime}}$, and $\sigma^{2}_{\boldsymbol{\psi}}$ are each set to 100, as detailed in Table S3, while r and M are provided in Table S1.
-The derivatives of $\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ in Equation S52 with respect to $\boldsymbol{\beta}$, ${\omega^{\prime}}$, ${\rho^{\prime}}$, ${\nu^{\prime}}$, ${\gamma^{\prime}}$ and $\tau_{i}^{\prime}$ are similar to that of without missing data given in Equations (S30),  (S31), (S32), (S33), (S34), and (S35), respectively.
-The derivative of $\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ with respect to $\boldsymbol{\psi}$ is similar to that of SEM-Gau given in Equation (S49).
-S5 Bayesian model comparison of spatial error models
-This section presents expressions for the term log $p(\mathbf{y}\mid\boldsymbol{\phi})$, which are used to compute the Deviance Information Criterion (DIC) (Celeux et al.,, 2006) for the various SEMs discussed in Section 3.3 of the main paper.
-Note that for SEM-Gau, and YJ-SEM-Gau, $\boldsymbol{\xi}=\boldsymbol{\phi}$, meaning for these two models $p(\mathbf{y}\mid\boldsymbol{\phi})=p(\mathbf{y}\mid\boldsymbol{\xi})$.
-For the SEM-Gau model, $p(\mathbf{y}\mid\boldsymbol{\xi})$ is Gaussian; see Section 2.1 of the main paper.
-The log-likelihood is given by:
+S6 Bayesian model comparison for SEMs
+This section presents expressions for the term log $p(\textbf{y}\mid\boldsymbol{\phi})$ for different SEMs.
+These terms are used to compute the deviance information criterion (DIC) (Celeux et al.,, 2006) values discussed in Section 5 of the main paper.
+Note that for the SEM-Gau, and YJ-SEM-Gau, $\boldsymbol{\xi}=\boldsymbol{\phi}$, meaning for these two models $p(\textbf{y}\mid\boldsymbol{\phi})=p(\textbf{y}\mid\boldsymbol{\xi})$.
+For the SEM-Gau, $p(\textbf{y}\mid\boldsymbol{\xi})$ is multivariate Gaussian; see Section 2.1 of the main paper.
+Corresponding log-likelihood is given by:
 $\text{log}~p(\textbf{y}\mid\boldsymbol{\phi})=-\frac{n}{2}\textrm{log}(2\pi)-\frac{n}{2}\textrm{log}(\sigma^{2}_{\boldsymbol{e}})+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{1}{2\sigma^{2}_{\boldsymbol{e}}}\textbf{r}^{\top}\textbf{M}\textbf{r},$ (S55)
-where M, and r are given in Table S1.
-For the YJ-SEM-Gau, $p(\mathbf{y}\mid\boldsymbol{\phi})=p(\mathbf{y}\mid\boldsymbol{\xi})$ is derived in Section 2.2 of the main paper, and the log-likelihood is given by:
+where M and r are given in Table S1.
+For the YJ-SEM-Gau, $p(\textbf{y}\mid\boldsymbol{\phi})=p(\textbf{y}\mid\boldsymbol{\xi})$ is derived in Section 2.2 of the main paper. Corresponding log-likelihood is given by:
 $\begin{split}\text{log}~p(\textbf{y}\mid\boldsymbol{\phi})&=-\frac{n}{2}\textrm{log}(2\pi)-\frac{n}{2}\textrm{log}(\sigma^{2}_{\boldsymbol{e}})+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{1}{2\sigma^{2}_{\boldsymbol{e}}}\textbf{r}^{\top}\textbf{M}\textbf{r}\\ &+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right),\end{split}$ (S56)
 where M, and r are given in Table S1, while $\frac{dt_{\gamma}(y_{i})}{dy_{i}}$ is given in Equation (S2).
-For the SEM-t model,
-[…]
-where $p(\mathbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})$ is multivariate Gaussian, with its mean vector and covariance matrix specified in Table S1.
-Each $p({\tau}_{i}\mid\boldsymbol{\phi})$ (for $i=1,\ldots,n$) follows an inverse gamma distribution with both shape and scale parameters equal to $\nu/2$; see Section 2.1 of the main paper.
+For the SEM-t,
+$p(\textbf{y}\mid\boldsymbol{\phi})=\int_{0}^{\infty}\ldots\int_{0}^{\infty}p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p({\tau}_{1}\mid\boldsymbol{\phi})\ldots p({\tau}_{n}\mid\boldsymbol{\phi})d{\tau}_{1}\ldots d{\tau}_{n}$ (S57)
+where $p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})=p(\textbf{y}\mid\boldsymbol{\xi})$ is multivariate Gaussian, with its mean vector and covariance matrix specified in Table S1.
+Each $p({\tau}_{i}\mid\boldsymbol{\phi})$ (for $i=1,\ldots,n$) follows an inverse gamma distribution where both the shape parameter and the rate parameter are set to $\nu/2$; see Section 2.1 of the main paper.
 This hierarchical representation implies that the marginal distribution $p(\mathbf{y}\mid\boldsymbol{\phi})$ is a multivariate Student’s $t$-distribution with mean vector $\textbf{X}\boldsymbol{\beta}$, scale matrix $\sigma^{2}_{\boldsymbol{e}}(\mathbf{A}^{\top}\mathbf{A})^{-1}$, and degrees of freedom $\nu$.
-The log-likelihood is given by:
+Then, the corresponding log-likelihood is given by:
 $\begin{split}\text{log}~p(\textbf{y}\mid\boldsymbol{\phi})&=\textrm{log}~\Gamma\left(\frac{\nu+2}{2}\right)-\textrm{log}~\Gamma\left(\frac{\nu}{2}\right)+\frac{1}{2}\textrm{log}|\textbf{A}^{\top}\textbf{A}|\\ &-\frac{n}{2}\textrm{log}(\pi\nu\sigma^{2}_{\textbf{e}})-\frac{(n+\nu)}{2}\textrm{log}\left(1+\frac{1}{\nu\sigma^{2}_{\boldsymbol{e}}}\textbf{r}^{\top}\textbf{A}^{\top}\textbf{A}\textbf{r}\right),\end{split}$ (S58)
-where $\textbf{r}=\textbf{y}-\textbf{X}\boldsymbol{\beta}$.
-Similarly, logarithm of $p(\textbf{y}\mid\boldsymbol{\phi})$ for YJ-SEM-t can be derived as
+where $\textbf{r}=\textbf{y}-\textbf{X}\boldsymbol{\beta}$, and $\Gamma(\cdot)$ is the gamma function.
+Similarly, the logarithm of $p(\textbf{y}\mid\boldsymbol{\phi})$ for YJ-SEM-t can be derived as
 $\begin{split}\text{log}~p(\textbf{y}\mid\boldsymbol{\phi})&=\textrm{log}~\Gamma\left(\frac{\nu+2}{2}\right)-\textrm{log}~\Gamma\left(\frac{\nu}{2}\right)+\frac{1}{2}\textrm{log}|\textbf{A}^{\top}\textbf{A}|\\ &-\frac{n}{2}\textrm{log}(\pi\nu\sigma^{2}_{\textbf{e}})-\frac{(n+\nu)}{2}\textrm{log}\left(1+\frac{1}{\nu\sigma^{2}_{\boldsymbol{e}}}\textbf{r}^{\top}\textbf{A}^{\top}\textbf{A}\textbf{r}\right)\\ &+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right),\end{split}$ (S59)
 where $\textbf{r}=t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$.
-S6 Simulation study with full data SEMs
-This section fits the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models to simulated datasets 1 and 2 (without missing data) in Section 4.2 of the main paper, and evaluates how effectively these models, along with their VB estimation methods, capture the characteristics of these datasets.
-The initial values for the VB algorithms are set as follows: The variational mean vector ($\boldsymbol{\mu}$) is initialised using the estimates of $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ obtained from fitting the SEM-Gau model via maximum likelihood estimation (MLE).
-The initial values for $\nu$ and $\gamma$ are set to 4 and 1, respectively.
+S7 Additional figures for simulation study: Assessing the accuracy of VB methods in Section 6
+Section 6 of the main paper compares the posterior density estimates of the model parameters for YJ-SEM-Gau obtained using VB (and HVB when missing values are present) and HMC (Neal,, 2011), based on a simulated dataset analysed both with and without missing values.
+To generate the data, the model parameters are set as follows: the fixed effects ($\boldsymbol{\beta}$) are randomly drawn from discrete uniform values between 3-3 and $3$ (excluding 0).
+The error variance parameter is set to $\sigma^{2}_{\mathbf{e}}=1$ and the spatial autoregressive parameter is $\rho=0.8$.
+The Yeo-Johnson transformation parameter is set to $\gamma=1.25$, which introduces moderate left skewness in the simulated response variable.
+The kernel density plot of the simulated response variable is provided in Figure S1.
+As expected, the density is moderately skewed to the left.
+Figure S1: Kernel density plot of the simulated response variable generated using the YJ-SEM-Gau with $\gamma=1.25$ in the simulation study described in Section 6 of the main paper.
+S8 Simulation study: Assessing the accuracy and robustness of proposed SEMs (full data case)
+This section fits SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t to simulated datasets 1 and 2 (without missing data) described in Section 7 of the main paper, and evaluates how effectively these models, along with their VB estimation methods, capture the characteristics of these datasets.
+The initial values for the VB algorithms are set as follows: the elements of the variational mean vector ($\boldsymbol{\mu}$) corresponding to $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ are initialised using estimates obtained by fitting the SEM-Gau via maximum likelihood (ML) estimation.
+The variational means corresponding to the parameters $\nu$ and $\gamma$ are initialised to 4 and 1, respectively.
+For the models with Student’s $t$ errors, the set of latent variables $\boldsymbol{\tau}=(\tau_{1},\dots,\tau_{n})^{\top}$ must also be initialised.
+The variational means corresponding to these latent variables are initialised by sampling from the prior distribution $p(\tau_{i}\mid\nu^{(0)})\sim\text{IG}(\frac{\nu^{(0)}}{2},\frac{\nu^{(0)}}{2})$} for $i=1,\dots,n$, where $\nu^{(0)}=4$ denotes the initial value of $\nu$.
 For the variational covariance matrix parameters, the elements of B and the diagonal elements of D are all initialised to 0.01.
 We use $p=4$ factors.
-After some experimentation, we found that the models with Student’s $t$ errors required a relatively large number of iterations to converge.
-This increased computational demand is attributed to the added complexity of estimating both the latent variables $\boldsymbol{\tau}$ and the degrees of freedom parameter $\nu$.
-For simulated dataset 1, the VB algorithms were run for 10,000 iterations for both SEM-Gau and YJ-SEM-Gau, and 20,000 iterations for SEM-t and YJ-SEM-t.
-For simulated dataset 2, convergence was achieved after 10,000 iterations for SEM-Gau and YJ-SEM-Gau, and 40,000 iterations for SEM-t and YJ-SEM-t.
-Convergence is assessed visually by inspecting the variational mean plots over iterations, which are presented in Section S9.1.
-Then, we generate 10,000 draws of model parameters from the variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$.
-Table S4 shows the posterior means and the 95% credible intervals for some of the model parameters of the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models obtained using the Gaussian variational approximation method described in Section S2 applied to the simulated dataset 1.
-The table also includes the computation times (in seconds) for one VB iteration, as well as the $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values for each model, calculated using Equations (18) and (19) of the main paper, respectively.
-The plots of posterior densities of the parameters are provided in Section S7.2.1.
-We begin by comparing the estimated values of the fixed effects (specifically $\beta_{0}$ and $\beta_{1}$), the variance parameter $\sigma_{\textbf{e}}^{2}$, and the spatial autocorrelation parameter $\rho$, as these parameters are present in all four models.
-Accurate estimation of these parameters is crucial for interpreting the underlying spatial relationships and the effect of covariates on the response variable in SEMs.
-The posterior means for $\beta_{0}$, $\beta_{1}$ and $\rho$ are nearly identical for YJ-SEM-Gau and YJ-SEM-t, closely matching the true values of $-1$ for $\beta_{0}$, $1$ for $\beta_{1}$, and $0.8$ for $\rho$.
-In contrast, the estimates from SEM-Gau and SEM-t show significant deviations from the true values.
-The estimated posterior means of $\sigma_{\textbf{e}}^{2}$ from SEM-Gau, SEM-t, and YJ-SEM-Gau deviate considerably from the true value, while the posterior mean obtained from YJ-SEM-t is much closer to the true value, as expected.
-The models incorporating the YJ transformation, YJ-SEM-Gau and YJ-SEM-t, successfully recover the parameter $\gamma$.
-The degrees of freedom ($\nu$) estimated by the SEM-t and YJ-SEM-t models show slight deviations from the true value.
-For the simulated dataset 1, the model fit assessment indicates that YJ-SEM-t is the best-performing model, as it achieves the lowest $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values among all four models.
-This outcome is expected, given that the dataset was simulated from YJ-SEM-t.
-The second-best model is YJ-SEM-Gau, which has the next lowest DIC values.
-Notably, the difference in DIC between YJ-SEM-t and YJ-SEM-Gau is minimal, suggesting that YJ-SEM-Gau provides a model fit nearly as good as YJ-SEM-t for this dataset.
+[…]
+S8.1 Results for the dataset 1
+Table S4 shows the posterior means and the 95% credible intervals for some of the model parameters of the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t obtained using the VB algorithm applied to the simulated dataset 1.
+The table also includes the computation times (in seconds) for one VB iteration, as well as the $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values for each model, calculated using Equations (21) and (22), respectively, in Section 5 of the main paper.
+Table S4: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 1 (full data), along with the computation time (CT) in seconds for one VB iteration.
+The table also presents the DIC values for each model.
+Parameters labelled as ’NA’ indicate that they are not applicable to the corresponding model.
 SEM-Gau SEM-t YJ-SEM-Gau YJ-SEM-t
@@
 CT 0.1322 0.1206 0.6715 0.7252
-Table S4: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 1 (full data), along with the computation time (CT) in seconds for one VB iteration.
-The table also presents the DIC values for each model.
-Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
-Table S5 presents the posterior means and 95% credible intervals for selected model parameters obtained from the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models applied to the simulated dataset 2, along with the computational cost for one VB iteration.
+We begin by comparing the estimated values of the fixed effects (specifically $\beta_{0}$ and $\beta_{1}$), the variance parameter $\sigma_{\textbf{e}}^{2}$, and the spatial autocorrelation parameter $\rho$, as these parameters are present in all four models.
+The posterior means for $\beta_{0}$, $\beta_{1}$ and $\rho$ are nearly identical for YJ-SEM-Gau and YJ-SEM-t, closely matching the true values of 1-1 for $\beta_{0}$, $1$ for $\beta_{1}$, and $0.8$ for $\rho$.
+In contrast, the estimates from SEM-Gau and SEM-t show significant deviations from the true values.
+The estimated posterior means of $\sigma_{\textbf{e}}^{2}$ from SEM-Gau, SEM-t, and YJ-SEM-Gau deviate considerably from the true value, while the posterior mean obtained from YJ-SEM-t is much closer to the true value.
+The models incorporating the YJ transformation, YJ-SEM-Gau and YJ-SEM-t, successfully recover the parameter $\gamma$.
+The degrees of freedom ($\nu$) estimated by the SEM-t and YJ-SEM-t models show slight deviations from the true value.
+For the simulated dataset 1, the model fit assessment indicates that YJ-SEM-t is the best-performing model, as it achieves the lowest $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values among all four models.
+This outcome is expected, given that the dataset is simulated from YJ-SEM-t.
+The second-best model is YJ-SEM-Gau, which has the next lowest DIC values.
+Notably, the difference in DIC between YJ-SEM-t and YJ-SEM-Gau is significantly small, suggesting that YJ-SEM-Gau provides a model fit nearly as good as YJ-SEM-t for the dataset 1.
+Figure S2 compares the posterior densities of selected model parameters obtained using the proposed VB method, applied to different SEMs for the simulated dataset 1.
+As expected, the YJ-SEM-t recovers the true parameter values more accurately than the other three models, since dataset 1 is simulated using the YJ-SEM-t.
+Figure S2: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t, fitted to the simulated dataset 1 (without missing data), using the VB method.
+The vertical dotted lines indicate the true parameter values.
+S8.2 Results for the dataset 2
+Table S5 presents the posterior means and 95% credible intervals for selected model parameters obtained from the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t applied to the simulated dataset 2, along with the computation cost for one VB iteration.
 The table also includes the $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values for each model.
-Corresponding posterior density plots of the parameters are provided in Section S7.2.1.
-Unlike the simulated dataset 1, which is skewed and heavy tails, the simulated dataset 2 is more symmetrical with lighter tails.
+Unlike the simulated dataset 1, which is skewed and has heavy tails, the simulated dataset 2 is more symmetrical with lighter tails.
 The posterior mean estimates for parameters $\beta_{0}$, $\beta_{1}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ across all four models closely match their true values.
 Likewise, the posterior mean estimate of $\nu$ for SEM-t and YJ-SEM-t is close to the true value of $30$, while the posterior mean estimate of $\gamma$ for YJ-SEM-Gau and YJ-SEM-t is close to the true value of $1$.
-The overall accuracy of parameter estimation across all models is further supported by the nearly identical $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values, with SEM-Gau having the lowest values, indicating the best fit for the simulated dataset 2, as expected.
-According to the computation times presented in Tables S4 and S5, VB algorithms for both SEM-t and YJ-SEM-t, which incorporate Student’s $t$ errors, require significantly more computation time compared to SEM-Gau and YJ-SEM-Gau.
-This is because, in models with $t$ errors, the VB optimisation is substantially more complex than in models with Gaussian errors.
-Specifically, for models with Student’s $t$ errors, the set of model parameters is $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$, whereas for models with Gaussian errors, it simplifies to $\boldsymbol{\xi}=\boldsymbol{\phi}$.
-Here, $\boldsymbol{\tau}$ is a latent variable vector of length $n$; see Section 2.1 of the main paper.
+The overall accuracy of parameter estimation across all models is further supported by the nearly identical $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values, with SEM-Gau having the lowest values, indicating the best fit for the simulated dataset 2.
+Table S5: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 2 (full data), along with the computation time (CT) in seconds for one VB iteration.
+The table also presents the DIC values for each model.
+Parameters labelled as ’NA’ indicate that they are not applicable to the corresponding model.
 SEM-Gau SEM-t YJ-SEM-Gau YJ-SEM-t
-${\beta}_{0}=-3$2.9976 (-3.0258, -2.9692) -3.0008 (-3.0305, -2.9707) -2.9781 (-3.0108, -2.9449) -2.9808 (-3.0033, -2.9584)
+${\beta}_{0}=-3$ -2.9976 (-3.0258, -2.9692) -3.0008 (-3.0305, -2.9707) -2.9781 (-3.0108, -2.9449) -2.9808 (-3.0033, -2.9584)
 ${\beta}_{1}=2$ 1.9968 (1.9773, 2.0161) 1.9947 (1.9774, 2.0121) 1.9920 (1.9739, 2.0104) 1.9950 (1.9716, 2.0183)
@@
 CT 0.1280 0.6718 0.1197 0.7241
-Table S5: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 2 (full data), along with the computation time (CT) in seconds for one VB iteration.
-The table also presents the DIC values for each model.
-Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
-S7 Additional figures for simulation study in Section 4
-This section provides additional figures related to the simulation study presented in Section 4 of the main paper, and Section S6.
-S7.1 Comparison of VB and HVB methods
-Section 4.1 of the main paper compares the estimated posterior densities of the model parameters for the YJ-SEM-Gau model using VB (including HVB) and HMC (Neal,, 1996) methods, both with and without missing data, based on a simulated dataset.
-To generate the data, the model parameters are set as follows: the fixed effects ($\boldsymbol{\beta}$) are randomly drawn from discrete uniform values between $-3$ and $3$ (excluding 00).
-The error variance parameter is set to $\sigma^{2}_{\mathbf{e}}=1$ and the spatial autoregressive parameter is $\rho=0.8$.
-The Yeo-Johnson transformation parameter is set to $\gamma=1.25$, which introduces moderate left skewness in the simulated response variable.
-The kernel density plot of the simulated response variable is provided in Figure S1.
-As expected, the density is moderately skewed to the left.
-Figure S1: Kernel density plot of the simulated response variable generated using the YJ-SEM-Gau model with $\gamma=1.25$ in the simulation study of Section 4.1 of the main paper.
-S7.2 Evaluating the performance of proposed SEMs
-S7.2.1 Simulation study with full data
-This section presents further figures from Section S6.
-Figures S2 and S3 compare the posterior densities of selected model parameters obtained using the proposed VB method, applied to different SEMs (without missing values) for simulated dataset 1 and simulated dataset 2, respectively.
-As shown in Figure S2, the YJ-SEM-t model is able to recover the true parameter values more accurately than the other three models, as expected, since the simulated dataset 1 was simulated using the YJ-SEM-t model.
-In contrast, as shown in Figure S3, all four models successfully recover the true values of the parameters.
-Figure S2: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the simulated dataset 1, using the VB method.
+Figure S3 compares the posterior densities of selected model parameters obtained using the proposed VB method, applied to different SEMs for simulated dataset 2.
+It can be seen that, all four models successfully recover the true values of the parameters.
+Figure S3: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t, fitted to the simulated dataset 2 (without missing data), using the VB method.
 The vertical dotted lines indicate the true parameter values.
-Figure S3: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the simulated dataset 2, using the VB method.
+S8.3 Summary and computation time comparison
+According to the computation times presented in Tables S4 and S5, VB algorithms for both SEM-t and YJ-SEM-t, which incorporate Student’s $t$ errors, require significantly more computation time per iteration compared to SEM-Gau and YJ-SEM-Gau.
+This is because, in models with Student’s $t$ errors, the VB optimisation is substantially more complex than in models with Gaussian errors.
+Specifically, for models with Student’s $t$ errors, the set of model parameters is $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$, whereas for models with Gaussian errors, it simplifies to $\boldsymbol{\xi}=\boldsymbol{\phi}$.
+Here, $\boldsymbol{\tau}$ is a latent variable vector of length $n$; see Section 2.1 of the main paper.
+In addition to the higher per-iteration cost, SEM-t and YJ-SEM-t also require more VB iterations to achieve convergence.
+Together, these factors lead to substantially higher computational costs for models with Student’s $t$ errors than for their Gaussian-error counterparts.
+In terms of estimation performance, the results indicate that both YJ-SEM-Gau and YJ-SEM-t are particularly effective at capturing the key characteristics of non-Gaussian data.
+For simulated dataset 1, which exhibits skewness and heavy tails, YJ-SEM-t achieves the lowest DIC values, indicating the best fit, while YJ-SEM-Gau attains the second-lowest values.
+The differences between the DIC values of YJ-SEM-t and YJ-SEM-Gau are small, suggesting that YJ-SEM-Gau provides a model fit comparable to that of YJ-SEM-t.
+For simulated dataset 2, which is closer to Gaussian, all four models recover the true parameter values accurately, and the DIC values are nearly identical, with SEM-Gau achieving the lowest values and therefore providing the best fit.
+S9 Additional figures for simulation study: Assessing the accuracy and robustness of proposed SEMs (missing data case in Section 7.1)
+This section provides additional figures supporting the analysis in Section 7.1 of the main paper, which evaluates the accuracy and robustness of the proposed SEMs and HVB (HVB-AllB) methods using two simulated datasets with missing values.
+S9.1 Comparison of posterior densities of model parameters
+Figures S4 and S5 present posterior densities of selected model parameters from different SEMs, estimated using the proposed HVB-AllB method on simulated datasets 1 and 2 with missing values.
+Figure S4 shows that YJ-SEM-t recovers the true parameter values most accurately for simulated dataset 1, as expected, since this dataset was generated from the YJ-SEM-t.
+In contrast, Figure S5 demonstrates that, for simulated dataset 2, all four models recover the true parameter values well.
+Figure S4: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t, fitted to the simulated dataset 1 with missing data, using the HVB-AllB method.
 The vertical dotted lines indicate the true parameter values.
-S7.2.2 Simulation study with missing data
-This section presents further figures from Section 4.2 of the main paper.
-Figures S4 and S5 compare the posterior densities of selected model parameters estimated using the proposed HVB method, applied to different SEMs with missing values, for simulated dataset 1 and simulated dataset 2, respectively.
-As shown in Figure S4, the YJ-SEM-t model is able to recover the true parameter values more accurately than the other three models, as expected, since the simulated dataset 1 was simulated using the YJ-SEM-t model.
-In contrast, as shown in Figure S5, all four models successfully recover the true values of the parameters.
-Figure S4: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the simulated dataset 1 with missing data, using the HVB method.
+Figure S5: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t, fitted to the simulated dataset 2 with missing data, using the HVB-AllB method.
 The vertical dotted lines indicate the true parameter values.
-Figure S5: Posterior densities of some of the parameters for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the simulated dataset 2 with missing data, using the HVB method.
-The vertical dotted lines indicate the true parameter values.
-The left panel of Figure S6 shows the kernel densities of the true missing values ($\mathbf{y}_{u}$) and the kernel density of posterior means of the missing values at locations $s_{1},s_{2},\ldots,s_{n}$, estimated from different SEMs for simulated dataset 2.
+S9.2 Comparison of posterior densities of missing values
+Figure S6 compares the posterior means and standard deviations of the missing values $\textbf{y}_{u}={(y_{u_{1}},\ldots,y_{u_{n_{u}}})}^{\top}$ estimated by the HVB-AllB algorithms under different SEMs for the simulated dataset 1 with missing values.
+The left panel compares the posterior means of the predicted missing values with their true values.
+The right panel compares the posterior standard deviations of the missing values estimated under the YJ-SEM-t (x-axis) with those obtained from the SEM-Gau, SEM-t, and YJ-SEM-Gau (y-axis).
+The posterior means of the missing values estimated under the YJ-SEM-Gau and YJ-SEM-t are closer to the true missing values compared to those estimated under the SEM-Gau and SEM-t.
+The posterior standard deviations of the missing values from the YJ-SEM-Gau and YJ-SEM-t are nearly identical, whereas those from the SEM-Gau and SEM-t deviate from the standard deviations obtained by the models with YJ transformation.
+Figure S6: Comparison of posterior means and standard deviations of missing values for the simulated dataset 1: The left panel displays the comparison of the posterior means of predicted missing values from the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t with true missing values.
+The right panel compares the posterior standard deviations of the missing values from SEM-Gau, SEM-t, and YJ-SEM-Gau against those obtained from YJ-SEM-t.
+Figure S7 compares the posterior means and standard deviations of the missing values $\textbf{y}_{u}={(y_{u_{1}},\ldots,y_{u_{n_{u}}})}^{\top}$ estimated by the HVB-AllB algorithms under different SEMs for the simulated dataset 2 with missing values.
+The left panel compares the posterior means of the predicted missing values with their true values.
+The right panel compares the posterior standard deviations of the missing values estimated under the YJ-SEM-t (x-axis) with those obtained from the SEM-Gau, SEM-t, and YJ-SEM-Gau (y-axis).
+The posterior means of the missing values estimated under all four models are close to the true values.
+Additionally, the posterior standard deviations from the SEM-Gau, SEM-t, and YJ-SEM-Gau closely align with those from the YJ-SEM-t, with YJ-SEM-Gau showing the closest match.
+Figure S7: Comparison of posterior means and standard deviations of predicted missing values for the simulated dataset 2: The left panel displays the comparison of the posterior means of missing values from the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t with true missing values.
+The right panel compares the posterior standard deviations of the missing values from SEM-Gau, SEM-t, and YJ-SEM-Gau against those obtained from YJ-SEM-t.
+The left four panels of Figure S8 compare the kernel density of the true missing values ($\textbf{y}_{u}$) with the kernel densities of the posterior means of the missing values estimated under different SEMs for the simulated dataset 2.
 All four models produce nearly identical densities of posterior means of missing values that closely match the density of the true missing values.
-The right panel of Figure S6 presents the posterior density of the maximum missing value ($\textbf{y}_{u_{\textrm{max}}}$) predicted by each SEM.
-The true value of $\textbf{y}_{u_{\textrm{max}}}$ lies within the posterior distribution estimated by all models.
-These results are consistent with expectations, as all four models produce similar parameter estimates and DIC values (see Table 5).
+The right panel of Figure S8 presents the posterior density of the maximum missing value, $\textrm{max}(\textbf{y}_{u})$, estimated under each SEM.
+The true value of $\textrm{max}(\textbf{y}_{u})$ lies within the posterior distribution estimated by all models.
+These results are consistent with expectations, as all four models produce similar parameter estimates and $\text{DIC}_{5}$ values (see Table 5 in Section 7.1 of the main paper).
+A corresponding figure for simulated dataset 1 is presented in Figure 4 in Section 7.1 of the main paper.
 Figure S8: Left panel: The kernel densities of the true missing values ($\textbf{y}_{u}$) and the posterior means of the missing values ($\textbf{y}_{u}$) obtained using the HVB-AllB method for different SEMs for the simulated dataset 2 with missing values.
-Right panel: The posterior density of the maximum missing value ($\textbf{y}_{u_{\textrm{max}}}$) predicted by different SEMs for the simulated dataset 2 with missing values.
+Right panel: The posterior density of the maximum missing value, $\textrm{max}(\textbf{y}_{u})$, predicted by different SEMs for the simulated dataset 2 with missing values.
 The true value of the maximum missing value is indicated by the vertical line.
-Figure S7 compares the posterior means and standard deviations of the missing values ($y_{u_{1}},\ldots,y_{u_{n_{u}}}$) estimated by different SEMs for the simulated dataset 1 with missing values.
-The left panel compares the posterior means of the predicted missing values with their true values.
-The right panel compares the posterior standard deviations of the missing values estimated by the YJ-SEM-t model (x-axis) against those obtained from SEM-Gau, SEM-t, and YJ-SEM-Gau models (y-axis).
-The posterior means of the missing values estimated by the YJ-SEM-Gau and YJ-SEM-t models are closer to the true missing values compared to those estimated by the SEM-Gau and SEM-t models.
-Additionally, the posterior standard deviations from the YJ-SEM-Gau and YJ-SEM-t models are nearly identical, whereas those from the SEM-Gau and SEM-t models deviate from the standard deviations obtained by the YJ-based models.
-Figure S7: Comparison of posterior means and standard deviations of missing values for the simulated dataset 1: The left plot shows the comparison of the posterior means of predicted missing values from the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t with true missing values.
-The right plot compares the posterior standard deviations of the missing values from SEM-Gau, SEM-t, and YJ-SEM-Gau against those obtained from YJ-SEM-t.
-Figure S8 compares the posterior means and standard deviations of the missing values ($y_{u_{1}},\ldots,y_{u_{n_{u}}}$) estimated by different SEMs for the simulated dataset 2 with missing values.
-The left panel compares the posterior means of the predicted missing values with their true values.
-The right panel compares the posterior standard deviations of the missing values estimated by the YJ-SEM-t model (x-axis) against those obtained from SEM-Gau, SEM-t, and YJ-SEM-Gau models (y-axis).
-The posterior means of the missing values estimated by all four models are close to the true values.
-Additionally, the posterior standard deviations from the SEM-Gau, SEM-t, and YJ-SEM-Gau models closely align with those from the YJ-SEM-t model, with YJ-SEM-Gau showing the closest match.
-Figure S8: Comparison of posterior means and standard deviations of predicted missing values for the simulated dataset 2: The left plot shows the comparison of the posterior means of missing values from the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t with true missing values.
-The right plot compares the posterior standard deviations of the missing values from SEM-Gau, SEM-t, and YJ-SEM-Gau against those obtained from YJ-SEM-t.
-S8 Additional figures for real data application in Section 5
+S10 Additional figures for real data application in Section 8
 This section provides additional figures related to the real data application discussed in Section 8 of the main paper.
@@
 The right panel shows the kernel density of the logarithm of house prices (in hundreds of thousands).
-Figures S10 and S11 compare the posterior densities of selected model parameters, estimated using the proposed VB and HVB-AllB methods, across different SEMs (without and with missing values) applied to the Lucas-1998-HP dataset.
-Figure S10: Posterior densities of some of the parameters of the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the Lucas-1998-HP dataset without missing values, using the VB method.
-Figure S11: Posterior densities of some of the parameters of the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models, fitted to the Lucas-1998-HP dataset with missing values, using the HVB-AllB method.
-S9 Convergence analysis
+Figures S10 and S11 compare the posterior densities of selected model parameters, estimated using the proposed VB and HVB-AllB methods, across the different SEMs applied to the Lucas-1998-HP dataset, without and with missing values, respectively.
+Figure S10: Posterior densities of some of the parameters of the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t, fitted to the Lucas-1998-HP dataset without missing values, using the VB method.
+Figure S11: Posterior densities of some of the parameters of the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t, fitted to the Lucas-1998-HP dataset with missing values, using the HVB-AllB method.
+S11 Convergence analysis
 For the variational Bayes (VB) and hybrid VB (HVB) algorithms, convergence is assessed by examining the trajectories of the variational means of model parameters across iterations.
 For the Hamiltonian Monte Carlo (HMC) algorithm (Neal et al.,, 2011), convergence is evaluated by inspecting the trace plots of the posterior draws of model parameters.
-Section S9.1 presents the convergence analysis plots for the simulation study described in Section 4 of the main paper.
-Similarly, Section S9.2 provides the convergence analysis plots for the real-data example discussed in Section 5 of the main paper.
-S9.1 Convergence analysis of simulation study
-S9.1.1 Convergence analysis for Section 4.1
-This section presents convergence plots from the simulation study designed to compare the performance of the VB approximation with the HMC algorithm, as described in Section 4.1 of the main paper.
-In this study, data are simulated from the YJ-SEM-Gau model, and the model is estimated using both VB and HMC methods.
-Figure S12 presents trace plots of posterior samples obtained using the HMC algorithm for selected parameters of the YJ-SEM-Gau model with full data, after excluding burn-in samples.
+Section S11.1 presents convergence analysis plots for the simulation study described in Section 6 of the main paper.
+Section S11.2 provides the corresponding convergence analysis plots for the simulation study in Section 7 of the main paper (with missing data), as well as for the full-data case in Section S8.
+Finally, Section S11.3 reports the convergence analysis plots for the real-data example discussed in Section 8 of the main paper.
+S11.1 Convergence analysis for simulation study: Assessing the accuracy of VB methods
+This section presents convergence plots from the simulation study designed to compare the performance of the VB approximation with the HMC algorithm, as described in Section 6 of the main paper.
+In this study, data are simulated from the YJ-SEM-Gau, and the model is estimated using both VB and HMC methods.
+Figure S12 presents trace plots of posterior samples obtained using the HMC algorithm for selected parameters of the YJ-SEM-Gau with full data, after excluding burn-in samples.
 Figure S13 displays the trajectories of the variational means for selected parameters across VB iterations.
 Both plots indicate that the algorithms have converged.
-Figure S12: Trace plots of posterior samples obtained using HMC for selected parameters of the YJ-SEM-Gau model (without missing values), after excluding burn-in samples.
-Figure S13: Trajectories of the variational means of selected parameters of the YJ-SEM-Gau model (without missing values) across VB iterations.
-Next, missing values are introduced into the simulated dataset, and the YJ-SEM-Gau model is fitted using both the HVB-NoB and HMC algorithms.
-Figure S14 presents trace plots of posterior samples obtained using the HMC algorithm for selected parameters of the YJ-SEM-Gau model with missing values, after excluding burn-in samples.
+Figure S12: Trace plots of posterior samples obtained using HMC for selected parameters of the YJ-SEM-Gau (without missing values), after excluding burn-in samples.
+Figure S13: Trajectories of the variational means of selected parameters of the YJ-SEM-Gau (without missing values) across VB iterations.
+Next, missing values are introduced into the simulated dataset, and the YJ-SEM-Gau is fitted using both the HVB-NoB and HMC algorithms.
+Figure S14 presents trace plots of posterior samples obtained using the HMC algorithm for selected parameters of the YJ-SEM-Gau with missing values, after excluding burn-in samples.
 Figure S15 displays the trajectories of the variational means for selected parameters across HVB-NoB iterations.
 Both plots indicate that the algorithms have converged.
-Figure S14: Trace plots of posterior samples obtained using HMC for selected parameters of the YJ-SEM-Gau model (with missing values), after excluding burn-in samples.
-Figure S15: Trajectories of the variational means of selected parameters of the YJ-SEM-Gau model (with missing values) across HVB-NoB iterations.
-S9.1.2 Convergence analysis for Section 4.2
-This section presents convergence plots from the simulation study conducted to evaluate the performance of the proposed SEMs, as described in Section 4.2 of the main paper.
+Figure S14: Trace plots of posterior samples obtained using HMC for selected parameters of the YJ-SEM-Gau (with missing values), after excluding burn-in samples.
+Figure S15: Trajectories of the variational means of selected parameters of the YJ-SEM-Gau (with missing values) across HVB-NoB iterations.
+S11.2 Convergence analysis for simulation study: Assessing the accuracy and robustness of proposed SEMs
+This section presents the convergence plots for the simulation study in Section 7 of the main paper (with missing data), as well as for the full-data case in Section S8.
 Figure S16 shows trajectories of the variational means for selected model parameters of different SEMs fitted to the simulated dataset 1, without missing values.
@@
 Figure S19: Trajectories of the variational means of selected parameters across HVB-AllB iterations for different SEMs fitted on the simulated dataset 2 with missing values.
-S9.2 Convergence analysis of real example
+S11.3 Convergence analysis of real data application
+This section presents the convergence analysis plots for the real data application discussed in Section 8 of the main paper.
 Figure S20 shows trajectories of the variational means for selected model parameters of different SEMs fitted to the Lucas-1998-HP dataset (full data).
+It can be observed that for each parameter in each model, the variational means converge by the end of the final iterations, indicating that the VB algorithms for all models have successfully converged.
+Figure S20: Trajectories of the variational means of selected parameters across VB iterations for different SEMs fitted on the Lucas-1998-HP dataset without missing values.
+Figure S21 shows trajectories of the variational means for selected model parameters of different SEMs fitted to the Lucas-1998-HP dataset, with missing values.
 It can be observed that for each parameter in each model, the variational means converge by the end of the final iterations, indicating that the HVB-AllB algorithms for all models have successfully converged.
-Figure S20: Trajectories of the variational means of selected parameters across HVB-AllB iterations for different SEMs fitted on the Lucas-1998-HP dataset without missing values.
-Figure S21 shows trajectories of the variational means for selected model parameters of different SEMs fitted to the Lucas-1998-HP dataset, with missing values.
-It can be observed that for each parameter in each model, the variational means converge by the end of the final iterations, indicating that the VB algorithms for all models have successfully converged.
 Figure S21: Trajectories of the variational means of selected parameters across HVB-AllB iterations for different SEMs fitted on the Lucas-1998-HP dataset with missing values.